### v1

# High-order Computation of Floquet Multipliers and Subspaces using Multistep Methods

Abstract
Accurate and efficient computation of Floquet multipliers and subspaces is essential for analyzing limit cycle in dynamical systems and periodic steady state in Radio Frequency (RF) simulation. This problem is typically addressed by solving a periodic linear eigenvalue problem, which is discretized from the linear periodic time-varying system using one-step methods. The backward Euler method offers a computationally inexpensive overall workflow but has limited accuracy. In contrast, one-step collocation methods achieve higher accuracy through over-sampling, explicit matrix construction, and condensation, thus become costly for large-scale sparse cases. We apply multistep methods to derive a periodic polynomial eigenvalue problem, which introduces additional spurious eigenvalues. Under mild smoothness assumptions, we prove that as the stepsize decreases, the computed Floquet multipliers and their associated invariant subspace converge with higher order, while the spurious eigenvalues converge to zero. To efficiently solve large-scale problems, we propose pTOAR, a memory-efficient iterative algorithm for computing the dominant Floquet eigenpairs. Numerical experiments demonstrate that multistep methods achieves high order accuracy, while its computational and memory costs are only marginally higher than those of the backward Euler method.
keywords:
Floquet Multiplier, Multistep Method, Periodic Polynomial Eigenvalue Problem, Subspace Perturbation
pacs:
MSC Classification]65L15, 65F15, 65L70
1 Introduction
A linear periodic time-varying (LPTV) system is described by eq.˜1:
$\frac{\mathrm{d}x(t)}{\mathrm{d}t}=G(t)x(t),$ (1)
where $G(t)\in\mathbb{R}^{n\times n}$ is a $T$-periodic matrix function, i.e., $G(t+T)=G(t)$, and $x(t)\in\mathbb{R}^{n}$ is the state vector. The state transition matrix $\Phi(T,0)$ of eq.˜1 is obtained by integrating eq.˜1 from $[0,T]$ with initial value matrix $X(0)=I_{n}$. The Floquet multipliers and eigenvectors are defined as the eigenpairs of $\Phi(T,0)$. This paper applies high-order multistep methods to discretize an LPTV system eq.˜1, for efficient and accurate computation of its Floquet multipliers and corresponding subspace.
The LPTV systems find applications in a wide range of areas, including dynamic simulation [1], communications [2], control theory [3, 4] and Radio Frequency circuit simulation [5]. A central problem in these fields is the stability analysis of limit cycles in nonlinear autonomous systems. The stability is determined by the variational equation along the cycle. Since the limit cycle is $T$-periodic, its associated Jacobian $G(t)$ is also $T$-periodic, making the variational equation an LPTV system. Floquet theory then provides the framework: the stability is completely characterized by the Floquet multipliers and eigenvectors of this system [6].
Depending on different application background and goals, there are two common scenarios in practice. First, when the matrix $G(t;\delta)$ is given as an explicit function of time $t$, with $\delta$ as a parameter, tracking how Floquet multipliers evolve with respect to this parameter provides insight into the system’s stability. In such cases, a multiplier crossing the unit circle from inside to outside typically signals the onset of a bifurcation or instability [6]. Second, when $G(t)$ is only available at discrete time samples, without an explicit analytic form, the analysis often focuses on computing the Floquet eigenvector associated with the dominant multiplier. In applications such as circuit oscillator design, this largest multiplier is typically equal to one [5]. The corresponding left Floquet eigenvector, known as the perturbation projection vector (PPV) [5, 7], characterizes the system’s phase sensitivity to external perturbations and plays a key role in modeling phase noise in Radio Frequency oscillator systems [5, 8].
The computation of Floquet multipliers and eigenvectors generally involves two key steps: discretization and solving the resulting eigenvalue problem. Existing high-precision software packages such as AUTO-07p [9] and MATCONT [10] adopt high-order one-step collocation methods for discretization, resulting in a periodic eigenvalue problem (periodic-EVP) [6], which can be solved by periodic QR algorithm [11, 12]. This framework makes them well-suited for the first scenario, where the parameterized periodic matrix function $G(t;\delta)$ is explicitly defined. However, the collocation method comes with several limitations for the second scenario, where the system matrix is only available at discrete time samples. It requires over-sampling at Gaussian collocation points within each subinterval, explicit construction of large system matrices, and a condensation step to eliminate over-sampled collocation points. This workflow is computationally expensive and memory-intensive for large-scale systems, and it is not well-suited for efficiently computing Floquet eigenvectors [6].
Motivated by these limitations, and aiming to address both scenarios, this paper applies a multistep method to discretize eq.˜1 that induces a polynomial periodic eigenvalue problem (periodic-PEP)
$\left(\theta^{d}+\sum_{j=0}^{d}\mathcal{A}_{j}\mathcal{C}^{d-j}\theta^{j}\right)u=0,$
where $\mathcal{A}_{j}\mathcal{C}^{d-j}$ are block cyclic matrices constructed from the discretization method. Compared to traditional collocation-based methods, our approach does not require an explicit expression of the system, avoids sampling at fixed Gaussian collocation points, eliminates the need for condensation, and does not explicitly construct or store large coefficient matrices. It naturally supports the simultaneous computation of Floquet multipliers and their eigenvectors, making it especially suitable for large-scale and data-driven systems.
The main contributions of this paper are as follows:
  1. 1.
    We apply $d$-step method to an LPTV system, corresponding to a periodic-PEP of degree $d$. We find that the $d$-step discretization generates d1d-1 spurious LPTV systems, whose designed $d$-step discretization corresponds to the same periodic-PEP.
  2. 2.
    We establish convergence analysis based on subspace perturbation theory, as the maximum stepsize $\Delta t$ goes to 0. Among $nd$ eigenpairs of the periodic-PEP, $n$ of them converge to Floquet eigenpairs at rate $\mathcal{O}(\Delta t^{s})$, where $s$ is the consistency order of the multistep method. The remaining n(d1)n(d-1) eigenvalues are related to spurious LPTV systems and decay to 0.
  3. 3.
    We propose an iterative algorithm pTOAR to solve the dominant eigenpairs of the periodic-PEP, and analyze its computational and memory costs. In particular, we show that using high-order multistep method incurs only negligible additional cost compared to the first-order one-step backward Euler method.
The remainder of this paper is organized as follows. In section˜2, the necessary background on Floquet theory, one-step collocation methods, multistep methods, and subspace perturbation theory are reviewed. In section˜3, we analyze how applying a multistep discretization to an LPTV system gives rise to a periodic-PEP. In section˜4, we introduce the spurious systems used to interpret the eigenpairs of the periodic-PEP and define the target objects to which they are expected to converge. In section˜5, the convergence results are established for the eigenpairs using tools from subspace perturbation theory. In section˜6, we present the proposed pTOAR algorithm for computing the dominant eigenpairs of large-scale periodic-PEP, together with an analysis of its computational and memory costs. Numerical experiments are reported in Section˜7, and conclusions are drawn in Section˜8.
2 Preliminaries
2.1 Floquet Theory via the State Transition Matrix of LPTV Systems
Our goal is to compute Floquet multipliers and eigenvectors, defined as the eigenpairs of the state transition matrix (STM) over one period. To this end, we first recall the definitions of these quantities and their key properties, which constitute the core of Floquet theory. The comprehensive Floquet theory can be found in [13, Chapter 5] and [14, Chapter 2].
A typical LPTV system with period $T$ is defined as eq.˜1. Let $X(t)$ denote a fundamental solution matrix, defined by $\dot{X}(t)=G(t)X(t)$, where the initial value matrix $X(0)=X_{0}$ is any full-rank matrix. The associated STM from $s$ to $t$ is $\Phi(t,s)=X(t)X(s)^{-1}$. Note that the definition of $\Phi(t,s)$ is independent of the particular choice of $X_{0}$.
The periodicity of the system matrix $G(t)$ implies that the STM $\Phi(t+T,t)$ is similar to $\Phi(T,0)$ for any $t$, as expressed in eq.˜2:
$\Phi(t+T,t)=\Phi(t,0)\,\Phi(T,0)\,\Phi^{-1}(t,0).$ (2)
To establish eq.˜2, we first note the periodic shift property of the STM,
$\Phi(t+T,T)=\Phi(t,0),\qquad\forall t\in\mathbb{R},$ (3)
which can be proven as follows. Define $Y(t)\equiv X(t+T)\,X^{-1}(T)\,X_{0},$ where $X(t)$ is the fundamental solution matrix. Then $\dot{Y}(t)=G(t)Y(t)$ and $Y(0)=X_{0}$. By the uniqueness of solutions to linear ODEs, we have $Y(t)\equiv X(t)$, which proves eq.˜3. Next, by decomposing $\Phi(t+T,0)$ at $T$ and $t$, we have
$\Phi(t+T,t)\Phi(t,0)=\Phi(t+T,T)\Phi(T,0).$
Right multiplying $\Phi(t,0)^{-1}$ on both sides and then using eq.˜3 yields eq.˜2.
The similarity relation eq.˜2 implies that all matrices $\Phi(t+T,t)$ share the same eigenvalue matrix $R$, whose diagonal entries are the Floquet multipliers. Moreover, there exists a $T$-periodic matrix $U(t)$ such that
$\Phi(t+T,t)=U(t)RU^{-1}(t),$ (4)
where the columns of $U(t)$ and the rows of $U^{-\mathrm{H}}(t)$ provide the corresponding right and left Floquet eigenvectors, respectively.
To construct $U(t)$ explicitly, we should use the the eigen-decomposition at $t=0$, that $\Phi(T,0)=U(0)RU^{-1}(0),$ and write $R=\exp(MT)$, where the diagonal entries of $M$ are the Floquet exponents. Now, we define $U(t)$ as:
$U(t)=\Phi(t,0)U(0)\exp(-Mt).$ (5)
In summary, Floquet multipliers and eigenvectors are obtained from the eigenvalue problem
$\Phi(T,0)u(0)=u(0)\rho,\quad\rho=e^{\mu T},$ (6)
at $t=0$, while the Floquet eigenvectors at any other time $t$ can be recovered using eq.˜5. Therefore, a single eigen-decomposition of $\Phi(T,0)$ suffices to determine all Floquet quantities.
2.2 One-step Methods for Solving Floquet Eigenpairs
Now, to numerically solve eq.˜6, we recall the existing one-step discretization of eq.˜1. The period $[0,T]$ is partitioned at time points $0=t_{0}<t_{1}<\cdots<t_{p}=T$, with stepsizes $\Delta t_{i}=t_{i}-t_{i-1}$. In general, we allow the stepsizes to be variable, and the maximum stepsize is $\Delta t$. Then, we write eq.˜6 on the time grid:
$\Phi(t_{p},t_{p-1})\cdots\Phi(t_{2},t_{1})\Phi(t_{1},t_{0})u(t_{0})=u(t_{0})e^{\mu T}.$
When $\Delta t$ is small enough, one-step methods approximate each $\Phi(t_{i},t_{i-1})$ by a matrix $A^{(i)}$ for $i=1,\ldots,p$, with $u_{i}$ denote the numerical solution at time $t_{i}$. This leads to a periodic eigenvalue problem to compute Floquet eigenpairs:
$A^{(p)}\cdots A^{(2)}A^{(1)}u_{0}=u_{0}\lambda.$ (7)
A well-known one-step method is backward Euler (BE), in which each transition matrix is given by $A^{(i)}=(I-G(t_{i})\Delta t_{i})^{-1}$ when applied to eq.˜1. This method only requires sampling of $G(t_{i})$ at time $t_{i}$. However, its first-order accuracy is usually insufficient for precise Floquet analysis.
High-order one-step methods, such as collocation, improve discretization accuracy. Existing public software packages for computing and analyzing limit cycles, such as MATCONT [10] and AUTO-07p [9], use collocation to discretize periodic systems, and then form the discrete local state transition matrices $A^{(i)}$ to compute Floquet multipliers. In what follows, we briefly outline this procedure; a comprehensive process can be found in [6].
Consider an interval $t_{i-1},t_{i}]$, we first choose the number of collocation points $k$, where a larger $k$ corresponds higher local approximation accuracy. Collocation methods introduce internal stage variables at $k$ equidistant nodes and constructs a Lagrange polynomial interpolant through these points to approximate the solution $x(t)$. The interpolant is then required to satisfy eq.˜1 at $k$ Gaussian collocation points within $t_{i-1},t_{i}]$, which needs sampling the system matrix $G(t)$ at these addtional Guassian collocation points, leading to oversampling.
The next step, known as condensation, is eliminating the intermediate stage variables to obtain the one-step relation $x_{i}=A^{(i)}x_{i-1}$, typically via an LU elimination [6]. This process yields an explicit construction of the local STM $A^{(i)}$.
Although collocation methods have high numerical accuracy, their limitations become apparent depending on the application scenario. In the first scenario, where $G(t)$ is analytically known, and its dimension is small, the overhead of collocation is manageable. However, in the second scenario, where $G(t)$ is only available on given time grids, the collocation methods become both impractical and costly:
  1. 1.
    they rely on temporal internal sampling on Gaussian points of each interval for accuracy, which is expensive and often infeasible in large-scale applications;
  2. 2.
    the condensation procedure incurs significant computational cost, often about twice that of a direct eigendecomposition [6];
  3. 3.
    the condensed transition matrix $A^{(i)}$ is typically dense and must be fo rmed explicitly, increasing memory usage and limiting the advantage of matrix-free solvers.
2.3 Multistep Methods
We now introduce multistep methods, a class of discretization schemes that attain high accuracy by reusing previously sampled values at adjacent grid points. We recall their general form and the standard definitions of consistency and stability for multistep methods [15, 16].
Definition 1 (Multistep Methods [15][Chapter 3.5] ).
For a fixed time interval $[0,T]$, we still consider a time grid $0=t_{0}<\cdots<t_{p}=T$. For an ODE $\dot{x}=f(t,x)$, a $d$-step method updates the solution at $t_{i}$ by combining the past $d$ steps as:
$\sum_{j=0}^{d}\alpha_{j,i}x_{i-d+j}=\Delta t_{i}\left(\sum_{j=0}^{d}\beta_{j,i}f(t_{i-d+j},x_{i-d+j})\right),$ (8)
where the coefficients $\alpha_{j,i}$ and $\beta_{j,i}$ depend on the past d1d-1 local stepsize ratios $\omega_{i},\omega_{i-1},\ldots,\omega_{i-d+1}$, with $\omega_{i}=\Delta t_{i}/\Delta t_{i-1}$. We normalize $\alpha_{d,i}=1$ and require $\left\lvert\alpha_{0,i}\right\rvert+\left\lvert\beta_{0,i}\right\rvert\neq 0$ to ensure all $d$ steps contribute.
For detailed derivations of these coefficients, see [15, Chapter 3]. As an illustration, the second-order BDF method (Gear2) has coefficients given by eq.˜9.
$\alpha_{0,i}=\frac{{\omega_{i}}^{2}}{1+2\omega_{i}},\,\alpha_{1,i}=-1-\alpha_{0,i},\alpha_{2,i}=1,\,\beta_{2,i}=\frac{1+\omega_{i}}{1+2\omega_{i}},\,\beta_{1,i}=\beta_{0,i}=0.$ (9)
We now turn to the fundamental notions of consistency and stability for such multistep methods.
Definition 2 (Consistency [15][Chapter 3, 5.17]).
A multistep method of the form eq.˜8 is said to be consistent of order $s$ if, when applied to a sufficiently smooth function $x(t)$, its local truncation error $R_{i}$ satisfies $\left\lVert R_{i}\right\rVert=\mathcal{O}(\Delta t_{i}^{s+1})$ as $\Delta t_{i}\to 0$, where
$R_{i}=\sum_{j=0}^{d}\alpha_{j,i}x(t_{i-d+j})-\Delta t_{i}\sum_{j=0}^{d}\beta_{j,i}\frac{\mathrm{d}x}{\mathrm{d}t}(t_{i-d+j}).$ (10)
The stability of a multistep method is to guarantee boundedness of the numerical solution. The trivial ODE $\dot{x}=0$ is usually chosen to be the test equation. Apply the multistep discretization, we obtain the homogeneous recurrence $\sum_{j=0}^{d}\alpha_{j,i}x_{i-d+j}=0.$ Now, we introduce $X_{i}=(x_{i-d+1},\cdots,x_{i-1},x_{i})^{\mathrm{T}}$ and the companion matrix
$\mathbf{F}_{i,i-1}=\begin{bmatrix}&1&&\\ &&\ddots&\\ &&&1\\ -\alpha_{0,i}&-\alpha_{1,i}&\cdots&-\alpha_{d-1,i}\end{bmatrix},\quad\mathbf{F}_{i+j,i}\equiv\mathbf{F}_{i+j,i+j-1}\cdots\mathbf{F}_{i+2,i+1}\mathbf{F}_{i+1,i}.$
Definition 3 (Stability [15][Chapter 3, Definition 5.4]).
The multistep method in eq.˜8 is called stable if for all $i$ and $j\geq 0$, there exists a constant $C$, such that
$\left\lVert\mathbf{F}_{i+j,i}\right\rVert\leq C.$ (11)
For a fixed multistep method, the matrices $\mathbf{F}_{i+j,i}$ in definition˜3 depend on the chosen time grid, because the coefficients $\alpha_{j,i}$ and $\beta_{j,i}$ are functions of the local stepsize ratios. As a result, stability can only be verified for each specific grid individually, which becomes cumbersome when the stepsizes vary. [15] shows that, if the method is stable under constant stepsizes and certain additional conditions are satisfied, there exist constants $0<\omega<1<\Omega$ such that, whenever the step ratios satisfy $\omega_{i}\in[\omega,\Omega]$, the multistep method remains stable on all corresponding grids. For simplicity, throughout this paper we assume that the method is stable on the chosen variable stepsize grid.
2.4 Subspace Perturbation Theory
In this subsection, we introduce the existing results about subspace perturbation theory, which is the main tool for analysis in this paper. We start with some notations: For any given $A\in\mathbf{C}^{n\times n}$ and $B\in\mathbf{C}^{m\times m}$, then:
  1. 1.
    Spectral gap: $\operatorname{d}(A,B)=\min\{|\lambda-\mu|:\lambda\in\lambda(A),\mu\in\lambda(B)\};$
  2. 2.
    Departure from normality of $A$ with respect to the matrix Frobenius norm:
[…]
  3. 3.
    The separation of $A$ and $B$ with respect to the matrix $\mathrm{F}$-norm:
    $\operatorname{sep}_{\mathrm{F}}(A,B)=\inf_{\left\lVert P\right\rVert=1}\left\lVert PA-BP\right\rVert_{\mathrm{F}}.$
Directly computing the separation is hard, but it can be bounded by the spectral gap and departure from normality. In theorem˜1, the upper bound is given in [17], and the lower bound is given by [18].
Theorem 1 (Bounds of $\operatorname{sep}(A,B)$ [18, 17]).
Let $A\in\mathbf{C}^{n\times n}$ and $B\in\mathbf{C}^{m\times m}$ with $\lambda(A)\cap\lambda(B)=\varnothing$. Then we have
$\operatorname{d}(A,B)\geqslant\operatorname{sep}_{\mathrm{F}}(A,B)\geqslant\frac{\operatorname{d}(A,B)}{1+\sum_{j=1}^{dn-2}\left[\frac{\Delta_{\mathrm{F}}(A)+\Delta_{\mathrm{F}}(B)}{\operatorname{d}(A,B)}\right]^{j}}.$
If $A$ and $B$ are both normal, then $\Delta(A,B)=0$, and we obtain $\operatorname{d}(A,B)=\operatorname{sep}_{\mathrm{F}}(A,B)$.
The second one is a perturbation theorem from [19, Theorem 3.5], and Betcke and Kressner gave a new proof in [20].
Theorem 2 (Invariant Subspace Perturbation Theorem [19]).
Let $A,E\in\mathbb{C}^{n\times n}$, $P=[P_{1},P_{2}]$ is a unitary matrix. $P_{1}\in\mathbb{C}^{n\times l},(1\leq l\leq n-1)$. Assume $\operatorname{span}\left\{P_{1}\right\}$ is the invariant subspace of $A$. Let $P^{\mathrm{H}}AP$ and $P^{\mathrm{H}}EP$ are consistently separated as
$P^{\mathrm{H}}AP=\begin{bmatrix}A_{11}&A_{12}\\ &A_{22}\end{bmatrix},\quad P^{\mathrm{H}}EP=\begin{bmatrix}E_{11}&E_{12}\\ E_{21}&E_{22}\end{bmatrix}$
If $\delta\equiv\operatorname{sep}(A_{11},A_{22})-\left(\left\lVert E_{11}\right\rVert+\left\lVert E_{22}\right\rVert\right)>0$, and $\left\lVert E_{21}\right\rVert\left(\left\lVert A_{12}\right\rVert+\left\lVert E_{12}\right\rVert\right)<\frac{1}{4}\delta^{2}$, then there exist a matrix $R\in\mathbb{C}^{(n-l)\times l}$ satisfying $\left\lVert R\right\rVert\leq 2\delta^{-1}\left\lVert E_{21}\right\rVert$, such that $\operatorname{span}\left\{P_{1}\begin{bmatrix}I\\ R\end{bmatrix}\right\}$ is an invariant subspace of $A+E$.
3 Periodic-PEP and Discrete State Transition Matrix
When a multistep method is applied to an LPTV system eq.˜1, it gives rise to a linear difference equation with periodic coefficient matrices, referred to as a periodic difference equation. In this section, we will show that this periodic difference equation corresponds to a periodic polynomial eigenvalue problem (periodic-PEP). Furthermore, the periodic-PEP can be transformed into a periodic eigenvalue problem whose coefficient matrix coincides with the discrete STM over a period, capturing the dynamics of the periodic difference equation.
3.1 Derivation of the Periodic-PEP
As in discussion of one-step method, we consider the period $[0,T]$ is partitioned at time points $0=t_{0}<t_{1}<\cdots<t_{p}=T$, In fact, this grid can be extended to an infinite sequence $\left\{t_{i}\right\}_{i\in\mathbb{Z}}$ satisfying $t_{i+p}=t_{i}+T$. This extension defines a periodic time grid. Throughout this paper, quantities that exhibit periodicity with period $p$ on this periodic discrete grid are indexed with a superscript $(i)$ to indicate their cyclic dependence. Hence, the coefficients $(\alpha_{j,i},\beta_{j,i})$ are denoted by $(\alpha_{j}^{(i)},\beta_{j}^{(i)})$, and the step size $\Delta t^{(i)}=t_{i}-t_{i-1}$, and the maximum step $\Delta t=\max_{i=1}^{p}\Delta t^{(i)}$.
$\sum_{j=0}^{d}\left(\alpha_{j}^{(i)}-\Delta t^{(i)}\beta_{j}^{(i)}G^{(i-d+j)}\right)x_{i-d+j}=0,$ (12)
Since $\left(I_{n}-\beta_{d}^{(i)}G^{(i)}\Delta t^{(i)}\right)$ is invertible for sufficiently small step size $\Delta t^{(i)}$, eq.˜12 can be expressed in the form of a $d$-step difference equation with $p$-periodic coefficients:
$x_{i}-\sum_{j=0}^{d-1}A_{j}^{(i)}x_{i-d+j}=0,\quad i=1,\ldots,p,$ (13)
where the coefficient matrices $A_{j}^{(i)}$ for $j=0,\ldots,d-1$ are defined by
$A_{j}^{(i)}=-\left(I_{n}-\beta_{d}^{(i)}G^{(i)}\Delta t^{(i)}\right)^{-1}\left(\alpha_{j}^{(i)}I_{n}-\beta_{j}^{(i)}G^{(i-d+j)}\Delta t^{(i)}\right).$ (14)
In this section, we only use the cyclicity of $A_{j}^{(i)}$ and do not require any of its specific properties from the multistep method.
Next, we define the periodic-PEP corresponding to eq.˜13.
Definition 4 (Periodic Polynomial Eigenvalue Problem).
Let $\{A_{0}^{(i)},A_{1}^{(i)},\ldots,A_{d-1}^{(i)}\}_{i=1}^{p}$ be $p$ sets of coefficient matrices in $\mathbb{R}^{n\times n}$. The periodic polynomial eigenvalue problem (periodic-PEP) seeks scalars $\theta\in\mathbb{C}$ and nonzero vectors $u^{(i)}\in\mathbb{C}^{n}$, $i=1,\ldots,p$, satisfying the periodicity condition $u^{(i+p)}=u^{(i)}$ such that
$u^{(i)}\theta^{d}-\sum_{j=0}^{d-1}A_{j}^{(i)}\,u^{(i-d+j)}\theta^{j}=0,\qquad i=1,\ldots,p.$ (15)
If $(\theta,\{u^{(i)}\}_{i=1}^{p})$ is an eigenpair, then for each $k=0,\ldots,p-1$, $\left(\theta\zeta_{p}^{k},\,\{u^{(i)}\zeta_{p}^{ik-k}\}_{i=1}^{p}\right)$ is also an eigenpair, where $\zeta_{p}=e^{\jmath 2\pi/p}$ is a primitive $p$th root of unity. Among them, we define the periodic eigenvalue of the periodic-PEP as $\lambda=\theta^{p}$, and the periodic eigenvector as the sequence $\{u^{(i)}\}_{i=1}^{p}$ associated with the root $\theta$ whose argument $\arg(\theta)\in[0,2\pi/p)$ is minimal.
For $d=1$, the periodic-PEP reduces to the standard periodic eigenvalue problem discussed in [11]. For $d>1$, the problem takes a polynomial form, which we now illustrate. As an example, when $d=2$, and $p=3$, definition˜4 can be written as:
$\left(\theta^{2}I_{3n}-\begin{bmatrix}&&A_{1}^{(1)}\\ A_{1}^{(2)}&&\\ &A_{1}^{(3)}&\\ \end{bmatrix}\theta-\begin{bmatrix}&A_{0}^{(1)}&\\ &&A_{0}^{(2)}\\ A_{0}^{(3)}&&\\ \end{bmatrix}\right)\begin{bmatrix}u^{(1)}\\ u^{(2)}\\ u^{(3)}\\ \end{bmatrix}=0.$ (16)
This example shows that stacking the $p$ component equations in definition˜4 naturally yields a PEP whose coefficient matrices possess a block-cyclic structure. To express this structure compactly, we introduce the stacking notation
$\operatorname{col}(x_{i})_{i=m+1}^{m+l}=\left[x_{m+1}^{\mathrm{H}},\,x_{m+2}^{\mathrm{H}},\,\ldots,\,x_{m+l}^{\mathrm{H}}\right]^{\mathrm{H}}\in\mathbb{R}^{ln},$ (17)
which collects lm+1l-m+1 consecutive state vectors ending at step $i$. We then define the block-diagonal matrices $\mathcal{A}_{j}=\operatorname{diag}_{i}(A_{j}^{(i)}),\,j=0,\ldots,d-1,$ and the cyclic shift operator $\mathcal{C}=C_{p}\otimes I_{n}$, $C_{p}=[e_{2},\dots,e_{p},e_{1}],$ where $e_{i}$ denotes the $i$-th standard basis vector of $\mathbb{R}^{p}$.
With these notations, the periodic-PEP in definition˜4 can be equivalently written in the compact matrix form.
$\left(\theta^{d}I_{np}-\sum_{j=0}^{d-1}\mathcal{A}_{j}\mathcal{C}^{d-j}\theta^{j}\right)\operatorname{col}(u^{(i)})_{i=1}^{p}=0.$ (18)
3.2 Equivalent Form of the Periodic-PEP using Discrete State Transition Matrix
This subsection transforms the periodic-PEP in eq.˜18 into an equivalent form to reveal its connection with the general solution of the underlying difference equation. We show that through a suitable linearization and a structured reordering of the companion-form matrices, eq.˜18 becomes an $nd$-dimensional periodic eigenvalue problem. The coefficient matrices of this new problem are identical to the state transition matrices of the difference equation eq.˜13. This identity establishes a direct correspondence between the eigenpairs of the periodic-PEP and the fundamental solutions of the periodic difference system.
The companion form of the PEP eq.˜18 is:
$\begin{bmatrix}&I_{np}&&\\ &&\ddots&\\ &&&I_{np}\\ \mathcal{A}_{0}\mathcal{C}^{d}&\mathcal{A}_{1}\mathcal{C}^{d-1}&\cdots&\mathcal{A}_{d-1}\mathcal{C}\\ \end{bmatrix}\begin{bmatrix}\operatorname{col}(u^{(i)})_{i=1}^{p}\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta\\ \vdots\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta^{d-1}\end{bmatrix}=\begin{bmatrix}\operatorname{col}(u^{(i)})_{i=1}^{p}\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta\\ \vdots\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta^{d-1}\end{bmatrix}\theta,$ (19)
The eigenvector in eq.˜19 exhibits a two-level structure: the inner structure encodes the periodicity of the coefficients, while the outer structure arises from the linearization of eq.˜18. To make this structure explicit, we introduce a permutation matrix $\mathcal{P}$ that reorders the stacked vector:
$\mathcal{P}\begin{bmatrix}\operatorname{col}(u^{(i)})_{i=1}^{p}\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta\\ \vdots\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta^{d-1}\end{bmatrix}=\begin{bmatrix}\operatorname{col}(u^{(1-j)}\theta^{d-j})_{j=1}^{d}\\ \operatorname{col}(u^{(2-j)}\theta^{d-j})_{j=1}^{d}\\ \vdots\\ \operatorname{col}(u^{(p-j)}\theta^{d-j})_{j=1}^{d}\\ \end{bmatrix},\,\operatorname{col}(u^{(i-j)}\theta^{d-j})_{j=1}^{d}=\begin{bmatrix}u^{(i-d+1)}\\ u^{(i-d+2)}\theta\\ \vdots\\ u^{(i)}\theta^{d-1}\\ \end{bmatrix}$.
After this reordering, the eigenvalue problem takes the following block-cyclic form:
$\begin{bmatrix}&&&L^{(1)}\\ L^{(2)}&&&\\ &\ddots&&\\ &&L^{(p)}&\\ \end{bmatrix}\begin{bmatrix}\operatorname{col}(u^{(1-j)}\theta^{d-j})_{j=1}^{d}\\ \operatorname{col}(u^{(2-j)}\theta^{d-j})_{j=1}^{d}\\ \vdots\\ \operatorname{col}(u^{(p-j)}\theta^{d-j})_{j=1}^{d}\\ \end{bmatrix}=\begin{bmatrix}\operatorname{col}(u^{(1-j)}\theta^{d-j})_{j=1}^{d}\\ \operatorname{col}(u^{(2-j)}\theta^{d-j})_{j=1}^{d}\\ \vdots\\ \operatorname{col}(u^{(p-j)}\theta^{d-j})_{j=1}^{d}\\ \end{bmatrix}\theta,$ (20)
where each block $L^{(i)}$ is the companion matrix
$L^{(i)}=\begin{bmatrix}&I_{n}&&\\ &&\ddots&\\ &&&I_{n}\\ A_{0}^{(i)}&A_{1}^{(i)}&\cdots&A_{d-1}^{(i)}\\ \end{bmatrix}$. (21)
The block-cyclic structure implies that, for every $i=1,\ldots,p$, the eigenvalue problem reduces to the $nd$-dimensional periodic eigenvalue problem:
$L^{(i)}L^{(i-1)}\cdots L^{(i-p+1)}\operatorname{col}(u^{(i-j)}\theta^{d-j})_{j=1}^{d}=\operatorname{col}(u^{(i-j)}\theta^{d-j})_{j=1}^{d}\theta^{p}.$ (22)
This transformation maps each $p$-fold set of eigenvalues of definition˜4 to a single eigenvalue $\theta^{p}$, effectively removing the rotational symmetry and revealing the essential dynamics.
Recall eq.˜12, we observe that each companion matrix $L^{(i)}$ describes one-step propagation of the stacked state vectors:
$\operatorname{col}(x_{j})_{j=i-d+1}^{i}=L^{(i)}\operatorname{col}(x_{j})_{j=i-d}^{i-1}.$ (23)
To describe the evolution, we define the discrete STM in a $dn$-dimensional state space [21]:
$F_{i+j,i}=L^{(i+j)}L^{(i+j-1)}\cdots L^{(i+1)}.$ (24)
Because the coefficients $L^{(i)}$ are $p$-periodic, we have $F_{i+kp,i+(k-1)p}=F_{i+p,i}$ for all integers $k$. Therefore, we partition the stacked sequence $\operatorname{col}(x_{j})_{j=i-d+1}^{i}$ for all $i\in\mathbb{Z}$ into $p$ groups according to the remainder of the index $i$ modulo $p$. Each group forms a phase slice $\operatorname{col}(x_{j})_{j=i+kp-d+1}^{i+kp}$, and within each slice the evolution satisfies a constant-coefficient difference equation:
$\operatorname{col}(x_{j})_{j=i+kp-d+1}^{i+kp}=F_{i+p,i}\,\operatorname{col}(x_{j})_{j=i+(k-1)p-d+1}^{i+(k-1)p}.$ (25)
Here, $F_{i+p,i}$ acts as the discrete state-transition matrix over one period for the corresponding slice, and the associated eigenvalue problem is precisely eq.˜22.
Moreover, these phase slices are not independent, because the transition matrices for different starting points are all similar:
$F_{i+p,i}=F_{i,0}F_{p,0}F_{i,0}^{-1},$ (26)
and therefore share the same spectrum. As a result, the eigen-decomposition of $F_{p,0}$ and state transition $F_{i,0}$ fully determines the eigen-structure of all $F_{i+p,i}$.
In this section, we have shown the discrete STM over a period $F_{i+p,i}$ reveals structural properties that closely mirror those of the continuous STM $\Phi(T,0)$ discussed in section˜2.1. The key distinction is that the discrete STM acts on an $nd$-dimensional stacked state corresponding to the difference equation, whereas the continuous STM evolves in the original $n$-dimensional state space of the LPTV system.
When the coefficient matrix $A_{j}^{(i)}$ are from multistep discretization of the LPTV system, we expect $n$ of the eigenvalues of $F_{p,0}$ to approximate the $n$ Floquet multipliers of the continuous system. However, the remaining n(d1)n(d-1) eigenvalues of $F_{p,0}$ require further explanation, which will be addressed by introducing spurious systems in the following section.
4 Spurious Systems from Multistep Methods
The continuous LPTV system eq.˜1 evolves in an $n$-dimensional state space. Discretizing it with a $d$-step multistep method yields a difference method whose state space is $nd$-dimensional and one-period STM is $F_{p,0}$. In this section, based on the given LPTV system and the adopted $d$-step multistep method, we design d1d-1 additional $n$-dimensional spurious systems together with their corresponding multistep discretizations. The discretizations of the physical and spurious systems share identical coefficient matrices, hence yield the same difference equation in coefficient form, but differ in the way the state variables enter the recursion. Utilizing the Floquet multipliers and eigenvectors of these spurious systems, we construct an $nd$-dimensional auxiliary STM $\Psi(t_{p},t_{0})$ that serves as the limiting counterpart of $F_{p,0}$ as $\Delta t\to 0$.
4.1 Spurious Systems from Discretizing Scalar System
To illustrate how multistep methods introduce the spurious systems, we first analyze a simple scalar system, $\dot{y}=0$, $t\in[0,T]$. Applying the $d$-step discretization on a periodic time grid leads to a high-order difference equation:
$y_{i}=-\sum_{j=0}^{d-1}\alpha_{j}^{(i)}y_{i-d+j},\quad i=1,\ldots,p.$ (27)
It corresponds to the following periodic-PEP:
$\left(\nu^{d}I_{p}-\sum_{j=0}^{d-1}\mathbf{A}_{j}C^{d-j}\nu^{j}\right)\eta=0,$ (28)
where $\mathbf{A}_{j}=-\operatorname{diag}\left\{\alpha_{j}^{(i)}\right\}_{i=1}^{p}$, and $\eta=[\eta^{(1)},\ldots,\eta^{(p)}]^{\mathrm{H}}$. According to definition˜4, the $d$ periodic eigenvalues of eq.˜28 are denoted by $\nu_{1}^{p},\ldots,\nu_{d}^{p}$. According to definition˜3, the trivial eigenvalue $\nu_{1}=1$ remains unchanged, while all other eigenvalues satisfy $\lvert\nu_{\tau}\rvert<1$ for $\tau=2,\ldots,d$. The corresponding eigenvector of $\nu_{1}$ is still $1,1,\ldots,1]^{\mathrm{H}}$, whereas the other eigenvectors are no longer uniform when the stepsize is not constant.
4.2 Spurious Systems from Discretizing LPTV Systems
Returning now to the general system eq.˜13, each coefficient matrix in eq.˜14 is a small perturbation of $\mathbf{A}_{j}\otimes I_{n}$ when $\Delta t$ is sufficiently small. Motivated by the scalar case, we assume that the general solution of eq.˜13 inherits the same structure, taking the form $x_{\tau,i}\eta^{(i)}\nu^{i}$. Substituting this back to eq.˜13 yields:
$x_{\tau,i}\eta_{\tau}^{(i)}\nu_{\tau}^{i}-\sum_{j=1}^{d}A_{j}^{(i)}x_{\tau,i-d+j}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{i-d+j}=0,\quad\tau=1,\ldots,d.$ (29)
The mode $\nu_{1}=1$ with $\eta_{1}^{(i)}\equiv 1$ corresponds to the physical dynamics and recovers eq.˜13. The remaining modes $\nu_{\tau}$ $(\tau>1)$ also describe valid dynamical components through $x_{\tau,i}$; however, their associated decaying envelopes render them spurious in the context of the original system.
To formalize how $x_{\tau,i}$ captures these spurious dynamics, we associate each mode $\tau$ with a corresponding spurious LPTV system and a spurious multistep method. In doing so, we show that eq.˜29 can be obtained from the discretization of these spurious systems. This motivates the construction of the corresponding spurious multistep methods, whose coefficients are then defined as
$\alpha_{\tau,j}^{(i)}=\frac{\alpha_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}}{\sum_{j=0}^{d}\alpha_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}h_{j}^{(i)}},\quad\beta_{\tau,j}=\frac{\beta_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}}{\sum_{j=0}^{d}\beta_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}}.$ (30)
where the envelope terms are incorporated into the numerators, and the normalization ensures the first-order consistency condition (definition˜2). The scaling factor $h_{j}^{(i)}$ in the denominator of $\alpha_{\tau,j}^{(i)}$ is defined by
$h_{j}^{(i)}\equiv\frac{t_{i+j}-t_{i}}{t_{i+d}-t_{i+d-1}}=\frac{\Delta t^{(i+1)}}{\Delta t^{(i+d)}}\sum_{l=1}^{j}\frac{\Delta t^{(i+l)}}{\Delta t^{(i+1)}}=\left(\prod_{l=1}^{d}{\omega^{(i+l)}}^{-1}\right)\sum_{l=1}^{j}\omega^{(i+l)}\cdots\omega^{(i+1)},$ (31)
where $\omega^{(i)}=\Delta t^{(i)}/\Delta t^{(i-1)}$ denotes the local stepsize ratio. As seen from eq.˜31, $h_{j}^{(i)}$ depends only on the stepsize ratios $\omega^{(i+l)}$ for $l=1,\ldots,d$, and reduces to $h_{j}^{(i)}=j$ under constant stepsizes.
We now construct the corresponding spurious LPTV systems:
$\frac{\mathrm{d}x_{\tau}}{\mathrm{d}t}$ $=\rho_{\tau}(t)G(t)x_{\tau}(t),$ (32)
$\rho_{\tau}(t)$ $=\frac{\sum_{j=0}^{d}\beta_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}}{\sum_{j=0}^{d-1}\alpha_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}h_{j}^{(i)}},\quad t\in[t_{i-1},t_{i}],\,\tau=2,\ldots,d.$
Although the underlying system is locally defined on each interval $t_{i-1},t_{i}]$ with potentially distinct dynamics, the $d$-step multistep discretization requires evaluating the equation over the combined interval $t_{i-d},t_{i}]$. To this end, we formally extend the equation form from the last interval $t_{i-1},t_{i}]$ to the preceding intervals $t_{i-d},t_{i-1}]$, effectively using the definition on $t_{i-1},t_{i}]$ to cover the earlier steps. Applying the spurious multistep coefficients from eq.˜30 to this extended system then yields
$\sum_{j=0}^{d}\left(\alpha_{j}^{(i)}I_{n}-\Delta t^{(i)}\beta_{j}^{(i)}G^{(i-d+j)}\right)x_{\tau,i-d+j}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}=0,\quad\tau=1,\ldots,d.$ (33)
Left-multiplying by $\left(I_{n}-\Delta t^{(i)}\beta_{d}^{(i)}G^{(i)}\right)^{-1}$ then recovers eq.˜29, and we can write eq.˜23 in an extended version
$L^{(i)}\operatorname{col}\left(x_{\tau,j}\,\eta^{(j)}\,\nu^{\,j}\right)_{j=i-d}^{i-1}=\operatorname{col}\left(x_{\tau,j}\,\eta^{(j)}\,\nu^{\,j}\right)_{j=i-d+1}^{i}.$ (34)
At last, we estimate the local truncation error of the spurious multistep method. Accroding to eq.˜10, we have the following lemma.
Lemma 3.
Consider the spurious $d$-step discretization eq.˜33 applied to the spurious system. When the stepsize is constant, the local truncation error of eq.˜33 is second-order, but its local truncation error drops to first order when stepsize is variable.
4.3 Construction of Auxiliary State Transition Matrix
Similar to the structure of $F_{i+1,i}=L^{(i)}$, we now consider a continuous representation. Let $X_{\tau}(t)$ denote a fundamental solution matrix of eq.˜32 for any fixed full rank initial matrix condition, and we define an $n$-dimensional subspace of an $nd$-dimensional space spanned by:
$\mathcal{X}_{\tau}\equiv\operatorname{span}\left\{\operatorname{col}\left(X_{\tau}^{(i)}(t_{j})\eta_{\tau}^{(j)}\nu_{\tau}^{j}\right)_{j=i-d+1}^{i}\right\}.$
Next, we show that the subspaces $\mathcal{X}_{1},\ldots,\mathcal{X}_{d}$ form a direct sum spanning the entire $nd$-dimensional space, that is $\mathcal{X}_{1}\oplus\mathcal{X}_{2}\oplus\cdots\oplus\mathcal{X}_{d}=\mathbb{R}^{(nd)\times(nd)}.$ To see this, we collect the bases of all subspaces into the block matrix
$\begin{bmatrix}X_{1}(t_{i-d+1})&X_{2}(t_{i-d+1})\eta_{2}^{(i-d+1)}&\cdots&X_{d}(t_{i-d+1})\eta_{d}^{(i-d+1)}\\ \vdots&\vdots&\ddots&\vdots\\ X_{1}(t_{i-1})&X_{2}(t_{i-1})\eta_{2}^{(i-1)}\nu_{2}^{i-1}&\cdots&X_{d}(t_{i-1})\eta_{d}^{(i-1)}\nu_{d}^{i-1}\\ X_{1}(t_{i})&X_{2}(t_{i})\eta_{2}^{(i)}\nu_{2}^{i}&\cdots&X_{d}(t_{i})\eta_{d}^{(i)}\nu_{d}^{i}\\ \end{bmatrix}$. (35)
Assume that $X_{\tau}(t_{i})$ is nonsingular for $\tau=1,\ldots,d$. For a fixed $\tau$ and sufficiently small $\Delta t$, the matrices $X_{\tau}(t_{j})$ with $j=i-d+1,\ldots,i$ differ from $X_{\tau}(t_{i})$ only by $\mathcal{O}(\Delta t)$, and the scaling factor $\eta^{(j)}$ vary smoothly as well. Consequently, the assembled block matrix can be regarded as a small perturbation of a block Vander monde matrix, which remains nonsingular provided that $\nu_{\tau}$ are pairwise distinct.
We define an auxiliary STM $\Psi(t_{i},t_{i-1})$ that satisfies
$\operatorname{col}\left(x_{\tau}^{\mathrm{H}}(t_{j})\,\eta_{\tau}^{(j)}\,\nu_{\tau}^{\,j}\right)_{j=i-d+1}^{i}=\Psi(t_{i},t_{i-1})\operatorname{col}\left(x_{\tau}^{\mathrm{H}}(t_{j})\,\eta_{\tau}^{(j)}\,\nu_{\tau}^{\,j}\right)_{j=i-d}^{i-1}.$ (36)
The block companion form of $\Psi(t_{i},t_{i-1})$ on subspace $\mathcal{X}_{\tau}$ is given by
$\left.\Psi(t_{i},t_{i-1})\right|_{\mathcal{X}_{\tau}}=\begin{bmatrix}0&I_{n}&&\\ &\ddots&\ddots&\\ &&0&I_{n}\\ 0&\cdots&0&\frac{\eta_{\tau}^{(i)}}{\eta_{\tau}^{(i-1)}}\nu_{\tau}\Phi_{\tau}(t_{i},t_{i-1})\\ \end{bmatrix}$. (37)
Repeated transition over one full period yields
$\left.\Psi(t_{i+p},t_{i})\right|_{\mathcal{X}_{\tau}}=$ $\begin{bmatrix}0&\cdots&0&\frac{\eta_{\tau}^{(i-d+1)}}{\eta_{\tau}^{(i)}}\Phi_{\tau}(t_{i+p-d+1},t_{i})\nu_{\tau}^{1-d}\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&0&\frac{\eta_{\tau}^{(i-1)}}{\eta_{\tau}^{(i)}}\Phi_{\tau}(t_{i+p-1},t_{i})\nu_{\tau}^{-1}\\ 0&\cdots&0&\Phi_{\tau}(t_{i+p},t_{i})\\ \end{bmatrix}$. (38)
Let $U_{\tau}(t_{i})$ be the eigenvector matrix of the STM of the $\tau$-th system $\Phi_{\tau}(t_{i+p},t_{i})$ associated with the Floquet multipliers $e^{M_{\tau}T}$ satisfying $\Phi_{\tau}(t_{i+p},t_{i})U_{\tau}(t_{i})=U_{\tau}(t_{i})e^{M_{\tau}T}.$ Denote the stacked eigenvector matrix:
$\mathcal{U}_{\tau}(t_{i})=\operatorname{col}\left(\Phi_{\tau}(t_{j},t_{i})U_{\tau}(t_{i})\eta_{\tau}^{(j)}\nu_{\tau}^{j}\right)_{j=i-d+1}^{i}.$ (39)
Then, it follows that $\mathcal{U}_{\tau}(t_{i+p})=\mathcal{U_{\tau}}(t_{i})\nu_{\tau}^{p}e^{M_{\tau}T}$, and
$\Psi(t_{i+p},t_{i})\mathcal{U_{\tau}}(t_{i})=\mathcal{U_{\tau}}(t_{i})\nu_{\tau}^{p}e^{M_{\tau}T}.$ (40)
The case $\tau=1$ corresponds to $\nu_{\tau}=1$, which corresponds to the physical system, whose eigenvalues $e^{M_{1}T}$ are the Floquet multipliers of the original LPTV system eq.˜1. The remaining $\tau=2,\ldots,d$ correspond to the spurious systems, whose eigenvalues $\nu_{\tau}^{p}e^{M_{\tau}T}$ are scaled versions of the Floquet multipliers of the spurious LPTV systems eq.˜32, and thus decay geometrically with increasing $p$ since $\lvert\nu_{\tau}\rvert<1$.
5 Spectral Convergence Analysis of Periodic-PEP from Multistep Discretization
To establish the convergence behavior of the discrete STM $F_{p,0}$ towards $\Psi(t_{p},t_{0})$, we analyze the difference $\Psi(t_{p},t_{0})-F_{p,0}$ acting on $\mathcal{U}_{\tau}$ for physical system $\tau=1$ and spurious systems $\tau=2,\ldots,d$. Since all spurious subspaces behave similarly, it suffices to present the proof for the representative two-step case ($d=2$). The proof proceeds in three steps:
  1. 1.
    Estimate the upper bound of $\left\lVert F_{i+l,i}\right\rVert$.
  2. 2.
    Estimate $\left\lVert[\Psi(t_{p},t_{0})-F_{p,0}][\mathcal{U}_{1},\mathcal{U}_{2}]\right\rVert$ by decomposing it into a sum of local differences across steps and bound each term.
  3. 3.
    Express the difference $\left\lVert[\Psi(t_{p},t_{0})-F_{p,0}][\mathcal{U}_{1},\mathcal{U}_{2}]\right\rVert$ in an orthonormal basis aligned with $\Psi(t_{p},t_{0})$’s invariant subspaces and apply the subspace perturbation theorem to conclude convergence of the physical and spurious subspaces.
Before we start the convergence analysis, we first introduce the following assumptions for the LPTV systems, periodic time grid and the multistep methods.
  1. A1.
    Consistency and stability of the multistep method: The $d$-step multistep method is of consistency order $s$ and zero-stable. Its coefficients $\alpha_{j}^{(i)}$ and $\beta_{j}^{(i)}$ are sufficiently smooth functions of the step-size ratios $\omega^{(i)}$.
  2. A2.
    Bounded variable step sizes: The stepsize ratios $\omega^{(i)}$ of the periodic time grid satisfy uniform bounds, i.e., $\omega<\omega^{(i)}<\Omega$, for some constants $0<\omega<1<\Omega$. Consequently, the coefficients $\alpha_{j}^{(i)}$ and $\beta_{j}^{(i)}$ admit uniform upper and lower bounds
    $\alpha_{\min}\leq\alpha_{j}^{(i)}\leq\alpha_{\max},\qquad\beta_{\min}\leq\beta_{j}^{(i)}\leq\beta_{\max},$
    for all $j=0,1,\ldots,d$ and $i=1,\ldots,p$.
  3. A3.
    Smoothness and boundedness of the LPTV system: The periodic matrix $G(t)$ in eq.˜1 is $s$-times continuously differentiable on $[0,T]$. In particular, smoothness implies that both $G(t)$ and the state-transition matrix $\Phi(t,s)$ are uniformly bounded:
    $\|G(t)\|_{\mathrm{F}}\leq C_{G},\qquad\|\Phi(t,s)\|\leq C_{\Phi},\quad\forall\,t-s\in[0,T].$
We begin by establishing the boundedness of the discrete STM $F_{i+p,i}$.
Lemma 4 (Stability and Boundedness of Discrete STM).
Under Assumption A1-A3, let $F_{i+l,i}$ denote the discrete STM over $l$ steps for the multistep discretization of eq.˜1. Then there exists a constant $C_{F}>0$, independent of $l$, such that
$\|F_{i+l,i}\|_{\mathrm{F}}\leq C_{F},\qquad l=1,\ldots,p.$
Proof.
Let $\mathbf{F}_{i+l,i}$ denote the discrete STM associated with the zero scalar system $\dot{y}=0$. By definition˜3, it satisfies $\|\mathbf{F}_{i+l,i}\|_{\mathrm{F}}\leq C_{0}$ for all $l\in\mathbb{N}$. Strictly speaking, the discrete STM acting on the $nd$-dimensional stacked state is $\mathbf{F}_{i+l,i}\otimes I_{n}$; however, for notational simplicity, we continue to denote it by $\mathbf{F}_{i+l,i}$.
We now regard $F_{i+l,i}$ as a perturbation of $\mathbf{F}_{i+l,i}$. For each step,
$\left\lVert F_{i+1,i}-\mathbf{F}_{i+1,i}\right\rVert_{\mathrm{F}}$ $\leq\sum_{j=0}^{d-1}\left\lVert\Delta A_{j}^{(i)}\right\rVert_{\mathrm{F}},$
where $\Delta A_{j}^{(i)}$ for $j=0,\ldots,d-1$ are defined as
$\Delta A_{j}^{(i)}$ $\equiv\left(\alpha_{d}^{(i)}I_{n}-\Delta t^{(i)}\beta_{d}^{(i)}G^{(i)}\right)^{-1}\left(\alpha_{j}^{(i)}I_{n}-\Delta t^{(i)}\beta_{j}^{(i)}G^{(i-d+j)}\right)-\alpha_{d}^{-1}\alpha_{j}I_{n}$
$=\Delta t^{(i)}\left(\alpha_{d}^{(i)}I_{n}-\Delta t^{(i)}\beta_{d}^{(i)}G^{(i)}\right)^{-1}{\alpha_{d}^{(i)}}^{-1}\left(\beta_{d}^{(i)}\alpha_{j}^{(i)}G^{(i)}-\alpha_{d}^{(i)}\beta_{j}^{(i)}G^{(i+j)}\right).$
For stepsize $\Delta t^{(i)}<\frac{\left\lvert\alpha_{d}^{(i)}\right\rvert}{2\left\lvert\beta_{d}^{(i)}\right\rvert C_{G}}\leq\frac{\alpha_{\max}}{2\beta_{\min}C_{G}}$, we have
$\left\lVert\Delta A_{j}^{(i)}\right\rVert_{\mathrm{F}}$ $<\left\lvert\alpha_{d}^{(i)}\right\rvert^{-1}\frac{2\alpha^{(i)}\beta^{(i)}C_{G}\Delta t^{(i)}}{\left\lvert\alpha_{d}^{(i)}\right\rvert-\Delta t\left\lvert\beta_{d}^{(i)}\right\rvert C_{G}}\leq 4\alpha_{\min}^{-2}\alpha_{\max}\beta_{\max}C_{G}\Delta t^{(i)},$
Hence, the bound on $\|\Delta A_{j}^{(i)}\|_{\mathrm{F}}$ holds uniformly for all $i$, leading to $\|F_{i+1,i}-\mathbf{F}_{i+1,i}\|_{\mathrm{F}}\leq C_{1}\,\Delta t^{(i)},$ for some constant $C_{1}$ independent of $i$.
The cumulative deviation over $l$ steps can be written as
$F_{i+l,i}-\mathbf{F}_{i+l,i}=\sum_{m=1}^{l}\mathbf{F}_{i+l,i+m}\,\big(F_{i+m,i+m-1}-\mathbf{F}_{i+m,i+m-1}\big)\,F_{i+m-1,i}.$
Taking norms and applying the triangle inequality yields
$\|F_{i+l,i}\|_{\mathrm{F}}$ $\leq\|\mathbf{F}_{i+l,i}\|_{\mathrm{F}}+\|\mathbf{F}_{i+l,i+m}\|_{\mathrm{F}}C_{1}\sum_{m=1}^{l}\Delta t^{(i+m-1)}\|F_{i+m-1,i}\|_{\mathrm{F}}$
$\leq C_{0}+C_{0}C_{1}\sum_{m=1}^{l}\Delta t^{(i+m-1)}\|F_{i+m-1,i}\|_{\mathrm{F}}.$
Finally, using the discrete Gronwall–Bellman inequality [22], we obtain
$\|F_{i+l,i}\|_{\mathrm{F}}\leq C_{0}e^{C_{0}C_{1}T}\equiv C_{F}.$
Thus the discrete STM remains uniformly bounded within each period. ∎
Lemma 5.
Under assumptions A1-A3, let $\Psi(t_{p},t_{0})$ and $F_{p,0}$ denote the auxiliary and discrete STM over $t_{0},t_{p}]$. Consider the difference $\Psi(t_{p},t_{0})-F_{p,0}$ acting on the normalized eigenvector matrices $\mathcal{U}_{\cdot}(t_{0})$ of $\Psi(t_{p},t_{0})$, with $\lVert\mathcal{U}_{\cdot}(t_{0})\rVert_{\mathrm{F}}=1$. Then the following bounds hold:
  1. 1.
    $\left\lVert\left[\Psi(t_{p},t_{0})-F_{p,0}\right]\mathcal{U}_{1}(t_{0})\right\rVert_{\mathrm{F}}\leq C_{1}\Delta t^{s}$,
  2. 2.
    $\left\lVert\left[\Psi(t_{p},t_{0})-F_{p,0}\right]\mathcal{U}_{2}(t_{0})\right\rVert_{\mathrm{F}}\leq C_{2}\Delta t$.
with constants $C_{1}$, $C_{2}$ independent of $\Delta t$.
Proof.
We decompose the global difference into sums of local difference:
$\left[\Psi(t_{p},t_{0})-F_{p,0}\right]\mathcal{U}_{\tau}(t_{0})$ (41)
$=$ $\sum_{i=1}^{p}F_{p,i}\cdot\left[\Psi(t_{i},t_{i-1})-F_{i,i-1}\right]\cdot\Psi_{i-1,0}\cdot\mathcal{U}_{\tau}(t_{0})$
$=$ $\sum_{i=1}^{p}F_{p,i}\cdot\left[\Psi(t_{i},t_{i-1})-F_{i,i-1}\right]\cdot\operatorname{col}\left(\Phi_{\tau}(t_{j},t_{0})U_{\tau}(t_{0})\eta_{\tau}^{(j)}\nu_{\tau}^{j}\right)_{j=i-d+1}^{i}.$
From eq.˜33 and lemma˜3, each local difference acts only on the last block of the column, giving
$\left[\Psi(t_{i},t_{i-1})-F_{i,i-1}\right]\operatorname{col}\left(\Phi_{\tau}(t_{j},t_{i})U_{\tau}(t_{i})\eta_{\tau}^{(j)}\nu_{\tau}^{j}\right)_{j=i-d+1}^{i}=\left[0,E_{i}\right],$
$\left\lVert E_{1,i}\right\rVert=\left\lVert\Phi_{1}(t_{i},t_{0})U_{1}(t_{0})\right\rVert\mathcal{O}(\Delta t^{s+1}),$
$\left\lVert E_{\tau,i}\right\rVert=\nu_{\tau}^{i}\left\lVert\Phi_{\tau}(t_{i},t_{0})U_{\tau}(t_{0})\right\rVert\mathcal{O}(\Delta t).$
For the physical mode ($\tau=1$), applying the triangle inequality and the uniform boundedness of $F_{p,i}$ and $\Phi_{1}(t_{i},t_{0})U_{1}(t_{0})$, we obtain
$\big\|\big[\Psi(t_{p},t_{0})-F_{p,0}\big]\mathcal{U}_{1}(t_{0})\big\|_{\mathrm{F}}$ $\leq\sum_{i=1}^{p}\|F_{p,i}\|_{\mathrm{F}}\,\|E_{1,i}\|\leq\sum_{i=1}^{p}C^{(i)}\,\Delta t^{s+1}\leq C_{1}\,\Delta t^{s},$
where $C_{1}$ is independent of $\Delta t$. This shows that the global error for the physical mode is $\mathcal{O}(\Delta t^{s})$.
For the spurious mode ($\tau=2$), we similarly have
$\big\|\big[\Psi(t_{p},t_{0})-F_{p,0}\big]\mathcal{U}_{2}(t_{0})\big\|_{\mathrm{F}}$ $\leq\sum_{i=1}^{p}\|F_{p,i}\|_{\mathrm{F}}\,\|E_{2,i}\|\leq\sum_{i=1}^{p}C_{F}\,C_{\mathcal{U}}\,\nu_{2}^{i}\,\Delta t\leq C_{2}\,\Delta t,$
where we have used that $\Phi_{2}(t_{i},t_{0})$ and the initial vector $U_{2}(t_{0})$ are uniformly bounded, and that the geometric series $\sum_{i=1}^{p}\nu_{2}^{i}$ converges to a constant independent of $\Delta t$. ∎
To apply the classical subspace perturbation theorem [19, Theorem 3.5], we express $\Psi(t_{p},t_{0})-F_{p,0}$ in an orthonormal basis of $\Psi(t_{p},t_{0})$’s invariant subspaces.
Lemma 6.
Under assumptions A1-A3, there exists an orthogonal matrix $P^{(0)}=[P_{1}^{(0)},P_{2}^{(0)}]$ of $\Psi(t_{p},t_{0})$, such that
${P^{(0)}}^{\mathrm{H}}\left[\Psi(t_{p},t_{0})-F_{p,0}\right]P^{(0)}=\begin{bmatrix}E_{11}&E_{12}\\ E_{21}&E_{22}\end{bmatrix}$, (42)
with blockwise norms: $\left\lVert E_{11}\right\rVert_{\mathrm{F}},\left\lVert E_{21}\right\rVert_{\mathrm{F}}=\mathcal{O}(\Delta t^{s}),\quad\left\lVert E_{12}\right\rVert_{\mathrm{F}},\left\lVert E_{22}\right\rVert_{\mathrm{F}}=\mathcal{O}(\Delta t).$
Proof.
Define $\operatorname{span}[V(t_{0})]:=\operatorname{span}[\mathcal{U}_{1}(t_{0}),\mathcal{U}_{2}(t_{0})]$ from the eigenvector matrix of $\Psi(t_{p},t_{0})$. Construct $P^{(0)}$ as its orthonormal basis of $\operatorname{span}[V(t_{0})]$ via QR decomposition, then $P^{(0)}R(t_{0})^{-1}=V(t_{0}),$ where $R(t_{0})^{-1}$ is an block upper-triangular matrix. Note that $R(t_{0})$ is uniformly bounded since $V(t_{0})$ has full column rank and $\|V(t_{0})\|_{\mathrm{F}}$ is uniformly bounded. From the definition of $P^{(0)}$, we have:
${P^{(0)}}^{\mathrm{H}}\left(\Psi(t_{p},t_{0})-F_{p,0}\right)P^{(0)}$
$=$ ${P^{(0)}}^{\mathrm{H}}\left(\Psi(t_{p},t_{0})-F_{p,0}\right)\left[\mathcal{U}_{1},\mathcal{U}_{2}\right](t_{0})\begin{bmatrix}R_{11}(t_{0})&R_{12}(t_{0})\\ &R_{22}(t_{0})\\ \end{bmatrix}$
$\equiv$ $\begin{bmatrix}E_{11}&E_{12}\\ E_{21}&E_{22}\end{bmatrix}(t_{0}).$
According to lemma˜5, the first block column can be bounded as $\mathcal{O}(\Delta t^{s})$ and the second block column as $\mathcal{O}(\Delta t)$. ∎
Since $\left\lVert E_{21}\right\rVert_{\mathrm{F}}=\mathcal{O}(\Delta t^{s})$, $P_{1}^{(0)}$ is an approximate invariant subspace of $F_{p,0}$. Our next work is to show that the difference between the invariant subspace of $F_{p,0}$ and $P_{1}^{(0)}$. This needs the help of the subspace perturbation theory.
From theorem˜1, we already have a quantitative lower bound for the subspace separation $\operatorname{sep}(\Psi_{11},\Psi_{22})$ based on the spectral information of $\Psi(t_{p},t_{0})$. This bound, together with the subspace perturbation theorem theorem˜2, yields the convergence of both the eigenvalues and corresponding subspace of $F_{p,0}$ towards those of $\Psi(t_{p},t_{0})$ as $\Delta t\to 0$. This result is summarized in theorem˜7.
Theorem 7 (Convergence of Eigenvalues and Invariant Subspaces).
Under assumptions A1-A3, the eigenpairs of $F_{p,0}$ admits the following convergence properties as $\Delta t\to 0$:
  1. 1.
    The $n$ largest eigenvalues of $F_{p,0}$ converge to the $n$ Floquet multipliers of $\Psi(t_{p},t_{0})$ with order $\mathcal{O}(\Delta t^{s})$.
  2. 2.
    The corresponding $n$-dimensional invariant subspace of $F_{p,0}$ converges to that of $\Psi(t_{p},t_{0})$ with order $\mathcal{O}(\Delta t^{s})$.
  3. 3.
    The remaining (d1)n(d-1)n eigenvalues of $F_{p,0}$ asymptotically converge to zero as $\Delta t\to 0$, with the convergence rate improving to $\mathcal{O}(\Delta t^{2})$ when the stepsize is constant.
Proof.
By theorem˜1, since $\Psi_{11}$ is independent of $\Delta t$ and the eigenvalues of $\Psi_{22}$ vanish as $\Delta t\to 0$, there exists a constant $C_{\text{sep}}$ such that $\operatorname{sep}(\Psi_{11},\Psi_{22})>2C_{\text{sep}}$ for sufficiently small $\Delta t$. From lemma˜6, we have $\left\lVert E_{11}\right\rVert+\left\lVert E_{22}\right\rVert=\mathcal{O}(\Delta t)$, hence $\left\lVert E_{11}\right\rVert+\left\lVert E_{22}\right\rVert\leq C_{\text{sep}},$ which ensures that the spectral gap condition $\delta\geq C_{\text{sep}}$ in theorem˜2 holds uniformly.
Next, the perturbation requirement $\left\lVert E_{21}\right\rVert\big(\left\lVert E_{12}\right\rVert+\left\lVert A_{12}\right\rVert\big)<\tfrac{1}{4}\delta^{2}$ is also satisfied for sufficiently small $\Delta t$, since $\left\lVert E_{21}\right\rVert=\mathcal{O}(\Delta t^{s})\to 0$ while the remaining terms stay bounded. Therefore, the conditions of theorem˜2 are met, implying the existence of a matrix $R$ such that $\left\lVert R\right\rVert\leq\frac{2}{C_{\text{sep}}}C_{1}\Delta t^{s}\equiv C_{2}\Delta t^{s},$ and
$F_{p,0}P_{1}^{(0)}\begin{bmatrix}I\\ R\end{bmatrix}=P_{1}^{(0)}\begin{bmatrix}I\\ R\end{bmatrix}(\Psi_{11}+\Delta\Psi_{11}),\qquad\Delta\Psi_{11}=E_{11}-(\Psi_{12}+E_{12})R.$ (43)
Hence, the invariant subspace of $F_{p,0}$ associated with the $n$ dominant eigenvalues deviates from that of $\Psi(t_{p},t_{0})$ by $\mathcal{O}(\Delta t^{s})$.
Finally, since $\Delta\Psi_{11}=\mathcal{O}(\Delta t^{s})$, the Bauer–Fike theorem [23] yields
$\max_{\lambda_{i}\in\lambda(\Psi_{11})}\max_{\mu_{j}\in\lambda(\Psi_{11}+\Delta\Psi_{11})}\left\lvert\lambda_{i}-\mu_{j}\right\rvert$ $\leq\left\lVert(\Psi_{12}+E_{12})R-E_{11}\right\rVert_{\mathrm{F}}$ (44)
$\leq\left[C_{R}\left\lVert\Psi_{12}+E_{12}\right\rVert_{\mathrm{F}}+C_{E}\right]\Delta t^{s}.$
The same reasoning applies to the remaining eigenpairs, which similarly converge to zero as $\Delta t\to 0$. ∎
The key mechanism underlying theorem˜7 is the use of the spectral separation between the physical and spurious subspaces of $\Psi(t_{p},t_{0})$: since all physical Floquet multipliers are strictly nonzero while the spurious ones vanish as $\Delta t\to 0$, a positive spectral gap guarantees the separation $\operatorname{sep}(\Psi_{11},\Psi_{22})>0$ and thus the convergence of the corresponding invariant subspaces.
In many applications, only the invariant subspace of the top $k$ Floquet multipliers is of interest. Assuming this subspace of $\Psi(t_{p},t_{0})$ remains separated from its complement as $\Delta t\to 0$, the same perturbation argument yields:
Corollary 8 (Convergence of Partial Floquet Multipliers and Subspaces).
Under the assumptions of theorem˜7, if the subspace separation between the invariant subspaces associated with the $k$ largest eigenvalues of $\Psi(t_{p},t_{0})$ and the remaining ones remains nonzero, then the $k$ largest eigenvalues of $F_{p,0}$ and their invariant subspace converge to the corresponding Floquet multipliers and invariant subspace of $\Psi(t_{p},t_{0})$ with rate $\mathcal{O}(\Delta t^{s})$.
In practice, the subspace separation $\operatorname{sep}$ is rarely computed explicitly. Instead, it is often estimated using the spectral gap, which, according to theorem˜1, serves as an upper bound for $\operatorname{sep}$. Hence, it is advisable to choose $k$ such that the $k$-th largest eigenvalue is well separated from the $(k+1)$-th one, with a spectral gap typically on the order of $\mathcal{O}(10^{-3})$ or larger. Empirically, when the spectral gap is sufficiently large, the actual separation tends to approach this upper bound rather than its lower limit, leading to markedly improved convergence behavior.
6 Algorithm and Complexity Analysis
We present pTOAR, an iterative method for computing the dominant $k$ eigenpairs of the periodic-PEP eq.˜18. It combines the periodic Arnoldi process [12] with TOAR-style basis compression [24] to reduce memory and computational cost, making it suitable for large-scale problems. The remainder of this section details the algorithmic steps and its underlying principles.
The periodic Krylov-Schur (pKS) algorithm [12] is well-suited for the linearized periodic-PEP eq.˜22. It avoids explicitly forming the full product $F_{i+p,i}=L^{(p)}\cdots\allowbreak L^{(1)}$, while maintaining accuracy even for eigenvalues spanning several orders of magnitude. The pKS algorithm constructs a periodic Arnoldi process:
$L^{(i)}V^{(i-1)}=V^{(i)}T^{(i)},$ (45)
where $V^{(i)}$ for $i=1,\ldots,p-1$ are $(nd)\times k$ orthonormal matrices, and $V^{(p)}\in\mathbb{C}^{nd\times(k+1)}$, which includes the additional Arnoldi residual vector. The projected matrices $T^{(i)}\in\mathbb{C}^{k\times k}$ are upper triangular, except for upper-Hessenberg matrix $T^{(p)}\in\mathbb{C}^{(k+1)\times k}$. Although the basis dimensions slightly differ at the boundaries, such variations do not affect the asymptotic memory complexity. For simplicity, we omit these boundary details in the subsequent discussion.
A limitation of pKS is that it disregards the linearized structure of $L^{(i)}$, resulting in a dimension increase from $n$ to $nd$ and leading to unnecessary computational overhead. To address this issue, we draw inspiration from the Two-level Orthogonal Arnoldi (TOAR) method [6], which was originally developed for polynomial eigenvalue problems to compress the basis arising from linearization.
We introduce pTOAR, a hybrid of the pKS method and TOAR-style compression. The key difference is that pTOAR stores the periodic Arnoldi basis $V^{(i)}$ in pKS in an implicit, compact form, greatly saving memory. We now show the details of the compression mechanism.
We explicitly substitute the block form of $L^{(i)}$ into eq.˜45, and partition the $V^{(i)}$ following the block structure of $L^{(i)}$, as shown in eq.˜46:
$\begin{bmatrix}&I_{n}&&\\ &&\ddots&\\ &&&I_{n}\\ A_{0}^{(i)}&A_{1}^{(i)}&\cdots&A_{d-1}^{(i)}\\ \end{bmatrix}\begin{bmatrix}V_{1}^{(i-1)}\\ \vdots\\ V_{d-1}^{(i-1)}\\ V_{d}^{(i-1)}\end{bmatrix}=\begin{bmatrix}V_{1}^{(i)}\\ \vdots\\ V_{d-1}^{(i)}\\ V_{d}^{(i)}\end{bmatrix}T^{(i)}.$ (46)
The key observation of pTOAR from eq.˜46 is an information overlap relationship across adjacent basis matrices:
$V_{j+1}^{(i)}T^{(i)}=V_{j}^{(i+1)},\,j=1,\ldots,d-1,\,i=1,\ldots,p.$
This implies that a single subspace can represent $d$ block segments drawn from $d$ consecutive basis matrices. Specifically, we define a shared basis $Q^{(i)}\in\mathbb{C}^{n\times k}$ such that
$\operatorname{span}(Q^{(i)})=\operatorname{span}(V_{d}^{(i)},V_{d-1}^{(i+1)},\ldots,V_{1}^{(i+d-1)}).$
Utilizing this structure, each Arnoldi basis $V^{(i)}$ can be compactly expressed by form:
$V^{(i)}=\begin{bmatrix}Q^{(i-d+1)}&&\\ &\ddots&\\ &&Q^{(i)}\end{bmatrix}\begin{bmatrix}U_{1}^{(i)}\\ \vdots\\ U_{d}^{(i)}\end{bmatrix}$,
where each $\left\{Q^{(j)}\right\}_{j=i-d+1}^{i}\in\mathbb{C}^{n\times k}$ is an orthonormal basis representing the shared subspace, and $\left\{U_{j}^{(i)}\right\}_{j=1}^{d}\in\mathbb{C}^{k\times k}$ are local representation matrices. It follows that the block-diagonal matrix $\operatorname{blkdiag}(Q^{(i-d+1)},\dots,Q^{(i)})$ is also orthogonal, which in turn implies that the tall matrix $\begin{bmatrix}(U_{1}^{(i)})^{\mathrm{T}}&\dots&(U_{d}^{(i)})^{\mathrm{T}}\end{bmatrix}^{\mathrm{T}}$ is also orthonormal.
The pTOAR algorithm proceeds through periodic Arnoldi basis construction, Ritz value extraction and reordering, and optional restarts, following a workflow analogous to that of pKS but differing in implementation details. We will discuss these differences step by step, and then estimate the memory and computational costs for both algorithms. Because a restart simply repeats this core procedure and its occurrence depends on the eigenvalue problem, and most practical cases requires no or only once restart, we estimate the costs by examining the non-restarted case.
Similar to pKS, the periodic Arnoldi basis update in pTOAR involves a matrix-vector product $L^{(i)}v^{(i-1)}$, whose result is used to update $Q^{(i)}$ and $U_{\cdot,j}^{(i)}$. Owing to the linearization structure of $L^{(i)}$, each product updates only the last block:
$L^{(i)}v^{(i-1)}=\left[{v_{2}^{(i-1)}}^{\mathrm{H}},\cdots,{v_{d}^{(i-1)}}^{\mathrm{H}},{v_{d}^{(i)}}^{\mathrm{H}}\right]^{\mathrm{H}},\,v_{d}^{(i)}={A_{d}^{(i)}}^{-1}\left(\sum_{j=0}^{d-1}A_{j}^{(i)}v_{j}^{(i-1)}\right),$
where $v^{(i-1)}$ is the last column of $V^{(i-1)}$, and $v_{j}^{(i-1)}$ is its $j$-th block of $v^{(i-1)}$, also the last column of $V_{j}^{(i-1)}$ in eq.˜46.
The second step in periodic Arnoldi basis construction is orthogonalization. Once $v_{d}^{(i)}$ is obtained, the pKS constructs and orthogonalizes $v^{(i)}$ against the current full basis $V^{(i)}$. In contrast, pTOAR first orthogonalizes $v_{d}^{(i)}$ against the current basis $Q^{(i)}$, and augments $Q^{(i)}$ with the normalized component. The representation of $v_{d}^{(i)}$ on the updated $Q^{(i)}$ is appended to the right of $U_{d}^{(i)}$. The remaining $U_{j}^{(i)}$ are extended from shifted representations $v_{j+1}^{(i-1)}$ in $Q^{(i-d+j)}$.
The computation of $v_{d}^{(i)}$ requires $d$ matrix-vector products, and one linear system solve, all in dimension $n$. Assuming matrix-vector products cost as $\mathcal{O}(n^{\gamma})$ and linear system solves cost as $\mathcal{O}(n^{\gamma+1})$, the total cost of constructing the periodic Arnoldi basis over $k$ iterations is $\mathcal{O}\left(pk(dn^{\gamma}+n^{\gamma+1})\right)$, where $\gamma=1$ for sparse and $\gamma=2$ for dense systems. In pTOAR, each $v_{j}^{(i-1)}$ needs to be explicitly reconstructed through $v_{j}^{(i-1)}=Q^{(i-d+j)}u_{j}^{(i-1)}$ which introduces an extra $2ndk$ cost per iteration. For orthogonalization, pKS applies classical Gram-Schmidt process to an $nd$-dimensional basis, incurring a cost of $2pndk^{2}+\mathcal{O}(pndk)$. In contrast, pTOAR orthogonalizes only the compressed basis $Q^{(i)}$, reducing this cost to $2pnk^{2}+\mathcal{O}(pnk)$. However, updating $U_{j}^{(i)}$ matrices introduces another $2pdk^{2}+\mathcal{O}(pdk)$ cost.
As for memory usage, instead of storing the all $p$ full basis $V^{(i)}\in\mathbb{C}^{nd\times k}$, pTOAR only stores $p$ subspace bases $Q^{(i)}$ and $pd$ representations $U_{j}^{(i)}$, reducing the overall memory cost from pndkpndk to $pnk+pdk^{2}$.
The extraction and reordering of Ritz values are identical for pKS and pTOAR. For a Krylov subspace of dimension $k$, this step applies the periodic QR algorithm, a direct method [11], to the projected periodic matrices $T_{k}^{(i)}$, incurring a computational cost of $\mathcal{O}(pk^{3})$ [6]. Each small projected matrix $T_{k}^{(i)}$ requires an additional $k^{2}$ memory, leading to a total of $pk^{2}$ across all $p$ slices.
At last, we evaluate memory requirement for storing the input data. For each $L^{(i)}$, it suffices to store the sampled matrices $G^{(i)}$ instead of the coefficient matrix $A_{j}^{(i)}$ in eq.˜14 and the full $nd\times nd$ matrix $L^{(i)}$. This leads to a memory cost of $n^{\gamma}$ per slice, where $\gamma=1$ if the matrices $A_{j}^{(i)}$ are sparse, and $\gamma=2$ if dense. Combining this with the basis memory and the Ritz extraction memory, the overall memory cost is thus $pn^{\gamma}+pnkd$ under assumption $k\ll n$.
Table˜1 compares the computational and memory costs of three algorithmic setups: one-step collocation with pKS, and multistep methods with pKS or pTOAR. Each leads to a distinct periodic eigenvalue problem for computing the dominant $k$ Floquet multipliers. Costs are measured per full Arnoldi process, as convergence mainly depends on the Floquet multipliers’ spectral distribution rather than the solver details. Complexity estimates for the condensation step in the one-step collocation method are from [6].
Algorithm Collocation Multistep Multistep
+pKS +pKS +pTOAR
memory Cost
Coefficient Matrices $pn^{2}$ $pn^{\gamma}$ $pn^{\gamma}$
Arnoldi Basis kpnkpn kdpnkdpn kpnkpn
Total Memory $pn^{2}$ $kdpn+pn^{\gamma}$ $kpn+pn^{\gamma}$
Computational Cost
[…]
Arnoldi Process $\mathcal{O}(pn^{3}k)$ $\mathcal{O}(pn^{\gamma}(n+d)k)$ $\mathcal{O}(pn^{\gamma}(n+d)k)$
Orthogonalization $2k^{2}pn$ $2k^{2}dpn$ $2k^{2}p(n+d)$
Ritz Value Extraction $4pk^{3}$ $4pk^{3}$ $4pk^{3}$
Total FLOPs $\mathcal{O}(pn^{3}(d^{3}+k))$ $\mathcal{O}(pn^{\gamma}(n+d)k)$ $\mathcal{O}(pn^{\gamma}(n+d)k)$
Table 1: Comparison of memory and computational costs for different methods, where $\gamma=1$ if input matrices are sparse, and $\gamma=2$ if they are dense.
As shown in table˜1, the collocation method suffers from the expensive condensation process and dense system matrices, while pKS and pTOAR avoid both. Compared to pKS, pTOAR further reduces memory and orthogonalization costs when $d>1$.
To compare with backward Euler, we can set $d=0$, since there is no extra matrix-vector product or memory cost in this case. For $d>1$, multistep methods incur extra matrix-vector products and memory for $pd$ matrices $U_{j}^{(i)}$, but this overhead is typically negligible compared to solving linear systems.
Finally, we are prepared to give the full algorithm for computing the Floquet multipliers via Discretization and iterative eigenvalue solvers.
Algorithm 1 Algorithm for Computing Dominant Floquet Eigenpairs with Multistep Discretization and Periodic Iterative Solvers
1:
2:Time samples $\{t_{i}\}_{i=0}^{p}$ with $0=t_{0}<t_{1}<\cdots<t_{p}=T$, where $T$ is the period;
3:Discrete evaluations $\{G(t_{i})\}_{i=0}^{p}$ of the system matrix;
4:Target number of eigenpairs $k$.
5:
6:Dominant $k$ Floquet multipliers and their corresponding eigenvectors.
7:Discretization: At each time $t_{i}$, define $\left\{A_{j}^{(i)}\right\}_{j=0}^{d}$ in eq.˜14 implicitly via matrix-vector products routines, and invoke them to define matrix-vector products of $L^{(i)}$, as ineq.˜21.
8:Eigenpair Computation: Apply the pTOAR to compute the dominant $k$ eigenpairs of the linearized periodic EVP eq.˜22.
7 Numerical Results
This section presents three numerical experiments designed to validate our proposed method. The first is an artificial example constructed from known Floquet eigenpairs to verify convergence theory and the distribution of Spurious eigenvalues. The second is a benchmark problem drawn from [25, 6], representative of scenario 1. We compare our pTOAR method with existing packages MATCONT [10] and AUTO-07p [9]. The third experiment targets an industrial circuit application, representative of scenario 2, where fixed sampling makes collocation infeasible and highlights the advantage of our multistep-based method. All numerical experiments are conducted using MATLAB 2024b on a system equipped with an Intel Core i5-1135G7 CPU at 2.4GHz and 16GB of RAM.
We assess the accuracy of the computed Floquet multipliers $\lambda$ and eigenvectors $x$ using relative errors:
$e_{\text{val}}=\frac{\left|\lambda-e^{\mu T}\right|}{\left|e^{\mu T}\right|},\quad e_{\text{vec}}=\frac{\|x_{p}-x(T)\|_{2}}{\|x(T)\|_{2}},$ (47)
where $e^{\mu T}$ and $x(T)$ are the exact Floquet multiplier and eigenvector. For the invariant subspace spanned by several eigenvectors, we define the subspace error between computed subspace $U$ and exact subspace $V$ by
$e_{\text{subspace}}=\Theta(U,V),$ (48)
where $\Theta(U,V)$ is the principal angle between the two subspaces. When exact solutions are not available, we use higher-order multistep methods to generate reference solutions, whose discrepancy from the exact operator is negligible for our evaluation purposes.
7.1 Convergence Behavior of Floquet multipliers and Spurious Eiegnvalues
Our first experiment is based on a two-dimensional LPTV system given by eq.˜49,
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}\alpha\cos^{2}t+(\beta-2)\sin^{2}t&1+(\beta-\alpha-2)\sin t\cos t\\ -1+(\beta-\alpha-2)\sin t\cos t&\alpha\sin^{2}t+(\beta-2)\cos^{2}t\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$. (49)
This model is obtained by linearizing the Stuart-Landau oscillator along its limit cycle. Its Floquet multipliers are $e^{2\pi\alpha}$ and $e^{2\pi(\beta-2)}$, and their corresponding eigenvectors are columns of
$U(t)=\begin{bmatrix}\cos t&\sin t\\ -\sin t&\cos t\end{bmatrix}$. (50)
We discretize one period $0,2\pi]$ with variable stepsize, and apply Gear2, Gear3 methods to obtain the periodic-EVP eq.˜22. The Floquet multiplier, its associated eigenvector, and the spurious eigenvalues from multistep methods are then computed using the periodic QR algorithm [11] with full precision. For reference, we also apply the backward Euler (BE) discretization, which does not produce spurious eigenvalues, to serve as a baseline for convergence comparison.
Through this experiment, we set $\alpha=0.1,\beta=0.1$. The system is discretized over one period into $2^{k}$ points with stepsizes alternating in the ratio $1:2:1:2:$... We study the convergence of the computed Floquet eigenpairs with respect to the largest stepsize.
(a) Floquet Eigenvalue Error vs. Maximum Discrete Stepsize
(b) Floquet Eigenvector Error vs. Maximum Discrete Stepsize
Figure 1: Convergence Behavior of Dominant Floquet Eigenpair
As shown in fig.˜1(a) and fig.˜1(b), BE, Gear2, and Gear3 exhibit first, second, and third order convergence, respectively, consistent with the theoretical prediction. Only the errors for the dominant Floquet multiplier and its eigenvector are shown here, as the results for the other mode are similar.
(a) Spurious Eigenvalues of Gear2
(b) Spurious Eigenvalues of Gear3
Figure 2: Convergence Behavior of Spurious Eigenvalues
Figures˜2(a) and 2(b) illustrate the convergence of the spurious eigenvalues associated with Gear2 (two spurious eigenvalues) and Gear3 (four spurious eigenvalues), respectively, as the discretization Stepsize decreases. Since these spurious eigenvalues are extremely small, their logarithms are plotted on the vertical axis, and different solid markers are used to distinguish individual eigenvalues. In addition, based on the multistep method and eq.˜28, we compute the theoretical reference $\nu^{p}$, which represents the expected magnitude of these spurious eigenvalues from the corresponding auxiliary STM, and plot its logarithm as red lines in Figures˜2(a) and 2(b). Strictly speaking, each spurious eigenvalue is equal to the product of the Floquet multiplier of the Spurious system and $\nu^{p}$, where the Floquet multiplier is of moderate size. Therefore, we only compare the order of magnitude, rather than the exact value. Consequently, the slope of the red line corresponds to $\log(\nu)$ and serves as a reference for expected convergence.
The figures show that the computed spurious eigenvalues match the expected magnitude very well, far exceeding the prediction of theorem˜7, which guarantees only linear decay with $\Delta t$.
7.2 Scenario 1: Coupled Stuart-Landau Oscillators
The first numerical experiment uses a coupled Stuart-Landau system adapted from [25, 6], whose analytical form is given by
[…]
We fix $\beta=0.5$ and use $\delta$ as the continuation parameter, approaching $\delta=0.25$ from below. Our focus is on the Floquet multipliers of the variational system associated to a limit cycle in eq.˜51, where two oscillators are $\pi$ radians out of phase, i.e., $x_{1}=-x_{2},y_{1}=-y_{2}$. As $\delta$ approaches $0.25$, this limit cycle approaches a heteroclinic loop, and the period grows raplidly, and the Floquet multipliers vary sharply. Our goal is to compute its Floquet multipliers with high accuracy to determine their position relative to the unit circle in the complex plane.
For a given pair of $(\delta,\beta)$, the periodic solution is first computed using existing continuation software—MATCONT [10] in MATLAB and AUTO-07p [26, 9] in FORTRAN. These tools provide not only the periodic orbit but also the Jacobian along the trajectory, enabling us to study the associated variational system.
We test three discretization and eigensolver combinations on the coupled Stuart-Landau system eq.˜51, as listed in table˜2.
Table 2: Comparison of Software, Discretization Methods, and Algorithms
Software Discretization Method Algorithm
AUTO-07p Collocation O.K.Floquet [25]
Modified MATCONT Collocation Periodic QR[11]
pTOAR Multistep Method pTOAR
The first method uses the built-in Floquet multiplier computation in AUTO-07p, which applies a collocation discretization and the O.K. Floquet algorithm [25]. The second method modifies MATCONT: it retains the collocation discretization but replaces the built-in MATLAB eig eigenvalue solver with a periodic QR algorithm [11] to compute all multipliers. The third method, proposed in this paper, discretizes the variational system using a multistep method and extracts the Floquet multipliers with a periodic Krylov-Schur algorithm.
To ensure a fair comparison, all methods use a total of 1024 sampling points per period. For AUTO-07p and the modified MATCONT, we set NTST = 512 and NCOL = 2.
For clarity of the figures, we plot the modulus of the computed Floquet multipliers on a logarithmic scale versus the period fig.˜3. Around suspected bifurcation points, the Floquet multipliers change very rapidly within narrow parameter intervals. Due to the limited number of computed points, these sharp transitions are not fully resolved in our figures.
(a) First Floquet Multiplier
(b) Second Floquet Multiplier
(c) Third Floquet Multiplier
(d) Fourth Floquet Multiplier
Figure 3: Evolution of Floquet Multipliers vs. Period of the Parameterized Limit Cycle
The three largest Floquet multipliers obtained via all three methods lie on very smooth curves, indicating that these multipliers are likely quite accurate. For the largest Floquet multipliers shown in fig.˜3, all three methods yield similar results. However, AUTO-07p fails to report the largest multiplier once it exceeds $10^{16}$. Aside from this limitation, its computed values remain consistent with other two methods. The second and third largest multipliers are nearly identical across all methods. For the smallest Floquet multiplier, both the modified MATCONT and our method indicate a monotonic decay. Notably, AUTO-07p incorrectly reports the fourth multiplier as an approximately constant value around $10^{-12}$ over a wide range of periods, while our results show a clear downward trend.
Figure 4: Error of Oscillatory Floquet Multiplier vs. Period of the Parameterized Limit Cycle
Figure˜4 shows the error in computing the oscillatory Floquet multiplier, which should be 1. Both AUTO-07p and the modified MATCONT maintain consistently high accuracy due to their use of fourth-order collocation. In comparison, our Gear4 method performs worse. The reason is as the period increases and the smallest multiplier approaches zero, the spectral gap between Floquet multipliers and spurious eigenvalues decreases. According to theorem˜1, this leads to the separation between Floquet subspace and spurious subspaces decreases, and thus reducing accuracy of each Floquet multipliers. Although the accuracy of the oscillatory multiplier computed by our method is slightly inferior, we are still able to maintain its value close to 1. Since the primary objective is to track the evolution of all four multipliers relative to the unit circle as a parameter varies, the overall performance remains satisfactory, as it still captures all four multipliers reasonably and correctly determines stability.
7.3 Scenario 2: RF Circuits
In radio frequency (RF) circuits, accurately predicting the phase noise of oscillators is essential for evaluating the performance and stability of oscillatory systems [5]. A critical step in this prediction involves computing the perturbation projection vector (PPV) [5], which corresponds to the eigenvector associated with the dominant Floquet multiplier of an LPTV system. The LPTV system takes the form:
$C(t)^{\mathrm{T}}\frac{\mathrm{d}y}{\mathrm{d}t}+G(t)^{\mathrm{T}}y=0,$ (52)
where $C(t)$ and $G(t)$ are $T$-periodic matrices with the following structure:
$C(t)=\begin{bmatrix}C_{11}(t)&0\\ 0&0\end{bmatrix},\quad G(t)=\begin{bmatrix}G_{11}(t)&G_{12}(t)\\ G_{21}(t)&G_{22}(t)\end{bmatrix},\quad y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}$.
Here, $C_{11}(t)$ and $G(t)$ are assumed to be nonsingular. The matrix $C_{11}(t)$ is referred to as the capacitance matrix, while $G(t)$ is the conductance matrix. The vector $y_{1}$ represents the part of the nodal voltages, and $y_{2}$ denotes the rest nodes. According to the DAE theory presented in [27], equation eq.˜52 is classified as a semi-explicit index-1 differential-algebraic equation (DAE). It can be systematically decoupled into an ODE and an algebraic constraint:
$\left\{\begin{aligned} &-\frac{\mathrm{d}}{\mathrm{d}t}y_{1}(t)+C_{11}^{-\mathrm{T}}(t)\left[G_{11}^{\mathrm{T}}(t)-G_{21}^{\mathrm{T}}(t)G_{22}^{-T}(t)G_{12}^{\mathrm{T}}(t)\right]y_{1}=0,\\ &y_{2}(t)=-G_{22}^{-T}(t)G_{12}^{\mathrm{T}}(t)y_{1}(t).\end{aligned}\right$. (53)
Our task is to solve the Floquet multiplier closest to 1 and its eigenvectors of the ODE part of eq.˜53, and thus recover the PPV. The physical nature of RF circuits leads to the following three characteristics of eq.˜53:
  1. 1.
    all Floquet multipliers have magnitudes strictly less than one, with only one or a few of them being close to $1$;
  2. 2.
    both the $C(t)$ and $G(t)$ are highly sparse;
  3. 3.
    only discrete samples of $C(t)$ and $G(t)$ are available at specific time points $0=t_{0}<t_{1}<t_{2}<\cdots<t_{p}=T$, and further sampling is not permitted.
Since MATCONT and AUTO-07p require explicit system equations to perform collocation at Gauss-Legendre points, they are not applicable in our setting where the sampling points are fixed. In contrast, our method is still applicable, and we avoid forming the explicit expression of $C_{11}^{-\mathrm{T}}(t)\left[G_{11}^{\mathrm{T}}(t)-G_{21}^{\mathrm{T}}(t)G_{22}^{-\mathrm{T}}(t)G_{12}^{\mathrm{T}}(t)\right]$.
Model Name Dimension Rank of $C(t)$ Total Sampled Points
Well Conditioned Case 18 16 448
Ill Conditioned Case 913 388 654
Table 3: PPV Cases Information
Table˜3 summarizes the basic properties of the two test cases. We use Gear4 as a benchmark and apply BE, Gear2, and Gear3 to discretize the systems, and then use the pTOAR algorithm to compute Floquet eigenpairs.
The goal is to verify as the discrete stepsize decrease, whether the convergence behavior observed in theory holds in practical, industry-scale circuits. Specifically, we examine whether the dominant eigenvector converges at the expected rate when it is well separated from the others, and whether the convergence instead occurs toward the invariant subspace when multiple Floquet multipliers cluster near 1.
(a) Floquet Multipliers Relative to the Unit Circle
(b) Separation between the Dominant Invariant Subspaces and the Remaining
(c) Dominant Eigenvector Error vs Stepsize
Figure 5: Floquet Multipliers, Subspace Separation, and Dominant Eigenvector Convergence for the Well Conditioned Case
In fig.˜5(a), the Floquet multipliers are plotted in blue relative to the unit circle in black. Most multipliers are clustered near the origin, except for a single one lying close to 1.
According to Theorem theorem˜1, the subspace separation between dominant $k$-dimension subspace (the invariant subspace of the dominant $k$ eigenvalues) and the remainder is bounded above by the spectral gap between the $k$-th and $(k+1)$-th largest eigenvalues. Consequently, the dominant one-dimensional subspace is expected to be well isolated, while this separation rapidly decay as more multipliers are included. This expectation is confirmed in fig.˜5(b), indicating this case is a well conditioned case for solving PPV.
To verify whether the theoretical convergence rates are achieved in practice, we computed the invariant subspace for the dominant Floquet multiplier using several multistep methods with varying maximum stepsizes. As shown in fig.˜5(c), all three methods converge at the expected theoretical rates.
In contrast, the ill-conditioned case features multiple Floquet multipliers near 1, as shown in fig.˜6(a). Figure˜6(b) reveals that although the dominant 13-dimensional invariant subspace is well-separated from the remainder, the dominant one-dimensional subspace itself is poorly isolated. This makes accurately computing the PPV particularly challenging.
(a) Floquet Multipliers Relative to the Unit Circle
(b) Separation between the Dominant Invariant Subspaces and the Remaining
Figure 6: Floquet Multipliers, Subspace Separation for the Ill Conditioned Case
In fig.˜7, we examine the convergence behavior of both the dominant eigenvector and the 13-dimensional subspace for the ill conditioned case. Figure˜7(a) shows that the dominant Floquet eigenvector fails to converge. In contrast, Figure˜7(b) demonstrates that the dominant 13-dimensional subspace converges at expected rate. This indicates that the clustering of the leading Floquet multipliers causes interference among their corresponding eigenvectors, degrading the convergence of any individual eigenvector. However, the collective subspace remains numerically recoverable with reasonable accuracy.
(a) Dominant Eigenvector Error vs. Maximum Stepsize
(b) Dominant 13-dim. Subspace Error vs. Maximum Stepsize
Figure 7: Convergence Behavior of Dominant Subspaces of the Ill Conditioned Case
(a) Well Conditioned Case
(b) Ill Conditioned Case
Figure 8: Computation Time vs. Step Numbers for the Well and Ill Conditioned Cases.
Figure˜8 presents the computational time required by our algorithm under different discretization methods and varying numbers of time steps. Excluding a few noticeable outliers, the results show that for a fixed discretization method, the computational time increases linearly with the number of time steps. Similarly, for a fixed number of time steps, the use of higher-order multistep methods does not lead to a significant increase in runtime. These observations are consistent with our earlier analysis of the algorithm’s computational complexity.
8 Conclusion
A multistep discretization of an LPTV system gives rise to a periodic-PEP for computing Floquet multipliers and their invariant subspaces. This periodic-PEP exhibits clear spectral separation: as the stepsize tends to zero, the dominant eigenvalues converge to the true Floquet multipliers, while the remaining ones decay to zero. By subspace perturbation theory, the computed multipliers and invariant subspaces converge to their exact counterparts at a rate matching the consistency order of the multistep method. Building on this, we propose the pTOAR algorithm to solve all or part of the Floquet eigenpairs efficiently. Numerical experiments demonstrate that the algorithm effectively addresses the two common scenarios, as outlined in the introduction.
Declarations
  • Funding: The research leading to these results received funding from National Key R&D Program of China under Grant Agreement No 2020YFA0711900, 2020YFA0711902.
  • Conflict of interest/Competing interests (check journal-specific guidelines for which heading to use): The authors have no conflicts of interest to declare that are relevant to the content of this article.
  • Ethics approval and consent to participate: Not applicable
  • Consent for publication:Not applicable.
  • Data availability: The datasets generated during and/or analyzed during the current study are available from the corresponding author on reasonable request.
  • Materials availability: Not applicable.
  • Code availability:The code used for the numerical experiments is available from the corresponding author upon reasonable request.
  • Author contribution: Yehao Zhang led the research and contributed to all aspects of this work, including the theoretical development, algorithm design, and numerical experiments. Yuncheng Xu and Yichen Tan participated in the theoretical discussions and provided suggestions for the design of the numerical experiments. Yangfeng Su supervised the research and provided guidance on the overall structure and direction of the study. All authors reviewed and approved the final manuscript.
If any of the sections are not relevant to your manuscript, please include the heading and write ‘Not applicable’ for that section.

### v2

@@
-High-order Computation of Floquet Multipliers and Subspaces using Multistep Methods
+
 Abstract
-Accurate and efficient computation of Floquet multipliers and subspaces is essential for analyzing limit cycle in dynamical systems and periodic steady state in Radio Frequency (RF) simulation.
-This problem is typically addressed by solving a periodic linear eigenvalue problem, which is discretized from the linear periodic time-varying system using one-step methods.
-The backward Euler method offers a computationally inexpensive overall workflow but has limited accuracy.
-In contrast, one-step collocation methods achieve higher accuracy through over-sampling, explicit matrix construction, and condensation, thus become costly for large-scale sparse cases.
-We apply multistep methods to derive a periodic polynomial eigenvalue problem, which introduces additional spurious eigenvalues.
-Under mild smoothness assumptions, we prove that as the stepsize decreases, the computed Floquet multipliers and their associated invariant subspace converge with higher order, while the spurious eigenvalues converge to zero.
-To efficiently solve large-scale problems, we propose pTOAR, a memory-efficient iterative algorithm for computing the dominant Floquet eigenpairs.
-Numerical experiments demonstrate that multistep methods achieves high order accuracy, while its computational and memory costs are only marginally higher than those of the backward Euler method.
+Accurate and efficient computation of Floquet multipliers and subspaces is essential for analyzing limit cycle in dynamical systems and periodic steady state in Radio Frequency simulation.
+This problem is typically addressed by solving a periodic linear eigenvalue problem, which is discretized from the linear time-periodic system using one-step collocation methods.
+Collocation methods become costly for large-scale cases.
+Our alternative approach is to use multistep methods.
+The multistep method leads to a periodic polynomial eigenvalue problem (pPEP), and introduces additional parasitic periodic eigenvalues.
+We prove that as the stepsize decreases, the computed Floquet multipliers and their associated invariant subspace converge with higher order, while the parasitic periodic eigenvalues converge to zero geometrically and therefore Floquet multipliers are not affected by those parasitic ones.
+A memory-efficient algorithm pTOAR is designed to solve the large-scale pPEP.
+Its computational and memory costs are almost independent of the choice of multistep methods.
+Numerical results coincide with our convergence analysis, and also demonstrate the efficiency of pTOAR.
 keywords:
 Floquet Multiplier, Multistep Method, Periodic Polynomial Eigenvalue Problem, Subspace Perturbation
-pacs:
-MSC Classification]65L15, 65F15, 65L70
+{MSCcodes}
+65L15, 65F15, 65L70
 1 Introduction
-A linear periodic time-varying (LPTV) system is described by eq.˜1:
-$\frac{\mathrm{d}x(t)}{\mathrm{d}t}=G(t)x(t),$ (1)
+A linear time-periodic (LTP) system is described by
+(1) $\frac{\mathrm{d}x(t)}{\mathrm{d}t}=G(t)x(t),$
 where $G(t)\in\mathbb{R}^{n\times n}$ is a $T$-periodic matrix function, i.e., $G(t+T)=G(t)$, and $x(t)\in\mathbb{R}^{n}$ is the state vector.
-The state transition matrix $\Phi(T,0)$ of eq.˜1 is obtained by integrating eq.˜1 from $[0,T]$ with initial value matrix $X(0)=I_{n}$.
-The Floquet multipliers and eigenvectors are defined as the eigenpairs of $\Phi(T,0)$.
-This paper applies high-order multistep methods to discretize an LPTV system eq.˜1, for efficient and accurate computation of its Floquet multipliers and corresponding subspace.
-The LPTV systems find applications in a wide range of areas, including dynamic simulation [1], communications [2], control theory [3, 4] and Radio Frequency circuit simulation [5].
-A central problem in these fields is the stability analysis of limit cycles in nonlinear autonomous systems.
-The stability is determined by the variational equation along the cycle.
-Since the limit cycle is $T$-periodic, its associated Jacobian $G(t)$ is also $T$-periodic, making the variational equation an LPTV system.
-Floquet theory then provides the framework: the stability is completely characterized by the Floquet multipliers and eigenvectors of this system [6].
-Depending on different application background and goals, there are two common scenarios in practice.
-First, when the matrix $G(t;\delta)$ is given as an explicit function of time $t$, with $\delta$ as a parameter, tracking how Floquet multipliers evolve with respect to this parameter provides insight into the system’s stability.
-In such cases, a multiplier crossing the unit circle from inside to outside typically signals the onset of a bifurcation or instability [6].
-Second, when $G(t)$ is only available at discrete time samples, without an explicit analytic form, the analysis often focuses on computing the Floquet eigenvector associated with the dominant multiplier.
-In applications such as circuit oscillator design, this largest multiplier is typically equal to one [5].
-The corresponding left Floquet eigenvector, known as the perturbation projection vector (PPV) [5, 7], characterizes the system’s phase sensitivity to external perturbations and plays a key role in modeling phase noise in Radio Frequency oscillator systems [5, 8].
-The computation of Floquet multipliers and eigenvectors generally involves two key steps: discretization and solving the resulting eigenvalue problem.
-Existing high-precision software packages such as AUTO-07p [9] and MATCONT [10] adopt high-order one-step collocation methods for discretization, resulting in a periodic eigenvalue problem (periodic-EVP) [6], which can be solved by periodic QR algorithm [11, 12].
-This framework makes them well-suited for the first scenario, where the parameterized periodic matrix function $G(t;\delta)$ is explicitly defined.
-However, the collocation method comes with several limitations for the second scenario, where the system matrix is only available at discrete time samples.
-It requires over-sampling at Gaussian collocation points within each subinterval, explicit construction of large system matrices, and a condensation step to eliminate over-sampled collocation points.
-This workflow is computationally expensive and memory-intensive for large-scale systems, and it is not well-suited for efficiently computing Floquet eigenvectors [6].
-Motivated by these limitations, and aiming to address both scenarios, this paper applies a multistep method to discretize eq.˜1 that induces a polynomial periodic eigenvalue problem (periodic-PEP)
-$\left(\theta^{d}+\sum_{j=0}^{d}\mathcal{A}_{j}\mathcal{C}^{d-j}\theta^{j}\right)u=0,$
-where $\mathcal{A}_{j}\mathcal{C}^{d-j}$ are block cyclic matrices constructed from the discretization method.
-Compared to traditional collocation-based methods, our approach does not require an explicit expression of the system, avoids sampling at fixed Gaussian collocation points, eliminates the need for condensation, and does not explicitly construct or store large coefficient matrices.
-It naturally supports the simultaneous computation of Floquet multipliers and their eigenvectors, making it especially suitable for large-scale and data-driven systems.
-The main contributions of this paper are as follows:
-  1. 1.
-    We apply $d$-step method to an LPTV system, corresponding to a periodic-PEP of degree $d$.
-We find that the $d$-step discretization generates d1d-1 spurious LPTV systems, whose designed $d$-step discretization corresponds to the same periodic-PEP.
-  2. 2.
-    We establish convergence analysis based on subspace perturbation theory, as the maximum stepsize $\Delta t$ goes to 0.
-Among $nd$ eigenpairs of the periodic-PEP, $n$ of them converge to Floquet eigenpairs at rate $\mathcal{O}(\Delta t^{s})$, where $s$ is the consistency order of the multistep method.
-The remaining n(d1)n(d-1) eigenvalues are related to spurious LPTV systems and decay to 0.
-  3. 3.
-    We propose an iterative algorithm pTOAR to solve the dominant eigenpairs of the periodic-PEP, and analyze its computational and memory costs.
+The monodromy matrix of Eq.˜1 is obtained by integrating Eq.˜1 from 0 to $T$ with initial value matrix $X(0)=I_{n}$.
+The Floquet multipliers and eigenvectors are the eigenpairs of the monodromy matrix.
+Efficient and accurate computation of Floquet multipliers and corresponding eigenvectors is essential for analyzing LTP systems.
+The LTP systems find applications in a wide range of areas, including dynamic simulation [1], communications [15], control theory [6, 4] and Radio Frequency circuit simulation [9].
+The central problem in these fields is the stability analysis for the limit cycles of nonlinear autonomous systems.
+The stability is completely characterized by the Floquet multipliers and eigenvectors of the variational equation along the limit cycle in the form of Eq.˜1 [20].
+The computation of Floquet multipliers and eigenvectors generally involves two key steps: discretization of Eq.˜1 and solving the resulting eigenvalue problem.
+The simplest implementation is to use backward Euler (BE) method to obtain a periodic eigenvalue problem (pEVP), which can be solved by periodic QR algorithm [5, 18].
+Existing high-precision software packages such as AUTO-07p [12] and MATCONT [11] adopt high-order one-step collocation methods for discretization.
+Collocation methods require extra function evaluations at Gaussian collocation points within each subinterval, and a condensation step to eliminate the extra collocation points.
+This framework makes collocation well-suited for modest-scale cases whose coefficient function $G(t)$ is given by an analytical expression.
+For large-scale systems, collocation methods become computationally expensive and memory-intensive [20], particularly as the condensation may break the inherent sparsity of the systems.
+In some practical cases, collocation methods also struggle if extra function evaluations are not permitted.
+This paper applies multistep methods to discretize Eq.˜1 that induces a periodic polynomial eigenvalue problem (pPEP).
+Multistep methods improve accuracy under the same sampling used by BE.
+However, applying a $d$-step method, the resulting pPEP possesses $nd$ periodic eigenpairs under cyclic symmetry; and the pPEP is much more complicated to solve.
+By introducing parasitic LTP systems, we prove that n(d1)n(d-1) periodic eigenvalues decay to 0 at rate $\mathcal{O}(\Delta t)$ as the maximum stepsize $\Delta t$ goes to 0.
+The remaining $n$ eigenpairs converge to Floquet eigenpairs at rate $\mathcal{O}(\Delta t^{s})$, where $s$ is the consistency order of the multistep method.
+Therefore, we only need to solve the $n$ dominant periodic eigenvalues of pPEP.
+To efficiently solve the eigenpairs of the pPEP, we propose the pTOAR.
 In particular, we show that using high-order multistep method incurs only negligible additional cost compared to the first-order one-step backward Euler method.
@@
 In Section˜2, the necessary background on Floquet theory, one-step collocation methods, multistep methods, and subspace perturbation theory are reviewed.
-In section˜3, we analyze how applying a multistep discretization to an LPTV system gives rise to a periodic-PEP.
-In section˜4, we introduce the spurious systems used to interpret the eigenpairs of the periodic-PEP and define the target objects to which they are expected to converge.
-In section˜5, the convergence results are established for the eigenpairs using tools from subspace perturbation theory.
-In section˜6, we present the proposed pTOAR algorithm for computing the dominant eigenpairs of large-scale periodic-PEP, together with an analysis of its computational and memory costs.
-Numerical experiments are reported in Section˜7, and conclusions are drawn in Section˜8.
+In Section˜3, we discuss the formulation of pPEP.
+In Section˜4, the convergence results are established for the eigenpairs using tools from subspace perturbation theory.
+In Section˜5, we present the pTOAR for solving pPEP, together with an analysis of its computational and memory costs.
+Numerical experiments are reported in Section˜6, and conclusions are drawn in Section˜7.
 2 Preliminaries
-2.1 Floquet Theory via the State Transition Matrix of LPTV Systems
-Our goal is to compute Floquet multipliers and eigenvectors, defined as the eigenpairs of the state transition matrix (STM) over one period.
-To this end, we first recall the definitions of these quantities and their key properties, which constitute the core of Floquet theory.
-The comprehensive Floquet theory can be found in [13, Chapter 5] and [14, Chapter 2].
-A typical LPTV system with period $T$ is defined as eq.˜1.
-Let $X(t)$ denote a fundamental solution matrix, defined by $\dot{X}(t)=G(t)X(t)$, where the initial value matrix $X(0)=X_{0}$ is any full-rank matrix.
-The associated STM from $s$ to $t$ is $\Phi(t,s)=X(t)X(s)^{-1}$.
-Note that the definition of $\Phi(t,s)$ is independent of the particular choice of $X_{0}$.
-The periodicity of the system matrix $G(t)$ implies that the STM $\Phi(t+T,t)$ is similar to $\Phi(T,0)$ for any $t$, as expressed in eq.˜2:
-$\Phi(t+T,t)=\Phi(t,0)\,\Phi(T,0)\,\Phi^{-1}(t,0).$ (2)
-To establish eq.˜2, we first note the periodic shift property of the STM,
-$\Phi(t+T,T)=\Phi(t,0),\qquad\forall t\in\mathbb{R},$ (3)
-which can be proven as follows.
-Define $Y(t)\equiv X(t+T)\,X^{-1}(T)\,X_{0},$ where $X(t)$ is the fundamental solution matrix.
-Then $\dot{Y}(t)=G(t)Y(t)$ and $Y(0)=X_{0}$.
-By the uniqueness of solutions to linear ODEs, we have $Y(t)\equiv X(t)$, which proves eq.˜3.
-Next, by decomposing $\Phi(t+T,0)$ at $T$ and $t$, we have
-$\Phi(t+T,t)\Phi(t,0)=\Phi(t+T,T)\Phi(T,0).$
-Right multiplying $\Phi(t,0)^{-1}$ on both sides and then using eq.˜3 yields eq.˜2.
-The similarity relation eq.˜2 implies that all matrices $\Phi(t+T,t)$ share the same eigenvalue matrix $R$, whose diagonal entries are the Floquet multipliers.
-Moreover, there exists a $T$-periodic matrix $U(t)$ such that
-$\Phi(t+T,t)=U(t)RU^{-1}(t),$ (4)
-where the columns of $U(t)$ and the rows of $U^{-\mathrm{H}}(t)$ provide the corresponding right and left Floquet eigenvectors, respectively.
-To construct $U(t)$ explicitly, we should use the the eigen-decomposition at $t=0$, that $\Phi(T,0)=U(0)RU^{-1}(0),$ and write $R=\exp(MT)$, where the diagonal entries of $M$ are the Floquet exponents.
-Now, we define $U(t)$ as:
-$U(t)=\Phi(t,0)U(0)\exp(-Mt).$ (5)
-In summary, Floquet multipliers and eigenvectors are obtained from the eigenvalue problem
-$\Phi(T,0)u(0)=u(0)\rho,\quad\rho=e^{\mu T},$ (6)
-at $t=0$, while the Floquet eigenvectors at any other time $t$ can be recovered using eq.˜5. Therefore, a single eigen-decomposition of $\Phi(T,0)$ suffices to determine all Floquet quantities.
+2.1 Floquet Theory
+Floquet theory provides the structure of the solutions of LTP systems Eq.˜1.
+We recall the key results of Floquet theory, and the comprehensive version can be found in [2, Chapter 5] and [14, Chapter 2].
+For LTP system Eq.˜1 with period $T$, let $X(t)$ denote a fundamental solution matrix, defined by $\dot{X}(t)=G(t)X(t)$, where the initial value matrix $X(0)=X_{0}$ is any nonsingular matrix.
+Define the state transition matrix $\Phi(t,s)$ by as the mapping from the principal fundamental solution matrix at time $s$ to that at time $t$, that $\Phi(t,s)X(s)=X(t)$.
+Theorem 2.1 (Floquet Representation Theorem [2][Theorem 20.9]).
+For LTP system Eq.˜1, there exists a nonsingular constant matrix $M$ and a $T$-periodic matrix function $U(t)\in C^{1}(\mathbb{R},\mathbb{C}^{n^{2}})$ such that its state transition matrix $\Phi(t,0)$ can be written as
+(2) $\Phi(t,0)=U(t)e^{Mt}.$
+Thus, the state transition matrix over a period $\Phi(t+T,t)$ can be decomposed by:
+(3) $\Phi(t+T,t)=U(t)e^{MT}U^{-1}(t),$
+Since $M$ is a constant matrix, and $T$ is the period, all matrices $\Phi(t+T,t)$ are similar.
+The eigenvalues of $M$ and $e^{MT}$ are called Floquet exponent and Floquet multipliers.
+For simplicity, we assume all Floquet multipliers are semi-simple.
+Thus, we assume $e^{MT}$ is diagonal, and Eq.˜3 becomes an eigenvalue decomposition of $\Phi(t+T,t)$.
+The columns of $U(t)$ and the rows of $U^{-\mathrm{H}}(t)$ provide the corresponding right and left Floquet eigenfunctions, respectively.
+To obtain the Floquet multipliers, we only need to consider the eigenvalue problem Eq.˜4, which corresponds to Eq.˜3 at $t=0$:
+(4) $\Phi(T,0)u(0)=u(0)e^{\mu T}.$
+$u(0)$ is called the Floquet eigenvector associated with the Floquet multiplier $\lambda=e^{\mu T}$.
+Though Eq.˜4 is only a slice of Eq.˜3 at $t=0$, it is sufficient to solve Floquet eigenfunctions $U(t)$ within the period.
+Once all the Floquet multipliers $e^{MT}$ are obtained, we can figure out all the Floquet exponents $M$, and thus solve Floquet eigenfunctions $U(t)$ from Eq.˜2.
+We should also note that $u(0)$ in Eq.˜4 is the value of the Floquet eigenfunction at $t=0$.
 2.2 One-step Methods for Solving Floquet Eigenpairs
-Now, to numerically solve eq.˜6, we recall the existing one-step discretization of eq.˜1.
-The period $[0,T]$ is partitioned at time points $0=t_{0}<t_{1}<\cdots<t_{p}=T$, with stepsizes $\Delta t_{i}=t_{i}-t_{i-1}$.
-In general, we allow the stepsizes to be variable, and the maximum stepsize is $\Delta t$.
-Then, we write eq.˜6 on the time grid:
+Let the period $[0,T]$ be partitioned at time points $0=t_{0}<t_{1}<\cdots<t_{p}=T$, with stepsizes $h_{i}=t_{i}-t_{i-1}$ allowed to be variable, and with maximum $h$. Rewrite Eq.˜3 on the time grid:
 $\Phi(t_{p},t_{p-1})\cdots\Phi(t_{2},t_{1})\Phi(t_{1},t_{0})u(t_{0})=u(t_{0})e^{\mu T}.$
-When $\Delta t$ is small enough, one-step methods approximate each $\Phi(t_{i},t_{i-1})$ by a matrix $A^{(i)}$ for $i=1,\ldots,p$, with $u_{i}$ denote the numerical solution at time $t_{i}$.
-This leads to a periodic eigenvalue problem to compute Floquet eigenpairs:
-$A^{(p)}\cdots A^{(2)}A^{(1)}u_{0}=u_{0}\lambda.$ (7)
-A well-known one-step method is backward Euler (BE), in which each transition matrix is given by $A^{(i)}=(I-G(t_{i})\Delta t_{i})^{-1}$ when applied to eq.˜1.
-This method only requires sampling of $G(t_{i})$ at time $t_{i}$.
-However, its first-order accuracy is usually insufficient for precise Floquet analysis.
-High-order one-step methods, such as collocation, improve discretization accuracy.
-Existing public software packages for computing and analyzing limit cycles, such as MATCONT [10] and AUTO-07p [9], use collocation to discretize periodic systems, and then form the discrete local state transition matrices $A^{(i)}$ to compute Floquet multipliers.
-In what follows, we briefly outline this procedure; a comprehensive process can be found in [6].
-Consider an interval $t_{i-1},t_{i}]$, we first choose the number of collocation points $k$, where a larger $k$ corresponds higher local approximation accuracy.
+When $h$ is small enough, one-step methods approximate each $\Phi(t_{i},t_{i-1})$ by a matrix $A^{(i)}$ for $i=1,\ldots,p$, and approximate the eigenpair $\left(u(t_{0}),e^{\mu T}\right)$ by $\left(u_{0},\lambda\right)$.
+This leads to a periodic eigenvalue problem [18]:
+(5) $A^{(p)}\cdots A^{(2)}A^{(1)}u_{0}=u_{0}\lambda.$
+One of the simplest one-step method is backward Euler (BE).
+BE only requires function evaluation of $G(t_{i})$ at time $t_{i}$, and $A^{(i)}=\left[I-G(t_{i})h_{i}\right]^{-1}$ when applied to Eq.˜1.
+However, its first-order accuracy is usually insufficient in practical.
+High-order one-step methods, such as collocation, improve discretization accuracy. In what follows, we briefly outline this procedure; a comprehensive process can be found in [20].
+The first step is to choose the number of collocation points $k$ within an interval $t_{i-1},t_{i}]$, where a larger $k$ corresponds higher local approximation accuracy.
 Collocation methods introduce internal stage variables at $k$ equidistant nodes and constructs a Lagrange polynomial interpolant through these points to approximate the solution $x(t)$.
-The interpolant is then required to satisfy eq.˜1 at $k$ Gaussian collocation points within $t_{i-1},t_{i}]$, which needs sampling the system matrix $G(t)$ at these addtional Guassian collocation points, leading to oversampling.
+The interpolant is then required to satisfy Eq.˜1 at $k$ Gaussian collocation points within $t_{i-1},t_{i}]$, which needs extra function evaluation of the system matrix $G(t)$ at these additional Gaussian collocation points.
 The next step, known as condensation, is eliminating the intermediate stage variables to obtain the one-step relation $x_{i}=A^{(i)}x_{i-1}$, typically via an LU elimination [20].
-This process yields an explicit construction of the local STM $A^{(i)}$.
-Although collocation methods have high numerical accuracy, their limitations become apparent depending on the application scenario.
-In the first scenario, where $G(t)$ is analytically known, and its dimension is small, the overhead of collocation is manageable.
-However, in the second scenario, where $G(t)$ is only available on given time grids, the collocation methods become both impractical and costly:
+This process yields an explicit construction of the local state transition matrix $A^{(i)}$.
+Collocation methods achieve higher accuracy than BE, and are efficient for cases that $G(t)$ is analytically known.
+It is also adopted by existing mature public software packages for computing and analyzing limit cycles, such as MATCONT [11] and AUTO-07p [12].
+However, if $G(t)$ is only available on given time grids, the collocation methods become both impractical and costly:
   1. 1.
-    they rely on temporal internal sampling on Gaussian points of each interval for accuracy, which is expensive and often infeasible in large-scale applications;
+    they rely on temporal internal function evaluation on Gaussian points of each interval for accuracy, which is expensive and often infeasible in large-scale applications;
   2. 2.
@@
     the condensed transition matrix $A^{(i)}$ is typically dense and must be formed explicitly, increasing memory usage and limiting the advantage of matrix-free solvers.
-2.3 Multistep Methods
-We now introduce multistep methods, a class of discretization schemes that attain high accuracy by reusing previously sampled values at adjacent grid points.
-We recall their general form and the standard definitions of consistency and stability for multistep methods [15, 16].
-Definition 1 (Multistep Methods [15][Chapter 3.5] ).
-For a fixed time interval $[0,T]$, we still consider a time grid $0=t_{0}<\cdots<t_{p}=T$.
-For an ODE $\dot{x}=f(t,x)$, a $d$-step method updates the solution at $t_{i}$ by combining the past $d$ steps as:
-$\sum_{j=0}^{d}\alpha_{j,i}x_{i-d+j}=\Delta t_{i}\left(\sum_{j=0}^{d}\beta_{j,i}f(t_{i-d+j},x_{i-d+j})\right),$ (8)
-where the coefficients $\alpha_{j,i}$ and $\beta_{j,i}$ depend on the past d1d-1 local stepsize ratios $\omega_{i},\omega_{i-1},\ldots,\omega_{i-d+1}$, with $\omega_{i}=\Delta t_{i}/\Delta t_{i-1}$.
-We normalize $\alpha_{d,i}=1$ and require $\left\lvert\alpha_{0,i}\right\rvert+\left\lvert\beta_{0,i}\right\rvert\neq 0$ to ensure all $d$ steps contribute.
-For detailed derivations of these coefficients, see [15, Chapter 3]. As an illustration, the second-order BDF method (Gear2) has coefficients given by eq.˜9.
-$\alpha_{0,i}=\frac{{\omega_{i}}^{2}}{1+2\omega_{i}},\,\alpha_{1,i}=-1-\alpha_{0,i},\alpha_{2,i}=1,\,\beta_{2,i}=\frac{1+\omega_{i}}{1+2\omega_{i}},\,\beta_{1,i}=\beta_{0,i}=0.$ (9)
+2.3 Multistep Methods for Solving Initial Value Problems
+Multistep methods offer an alternative approach to enhance discretization accuracy by utilizing previously computed values at neighboring grid points.
+We review their general form and recall standard notions of consistency and stability [17].
+These results are essential for solving initial value problems, and also provide the foundation for our later analysis of Floquet multipliers using multistep discretizations.
+Definition 2.2 (Variable Stepsize Multistep Methods [17, Chapter 3.5]).
+Let $0=t_{0}<\cdots<t_{p}=T$ be a time grid on the interval $[0,T]$.
+We define the stepsize $h_{i}=t_{i}-t_{i-1}$ and the stepsize ratio $\omega_{i}=h_{i}/h_{i-1}$.
+We define the ratio vector $\boldsymbol{\omega}_{i}\in\mathbb{R}^{d}$ as: $\boldsymbol{\omega}_{i}=(\omega_{i-d+1},\ldots,\omega_{i}).$ For an ODE $\dot{x}=f(t,x)$, a $d$-step method updates the solution at $t_{i}$ according to:
+(6) $\sum_{j=0}^{d}\alpha_{j}^{(i)}x_{i-d+j}=h_{i}\sum_{j=0}^{d}\beta_{j}^{(i)}f(t_{i-d+j},x_{i-d+j}),$
+where the coefficients are determined by smooth functions of the local step ratios, i.e., $\alpha_{j,i}=\alpha_{j}(\boldsymbol{\omega}_{i})$ and $\beta_{j,i}=\beta_{j}(\boldsymbol{\omega}_{i})$, and $\left\lvert\alpha_{0,i}\right\rvert+\left\lvert\beta_{0,i}\right\rvert\neq 0$ for irreducibility condition.
+For detailed derivations of these coefficients, see [17, Chapter 3]. As an illustration, the second-order BDF method (Gear2) has coefficients given by Eq.˜7.
+(7) $\alpha_{0,i}=\frac{{\omega_{i}}^{2}}{1+2\omega_{i}},\,\alpha_{1,i}=-1-\alpha_{0,i},\alpha_{2,i}=1,\,\beta_{2,i}=\frac{1+\omega_{i}}{1+2\omega_{i}},\,\beta_{1,i}=\beta_{0,i}=0.$
 We now turn to the fundamental notions of consistency and stability for such multistep methods.
-Definition 2 (Consistency [15][Chapter 3, 5.17]).
-A multistep method of the form eq.˜8 is said to be consistent of order $s$ if, when applied to a sufficiently smooth function $x(t)$, its local truncation error $R_{i}$ satisfies $\left\lVert R_{i}\right\rVert=\mathcal{O}(\Delta t_{i}^{s+1})$ as $\Delta t_{i}\to 0$, where
-$R_{i}=\sum_{j=0}^{d}\alpha_{j,i}x(t_{i-d+j})-\Delta t_{i}\sum_{j=0}^{d}\beta_{j,i}\frac{\mathrm{d}x}{\mathrm{d}t}(t_{i-d+j}).$ (10)
+Definition 2.3 (Consistency [17][Chapter 3, 5.17]).
+A multistep method Eq.˜6 is consistent of order $s$ if, when applied to a sufficiently smooth function $x(t)$, its local truncation error $E_{i}$ satisfies $\left\lVert E_{i}\right\rVert=\mathcal{O}(h_{i}^{s+1})$ as $h_{i}\to 0$, where
+(8) $E_{i}=\sum_{j=0}^{d}\alpha_{j,i}x(t_{i-d+j})-h_{i}\sum_{j=0}^{d}\beta_{j,i}\frac{\mathrm{d}x}{\mathrm{d}t}(t_{i-d+j}).$
 The stability of a multistep method is to guarantee boundedness of the numerical solution.
@@
 Apply the multistep discretization, we obtain the homogeneous recurrence $\sum_{j=0}^{d}\alpha_{j,i}x_{i-d+j}=0.$ Now, we introduce $X_{i}=(x_{i-d+1},\cdots,x_{i-1},x_{i})^{\mathrm{T}}$ and the companion matrix
-$\mathbf{F}_{i,i-1}=\begin{bmatrix}&1&&\\ &&\ddots&\\ &&&1\\ -\alpha_{0,i}&-\alpha_{1,i}&\cdots&-\alpha_{d-1,i}\end{bmatrix},\quad\mathbf{F}_{i+j,i}\equiv\mathbf{F}_{i+j,i+j-1}\cdots\mathbf{F}_{i+2,i+1}\mathbf{F}_{i+1,i}.$
-Definition 3 (Stability [15][Chapter 3, Definition 5.4]).
-The multistep method in eq.˜8 is called stable if for all $i$ and $j\geq 0$, there exists a constant $C$, such that
-$\left\lVert\mathbf{F}_{i+j,i}\right\rVert\leq C.$ (11)
-For a fixed multistep method, the matrices $\mathbf{F}_{i+j,i}$ in definition˜3 depend on the chosen time grid, because the coefficients $\alpha_{j,i}$ and $\beta_{j,i}$ are functions of the local stepsize ratios.
+(9) $L_{\alpha;i}=\begin{bmatrix}&1&&\\ &&\ddots&\\ &&&1\\ -\alpha_{0,i}&-\alpha_{1,i}&\cdots&-\alpha_{d-1,i}\end{bmatrix},\,F_{\alpha;i+j,i}\equiv L_{\alpha;i+j}\cdots L_{\alpha;i+2}L_{\alpha;i+1}.$
+Definition 2.4 (Stability [17][Chapter 3, Definition 5.4]).
+The multistep method in Eq.˜6 is called stable if for all $i$ and $j\geq 0$, there exists a constant $C$, such that
+(10) $\left\lVert F_{\alpha;i+j,i}\right\rVert\leq C.$
+For a fixed multistep method, the matrices $F_{\alpha;i+j,i}$ in Definition˜2.4 depend on the chosen time grid, because the coefficients $\alpha_{j,i}$ and $\beta_{j,i}$ are functions of the local stepsize ratios.
 As a result, stability can only be verified for each specific grid individually, which becomes cumbersome when the stepsizes vary.
@@
 For simplicity, throughout this paper we assume that the method is stable on the chosen variable stepsize grid.
-2.4 Subspace Perturbation Theory
-In this subsection, we introduce the existing results about subspace perturbation theory, which is the main tool for analysis in this paper.
-We start with some notations: For any given $A\in\mathbf{C}^{n\times n}$ and $B\in\mathbf{C}^{m\times m}$, then:
+3 pPEP Derived by Multistep Discretization of LTP System
+When a multistep method is applied to an LTP system Eq.˜1, it gives rise to a linear difference equation with periodic coefficient matrices, referred to as a periodic difference equation.
+In this section, we will show that this periodic difference equation corresponds to a periodic polynomial eigenvalue problem (pPEP).
+Furthermore, the pPEP admits equivalent representations either as a polynomial eigenvalue problem with block-cyclic coefficient matrices or as a periodic eigenvalue problem with coefficients taking the companion form of the underlying difference equation.
+3.1 Derivation of pPEP and its Linearization
+Consider a partition of one period $[0,T]$ at time points $0=t_{0}<t_{1}<\cdots<t_{p}=T$.
+This grid is extended to an infinite sequence $\left\{t_{i}\right\}_{i\in\mathbb{Z}}$ satisfying $t_{i+p}=t_{i}+T$, which defines a periodic time grid.
+Throughout this paper, quantities defined on this periodic time grid that exhibit $p$-periodicity are indexed with a superscript $(i)$ to indicate their cyclic dependence.
+Accordingly, the multistep coefficients $\left(\alpha_{j,i},\beta_{j,i}\right)$ are denoted by $\left(\alpha_{j}^{(i)},\beta_{j}^{(i)}\right)$, the stepsizes by $h^{(i)}=t_{i}-t_{i-1}$, the maximum step by $h=\max_{i=1}^{p}h^{(i)}$, and the step ratio by $\omega^{(i)}$.
+Applying a linear $d$-step method to Eq.˜1 yields a linear $d$-order difference equation of the form
+(11) $\sum_{j=0}^{d}A_{j}^{(i)}x_{i-d+j}=0,$
+where the coefficient matrices $A_{j}^{(i)}\in\mathbb{C}^{n\times n}$ are $p$-periodic in $i$ and given by
+(12) $A_{j}^{(i)}=\alpha_{j}^{(i)}I_{n}-\beta_{j}^{(i)}G^{(i-d+j)}h^{(i)},\quad j=0,\ldots,d.$
+To compute Floquet multipliers, we seek an eigenvalue formulation associated with Eq.˜11.
+Motivated by the discrete Floquet theorem for one-step periodic recursions [7], we substitute the Floquet-type Ansatz $x_{i}=u^{(i)}\theta^{i}$ into the general $d$-step recursion Eq.˜11, where $u^{(i)}$ is $p$-periodic.
+This leads to a periodic polynomial eigenvalue problem characterizing the multistep dynamics.
+Definition 3.1 (Periodic Polynomial Eigenvalue Problem (pPEP)).
+Let $\{A_{j}^{(i)}\}_{j=0,\ldots,d}^{i=1,\ldots,p}$ denote $d+1$ sequences of matrices in $\mathbb{C}^{n\times n}$, each of period $p$.
+The periodic polynomial eigenvalue problem (pPEP) consists of finding a scalar $\theta\in\mathbb{C}$ and a nonzero $p$-periodic sequence $\{u^{(i)}\}_{i\in\mathbb{Z}}$, $u^{(i)}\subset\mathbb{C}^{n}$ such that
+(13) $\sum_{j=0}^{d}A_{j}^{(i)}\,u^{(i-d+j)}\theta^{j}=0,\qquad i=1,\ldots,p,$
+where the indices $(i)$ in the superscript are understood $i$ modulo $p$.
+Stacking $p$ components in Eq.˜13 together leads to a polynomial eigenvalue problem (PEP) Eq.˜14 of degree $d$ with block-cyclic coefficient matrices to reflect the periodicity.
+To illustrate this, we define the block-diagonal matrices $\mathcal{A}_{j}=\operatorname{diag}_{i}(A_{j}^{(i)}),\,j=0,\ldots,d,$ and the cyclic shift operator $\mathcal{C}=C_{p}\otimes I_{n}$, $C_{p}=[e_{2},\dots,e_{p},e_{1}],$ where $e_{i}$ denotes the $i$-th standard basis vector of $\mathbb{R}^{p}$.
+(14) $\left(\sum_{j=0}^{d}\mathcal{A}_{j}\mathcal{C}^{d-j}\theta^{j}\right)\operatorname{col}(u^{(i)})_{i=1}^{p}=0,\quad\operatorname{col}(u^{(i)})_{i=1}^{p}\equiv\begin{bmatrix}u^{(1)}\\ u^{(2)}\\ \vdots\\ u^{(p)}\end{bmatrix}\in\mathbb{C}^{np},$
+A simple example monic pPEP for $d=2$ and $p=3$ is given by Eq.˜15:
+(15) $\left(I_{3n}\theta^{2}+\begin{bmatrix}&&A_{1}^{(1)}\\ A_{1}^{(2)}&&\\ &A_{1}^{(3)}&\\ \end{bmatrix}\theta+\begin{bmatrix}&A_{0}^{(1)}&\\ &&A_{0}^{(2)}\\ A_{0}^{(3)}&&\\ \end{bmatrix}\right)\begin{bmatrix}u^{(1)}\\ u^{(2)}\\ u^{(3)}\\ \end{bmatrix}=0.$
+If $(\theta,\operatorname{col}\{u^{(i)}\}_{i=1}^{p})$ is an eigenpair, then for each $k=0,\ldots,p-1$,
+$\left(\theta\zeta_{p}^{k},\operatorname{col}\{u^{(i)}\zeta_{p}^{ik-k}\}_{i=1}^{p}\right)$
+is also an eigenpair, where $\zeta_{p}=e^{\jmath 2\pi/p}$ is a primitive $p$-th root of unity, that $\jmath^{2}=-1$.
+Among these, the periodic eigenvalue is defined as $\lambda=\theta^{p}$, and the periodic eigenvector as the sequence $\operatorname{col}\{u^{(i)}\}_{i=1}^{p}$ corresponding to the root $\theta$ with minimal argument $\arg(\theta)\in[0,2\pi/p)$, the periodic eigenpair is referred to as $(\theta^{p},\operatorname{col}\{u^{(i)}\}_{i=1}^{p})$.
+The pPEP is called monic if its leading coefficient matrices satisfy $A_{d}^{(i)}=I_{n}$ for all $i=1,\ldots,p$.
+For pPEPs arising from multistep discretizations of LTP ODE systems, this assumption entails no loss of generality.
+For sufficiently small step sizes $h^{(i)}$, the matrices $A_{d}^{(i)}$ defined in Eq.˜12 are nonsingular.
+Thus, by left-multiplying the $i$-th equation in Eq.˜13 by $\left(A_{d}^{(i)}\right)^{-1}$, the pPEP can always be transformed into an equivalent monic form.
+The linearization of the PEP Eq.˜14 is
+(16) $\begin{bmatrix}&I_{np}&&\\ &&\ddots&\\ &&&I_{np}\\ -\mathcal{A}_{0}\mathcal{C}^{d}&-\mathcal{A}_{1}\mathcal{C}^{d-1}&\cdots&-\mathcal{A}_{d-1}\mathcal{C}\\ \end{bmatrix}\begin{bmatrix}\operatorname{col}(u^{(i)})_{i=1}^{p}\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta\\ \vdots\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta^{d-1}\end{bmatrix}=\begin{bmatrix}\operatorname{col}(u^{(i)})_{i=1}^{p}\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta\\ \vdots\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta^{d-1}\end{bmatrix}\theta.$
+3.2 Linearized pPEP as a Periodic Eigenvalue Problem
+The pPEP Eq.˜14 in PEP form can be transformed into a periodic eigenvalue problem (pEVP) of dimension $nd$ and period $p$.
+The transformation is obtained by a linearization of PEP followed by a permutation.
+The eigenvector in Eq.˜16 exhibits a two-level structure: the inner structure encodes the periodicity of the coefficients, while the outer structure arises from the linearization of Eq.˜14.
+To make this structure explicit, we introduce a permutation matrix $\mathcal{P}$ that reorders the stacked vector:
+$\mathcal{P}\begin{bmatrix}\operatorname{col}(u^{(i)})_{i=1}^{p}\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta\\ \vdots\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta^{d-1}\end{bmatrix}=\begin{bmatrix}\operatorname{col}(u^{(1-j)}\theta^{d-j})_{j=1}^{d}\\ \operatorname{col}(u^{(2-j)}\theta^{d-j})_{j=1}^{d}\\ \vdots\\ \operatorname{col}(u^{(p-j)}\theta^{d-j})_{j=1}^{d}\\ \end{bmatrix},\,\operatorname{col}(u^{(i-j)}\theta^{d-j})_{j=1}^{d}=\begin{bmatrix}u^{(i-d+1)}\\ u^{(i-d+2)}\theta\\ \vdots\\ u^{(i)}\theta^{d-1}\\ \end{bmatrix}$.
+After this permutation, Eq.˜16 takes the following block-cyclic form:
+(17) $\begin{bmatrix}&&&L^{(1)}\\ L^{(2)}&&&\\ &\ddots&&\\ &&L^{(p)}&\\ \end{bmatrix}\begin{bmatrix}\operatorname{col}(u^{(1-j)}\theta^{d-j})_{j=1}^{d}\\ \operatorname{col}(u^{(2-j)}\theta^{d-j})_{j=1}^{d}\\ \vdots\\ \operatorname{col}(u^{(p-j)}\theta^{d-j})_{j=1}^{d}\\ \end{bmatrix}=\begin{bmatrix}\operatorname{col}(u^{(1-j)}\theta^{d-j})_{j=1}^{d}\\ \operatorname{col}(u^{(2-j)}\theta^{d-j})_{j=1}^{d}\\ \vdots\\ \operatorname{col}(u^{(p-j)}\theta^{d-j})_{j=1}^{d}\\ \end{bmatrix}\theta,$
+where each block $L^{(i)}$ is the companion matrix of Eq.˜11 at step $i$:
+(18) $L^{(i)}=\begin{bmatrix}&I_{n}&&\\ &&\ddots&\\ &&&I_{n}\\ -A_{0}^{(i)}&-A_{1}^{(i)}&\cdots&-A_{d-1}^{(i)}\\ \end{bmatrix}$.
+The block-cyclic structure implies that, for every $i=1,\ldots,p$, Eq.˜17 reduces to $p$ periodic eigenvalue problems of dimension $nd$:
+(19) $L^{(i+p)}L^{(i+p-1)}\cdots L^{(i+1)}\operatorname{col}(u^{(i-j)}\theta^{d-j})_{j=1}^{d}=\operatorname{col}(u^{(i-j)}\theta^{d-j})_{j=1}^{d}\theta^{p},\,i=1,\ldots,p.$
+4 Convergence Analysis of Solving Floquet eigenpairs from pPEP
+In this section, we analyze whether and in what sense the periodic eigenpairs obtained from the linearized form of the pPEP Eq.˜19 approximate the Floquet eigenpairs of the underlying LTP system, and characterize the role of the additional eigenvalues introduced by the multistep discretization.
+4.1 Connection between Periodic Eigenpairs of pPEP and Floquet Eigenpairs
+This subsection motivates the subsequent convergence analysis by examining the structural relationship between the linearized pPEP Eq.˜14 and the continuous Floquet eigenvalue problem Eq.˜4.
+Each companion matrix $L^{(i)}$ in Eq.˜18 performs one-step propagation of the $d$-stacked state vectors of Eq.˜11 [16]:
+(20) $\operatorname{col}(x_{j})_{j=i-d+1}^{i}=L^{(i)}\operatorname{col}(x_{j})_{j=i-d}^{i-1}.$
+Consequently, the discrete stacked propagator of Eq.˜11 from step $i$ to $i+j$ is
+(21) $F_{i+j,i}=L^{(i+j)}L^{(i+j-1)}\cdots L^{(i+1)},$
+and in particular $F_{i+p,i}$ represents the one-period discrete stacked propagator.
+Because $L^{(i)}$ is $p$-periodic, $F_{i+p,i}$ for all $i\in\mathbb{Z}$ satisfies similarity relation $F_{i+p,i}=F_{i,0}F_{p,0}F_{i,0}^{-1},$ so it suffices to consider $i=0$.
+The propagation from step 0 to $p$ is given by
+(22) $\operatorname{col}(x_{j})_{j=p-d+1}^{p}=F_{p,0}\,\operatorname{col}(x_{j})_{j=-d+1}^{0}.$
+and the pPEP Eq.˜19 amounts to finding a sequence $x_{i}$ satisfying $x_{i+p}=x_{i}\lambda,$ i.e. $\operatorname{col}(x_{j})_{j=p-d+1}^{p}=\operatorname{col}(x_{j})_{j=-d+1}^{0}\lambda$
+By comparison, the continuous Floquet eigenvalue problem Eq.˜4 seeks eigenpairs $(u(0),e^{\mu T})$.
+The solution $x(t)$ of Eq.˜1 with $x(0)=u(0)$ satisfies $x(T)=e^{\mu T}x(0)$.
+Accordingly, one may associate the Floquet multipliers $e^{\mu T}$ with a subset of eigenvalues of $F_{p,0}$, in the sense that such eigenvalues are expected to approximate $e^{\mu T}$ as the grid is refined.
+Importantly, the $n$ Floquet multipliers and their associated eigenvectors determine only $n$ reference eigenpairs of $F_{p,0}$.
+The remaining (d1)n(d-1)n eigenpairs do not admit a direct continuous-time interpretation and arise from the augmented state space induced by the $d$-step discretization.
+4.2 Convergence Results of Periodic Eigenpairs of pPEP
+The discussion in the previous subsection suggests that only $n$ eigenpairs of the one-period stacked propagator $F_{p,0}$ are relevant for approximating the Floquet multipliers of the underlying LTP system, while the remaining eigenvalues arise from the augmented state space induced by the multistep discretization.
+We now make this correspondence precise by establishing rigorous convergence results for the periodic eigenpairs of the pPEP.
+To ensure that these convergence results are well-defined under time-grid refinement, we impose several standard assumptions on the multistep discretization.
+  1. (A1)
+    Smoothness of Multistep Method Coefficients.
+We assume that the functions $\alpha_{j}(\boldsymbol{\omega})$ and $\beta_{j}(\boldsymbol{\omega})$ are $C^{1}$-continuous with respect to $\boldsymbol{\omega}$ in a neighborhood of the uniform grid ratio $\mathbf{1}=(1,\ldots,1)$.
+  2. (A2)
+    Consistency of the Multistep Method. The $d$-step multistep method has consistency order $s\geq 1$. Thus, the coefficients satisfy $\sum_{j=0}^{d}\alpha_{j}^{(i)}=0$.
+  3. (A3)
+    Strong Stability of the Multistep Method. The underlying constant-stepsize method is strongly stable in the sense of [8]. Specifically, the first generating polynomial
+    $\rho(\nu)=\sum_{j=0}^{d}\alpha_{j}(\mathbf{1})\nu^{j}$
+    has a simple root at $\nu_{1}=1$, referred to as the principal root, while the remaining roots $\nu_{j}$, $j=2,\ldots,d$, referred to as the parasitic roots, satisfy $|\nu_{j}|<1$ [17].
+Moreover, all parasitic roots are assumed to be distinct.
+  4. (A4)
+    Bounded Total Deviation of the Periodic Time Grid.
+The step ratios are uniformly bounded, more over, they satisfy $\sum_{i=0}^{p}\lvert\omega^{(i)}-1\rvert\leq C_{T},$ where $C_{T}>0$ is independent of the time grid refinement.
+The assumptions (A1)–(A4) are standard in the classical theory of variable stepsize multistep methods for initial value problems [17].
+Under these assumptions, the following theorem states the main convergence result for the periodic eigenpairs of the pPEP.
+Theorem 4.1 (Convergence of Floquet Multipliers).
+Assume that the periodic matrix $G(t)$ in Eq.˜1 is continuously differentiable on $[0,T]$, and hence uniformly bounded,
+$\|G(t)\|_{2}\leq C_{G}.$
+All Floquet multipliers $e^{\mu_{k}T}$, $k=1,\ldots,n$, are semisimple and ordered by non-increasing modulus. Suppose that the multistep method satisfies Assumptions (A1)(A4).
+Then, as the maximum stepsize $h\to 0$ (equivalently, as $p\to\infty$), the spectrum of the one-period stacked propagator $F_{p,0}$ admits a decomposition into a principal part and a parasitic part with the following properties.
   1. 1.
-    Spectral gap: $\operatorname{d}(A,B)=\min\{|\lambda-\mu|:\lambda\in\lambda(A),\mu\in\lambda(B)\};$
+    There exist $n$ eigenvalues of $F_{p,0}$, denoted by $\{\lambda_{k}(h)\}_{k=1}^{n}$, such that, after suitable ordering,
+    $\lambda_{k}(h)=e^{\mu_{k}T}+O(h^{s}),\qquad k=1,\ldots,n,$
+    where $\{e^{\mu_{k}T}\}_{k=1}^{n}$ are the Floquet multipliers of the underlying LTP system.
   2. 2.
-    Departure from normality of $A$ with respect to the matrix Frobenius norm:
+    For any fixed $m$, $1\leq m\leq n$, the invariant subspace spanned by the eigenvectors associated with the $m$ dominant eigenvalues $\{\lambda_{k}(h)\}_{k=1}^{m}$, ordered by non-increasing modulus, converges to the subspace spanned by the first $m$ columns of $\mathcal{X}(t_{0})$ at the rate $O(h^{s}/\left\lvert e^{\mu_{m}T}-e^{\mu_{m+1}T}\right\rvert>0)$ if $\left\lvert e^{\mu_{m}T}-e^{\mu_{m+1}T}\right\rvert>0$.
+  3. 3.
+    The remaining (d1)n(d-1)n eigenvalues of $F_{p,0}$, referred to as parasitic periodic eigenvalues, satisfy
+    $|\lambda_{k}(h)|\leq C\,\nu_{\max}^{p},\qquad k=n+1,\ldots,nd,\quad\nu_{\max}=\max_{2\leq j\leq d}|\nu_{j}|<1.$
+    They therefore converge to 0 at a geometric rate as $p\to\infty$.
+4.3 Construction of the Moving Basis and the Reference Transition Operator
+To analyze the variable stepsize case, we construct a moving basis $\mathcal{Z}(t_{i})$ in the augmented state space $\mathbb{R}^{nd}$.
+Let $X(t)=\Phi(t,0)U(0)\in\mathbb{R}^{n\times n}$ represent the principal solution component, where $\Phi(t,0)$ is the state transition matrix of the ODE and $U(0)$ contains the Floquet eigenvectors.
+The parasitic components are modeled explicitly by the roots $\nu_{j}$.
+We define the moving basis matrix $\mathcal{Z}(t_{i})\in\mathbb{R}^{nd\times nd}$ by concatenating the snapshots of the principal solution and the parasitic modes:
+(23) $\mathcal{Z}(t_{i})\equiv$ $\begin{bmatrix}X(t_{i-d+1})&\nu_{2}^{i-d+1}I_{n}&\cdots&\nu_{d}^{i-d+1}I_{n}\\ \vdots&\vdots&\ddots&\vdots\\ X(t_{i-1})&\nu_{2}^{i-1}I_{n}&\cdots&\nu_{d}^{i-1}I_{n}\\ X(t_{i})&\nu_{2}^{i}I_{n}&\cdots&\nu_{d}^{i}I_{n}\end{bmatrix}$
+$=$ $\begin{bmatrix}\mathcal{X}(t_{i})&\mathcal{Y}_{2}(t_{i})&\cdots&&\mathcal{Y}_{d}(t_{i}),\end{bmatrix}$
+where $\mathcal{X}(t_{i})$ is the first block column of $\mathcal{Z}(t_{i})$ and $\mathcal{Y}_{\tau}$ denotes the $\tau$-th block column of $\mathcal{Z}(t_{i})$ that $\tau=2,\ldots,d$.
+The basis $\mathcal{Z}(t_{i})$ captures the state of the multistep method at step $i$, effectively separating the principal subspace from the parasitic subspaces.
+We observe that $\mathcal{Z}(t_{i})$ admits the decomposition:
+(24) $\mathcal{Z}(t_{i})=\underbrace{\begin{bmatrix}\Phi(t_{i-d+1},t_{i})&I_{n}&\cdots&I_{n}\\ \vdots&\vdots&\ddots&\vdots\\ \Phi(t_{i-1},t_{i})&\nu_{2}^{d-1}I_{n}&\cdots&\nu_{d}^{d-1}I_{n}\\ I_{n}&\nu_{2}^{d}I_{n}&\cdots&\nu_{d}^{d}I_{n}\end{bmatrix}}_{\mathcal{V}(t_{i})}\operatorname{diag}\left(X(t_{i}),\,\nu_{2}^{i-d}I_{n},\,\ldots,\,\nu_{d}^{i-d}I_{n}\right).$
+As the stepsize approaches zero, $\Phi(t_{i-j},t_{i})$ approaches to $I_{n}$ with $\mathcal{O}(h)$.
+Consequently, the block matrix $\mathcal{V}(t_{i})$ converges to a block Vandermonde matrix.
+Since the roots $\{1,\nu_{2},\dots,\nu_{d}\}$ are distinct, $\mathcal{V}(t_{i})$ is nonsingular for sufficiently small stepsizes.
+This guarantees that $\mathcal{Z}(t_{i})$ is invertible.
+Given the nonsingular bases at consecutive time steps, the corresponding reference state-transition operator $\Psi(t_{i},t_{i-1})$ is well defined.
+(25) $\Psi(t_{i},t_{i-1})\equiv\mathcal{Z}(t_{i})\,[\mathcal{Z}(t_{i-1})]^{-1}.$
+Noting that $X(t_{i+p})=X(t_{i})e^{MT}$, the moving basis $\mathcal{Z}(t_{i})$ is periodic up to a multiplying a diagonal matrix:
+(26) $\mathcal{Z}(t_{i+p})=\mathcal{Z}(t_{i})\,\begin{bmatrix}e^{MT}&\\ &D^{p}\end{bmatrix},\quad D=\operatorname{diag}\left(\nu_{2}^{p},\ldots,\nu_{d}^{p}\right)\otimes I_{n}.$
+Thus, the reference state-transition operator over one period possesses the eigendecomposition:
+(27) $\Psi(t_{i+p},t_{i})\,\mathcal{Z}(t_{i})=\mathcal{Z}(t_{i})\begin{bmatrix}e^{MT}&\\ &D^{p}\end{bmatrix}$.
+4.4 Convergence Analysis
+The construction of the reference stacked propagator $\Psi(t_{i+p},t_{i})$ establishes the framework for analyzing the discrete stacked propagator $F_{i+p,i}=L^{(i+p)}\cdots L^{(i+1)}$.
+Setting $i=0$ without loss of generality, the analysis proceeds by comparing $F_{p,0}$ against $\Psi(t_{p},t_{0})$ in a three-stage process:
+  1. 1.
+    Local Estimation: Compute the local deviation between the local discrete transition $L^{(i)}$ and its reference $\Psi(t_{i},t_{i-1})$ on the moving basis:
+    (28) $\left[\mathcal{Z}(t_{i})\right]^{-1}L^{(i)}\mathcal{Z}(t_{i-1})=I_{dn}+\Delta_{i}.$
+  2. 2.
+    Global Accumulation: Compute the global deviation between the one-period discrete propagator $F_{p,0}=L^{(p)}\cdots L^{(1)}$ and its reference $\Psi(t_{p},t_{0})$ on the moving basis:
+    $\mathcal{Z}(t_{0})]^{-1}F_{p,0}\mathcal{Z}(t_{0})=\begin{bmatrix}e^{MT}&\\ &D^{p}\end{bmatrix}(I_{dn}+\Delta),$
+    where $\begin{bmatrix}e^{MT}&\\ &D^{p}\end{bmatrix}$ is the eigenvalues of $\Psi(t_{p},t_{0})$ and $\Delta$ represents the accumulated global perturbation.
+  3. 3.
+    Eigenpair Perturbation Estimation: Since the representation of $\Psi(t_{p},t_{0})$ on the moving basis is an eigendecomposition Eq.˜27, then apply the subspace and eigenvalue perturbation theory to estimate how eigenpair of $F_{p,0}$ deviates from $\Psi(t_{p},t_{0})$.
+Lemma 4.2 (Local Difference Estimation).
+Suppose that Assumptions (A1)–(A4) hold.
+Then there exists a threshold $\delta>0$ such that, for any periodic grid satisfying $\max_{i}h^{(i)}\leq\delta$ the local deviation of $L^{(i)}$ from $\Psi(t_{i},t_{i-1})$ admits the representation
+$\Delta_{i}\equiv\mathcal{Z}^{-1}(t_{i})\bigl[L^{(i)}-\Psi(t_{i},t_{i-1})\bigr]\mathcal{Z}(t_{i-1})=\begin{bmatrix}E_{i,\mathcal{X}\mathcal{X}}&E_{i,\mathcal{X}\mathcal{Y}}\\ E_{i,\mathcal{Y}\mathcal{X}}&E_{i,\mathcal{Y}\mathcal{Y}}\end{bmatrix}$,
+where the block partition is consistent with the decomposition of $\mathcal{Z}$ into its principal and parasitic components, with $E_{i,\mathcal{X}\mathcal{X}}\in\mathbb{R}^{n\times n}$ and $E_{i,\mathcal{Y}\mathcal{Y}}\in\mathbb{R}^{(d-1)n\times(d-1)n}.$
+Moreover, there exist constants $C_{1},C_{2},K>0$, independent of $i$, such that
+$\left\|\begin{bmatrix}E_{i,\mathcal{X}\mathcal{X}}\\ E_{i,\mathcal{X}\mathcal{Y}}\end{bmatrix}\right\|_{2}\leq C_{1}\,{h^{(i)}}^{s+1},\qquad\left\|\begin{bmatrix}E_{i,\mathcal{Y}\mathcal{X}}\\ E_{i,\mathcal{Y}\mathcal{Y}}\end{bmatrix}\right\|_{2}\leq C_{2}\,h^{(i)}+K\sum_{l=0}^{d}\bigl\lvert\omega^{(i-d+l)}-1\bigr\rvert$.
+Proof 4.3.
+By Assumption (A4), the step ratios $\omega^{(i)}$ remain uniformly bounded.
+Together with the smooth dependence of the coefficients $\alpha_{j}(\boldsymbol{\omega})$ and $\beta_{j}(\boldsymbol{\omega})$ Assumption (A1), this implies uniform boundedness of $\alpha_{j}^{(i)}$ and $\beta_{j}^{(i)}$.
+Hence, there exists $\delta>0$ sufficiently small such that for $h^{(i)}<\delta$, the matrix $\alpha_{d}^{(i)}I_{n}-h^{(i)}\beta_{d}^{(i)}G^{(i)}$ is invertible with uniformly bounded inverse.
+We first examine the action of $L^{(i)}-\Psi(t_{i},t_{i-1})$ on each block column of $\mathcal{Z}(t_{i-1})$.
+$\left[L^{(i)}-\Psi(t_{i},t_{i-1})\right]\mathcal{Z}(t_{i-1})$
+$=$ $\left[L^{(i)}-\Psi(t_{i},t_{i-1})\right]\left[\mathcal{X}_{1}(t_{i-1}),\,\mathcal{Y}_{2}(t_{i-1}),\,\ldots,\,\mathcal{Y}_{d}(t_{i-1})\right]$
+$=$ $\begin{bmatrix}0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\\ R_{\mathcal{X}_{i}}&R_{\mathcal{Y}_{2,i}}\nu_{2}^{i-d}&\cdots&R_{\mathcal{Y}_{d,i}}\nu_{d}^{i-d}\end{bmatrix}$.
+By Assumption (A2) and the uniform boundedness of the inverse matrix,
+$\left\lVert R_{\mathcal{X}_{i}}\right\rVert_{2}\leq C_{1}\,{h^{(i)}}^{s+1}.$
+For $\tau=2,\ldots,d$, the residual takes the form
+$R_{\mathcal{Y}_{\tau,i}}=-\left(\alpha_{d}^{(i)}I_{n}-h^{(i)}\beta_{d}^{(i)}G^{(i)}\right)^{-1}\left(\sum_{j=0}^{d}h^{(i)}\beta_{j}^{(i)}G^{(i-d+j)}\nu_{\tau}^{j}+\sum_{j=0}^{d}\alpha_{j}^{(i)}\nu_{\tau}^{j}I_{n}\right).$
+The first term inside the right bracket is $O(h^{(i)})$ by boundedness of the coefficients $\beta_{j}^{(i)}$ and $G(t)$.
+For the second term, using the parasitic root condition $\sum_{j}\alpha_{j}(\mathbf{1})\nu_{\tau}^{j}=0$ and the smoothness assumption on the coefficients (A1), we obtain
+$\left\lvert\sum_{j=0}^{d}\alpha_{j}^{(i)}\nu_{\tau}^{j}\right\rvert=\left\lvert\sum_{j=0}^{d}\left(\alpha_{j}^{(i)}-\alpha_{j}(\boldsymbol{1})\right)\nu_{\tau}^{j}\right\rvert\leq K\sum_{l=0}^{d}|\omega^{(i-d+l)}-1|.$
+Consequently,
+$\left\lVert R_{\mathcal{Y}_{\tau,i}}\right\rVert_{2}\leq C_{2}h^{(i)}+K\sum_{l=0}^{d}|\omega^{(i-d+l)}-1|.$
+We now transform the residual matrix to the moving basis by left-multiplying with $\mathcal{Z}^{-1}(t_{i})$. Therefore,
+$\Delta_{i}$ $=\mathcal{Z}^{-1}(t_{i})\left[L^{(i)}\mathcal{Z}(t_{i-1})-\mathcal{Z}(t_{i})\right]$
+$=\begin{bmatrix}X(t_{i-d+1})&I_{n}&\cdots&I_{n}\\ \vdots&\vdots&\ddots&\vdots\\ X(t_{i-1})&I_{n}\nu_{2}^{d-1}&\cdots&I_{n}\nu_{d}^{d-1}\\ X(t_{i})&I_{n}\nu_{2}^{d}&\cdots&I_{n}\nu_{d}^{d}\end{bmatrix}^{-1}\begin{bmatrix}0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\\ R_{\mathcal{X}_{i}}&R_{\mathcal{Y}_{2,i}}&\cdots&R_{\mathcal{Y}_{d,i}}\end{bmatrix}$
+inherits the same block-wise bounds as the residual matrix.
+Finally, transforming to the $\mathcal{Z}$-coordinates via $\Delta_{i}=\mathcal{Z}^{-1}(t_{i})R_{i}$ and using the uniform boundedness of $\mathcal{Z}^{-1}(t_{i})$ yields the stated block-wise estimates.
+Lemma 4.4 (Global Difference Estimation).
+Suppose that Assumptions (A1)–(A4) hold. Then, with respect to the moving basis $\mathcal{Z}(t_{0})$, the discrete stacked propagator over one period $F_{p,0}$ admits the representation
 […]
-  3. 3.
-    The separation of $A$ and $B$ with respect to the matrix $\mathrm{F}$-norm:
-    $\operatorname{sep}_{\mathrm{F}}(A,B)=\inf_{\left\lVert P\right\rVert=1}\left\lVert PA-BP\right\rVert_{\mathrm{F}}.$
-Directly computing the separation is hard, but it can be bounded by the spectral gap and departure from normality. In theorem˜1, the upper bound is given in [17], and the lower bound is given by [18].
-Theorem 1 (Bounds of $\operatorname{sep}(A,B)$ [18, 17]).
-Let $A\in\mathbf{C}^{n\times n}$ and $B\in\mathbf{C}^{m\times m}$ with $\lambda(A)\cap\lambda(B)=\varnothing$. Then we have
-$\operatorname{d}(A,B)\geqslant\operatorname{sep}_{\mathrm{F}}(A,B)\geqslant\frac{\operatorname{d}(A,B)}{1+\sum_{j=1}^{dn-2}\left[\frac{\Delta_{\mathrm{F}}(A)+\Delta_{\mathrm{F}}(B)}{\operatorname{d}(A,B)}\right]^{j}}.$
-If $A$ and $B$ are both normal, then $\Delta(A,B)=0$, and we obtain $\operatorname{d}(A,B)=\operatorname{sep}_{\mathrm{F}}(A,B)$.
-The second one is a perturbation theorem from [19, Theorem 3.5], and Betcke and Kressner gave a new proof in [20].
-Theorem 2 (Invariant Subspace Perturbation Theorem [19]).
+where the block structure is consistent with the decomposition of $\mathcal{Z}(t_{0})$ into principal and parasitic components.
+Moreover, under Assumptions (A1)–(A4), there exist constants $C_{1},C_{2}>0$, independent of the periodic time grid, such that
+$\left\lVert\begin{bmatrix}E_{\mathcal{X}\mathcal{X}}\\ E_{\mathcal{X}\mathcal{Y}}\end{bmatrix}\right\rVert_{2}\leq C_{1}h^{s},\qquad\left\lVert\begin{bmatrix}E_{\mathcal{Y}\mathcal{X}}\\ E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}\right\rVert_{2}\leq C_{2}.$
+Proof 4.5.
+We begin by expressing $F_{p,0}$ in the basis $\mathcal{Z}(t_{0})$.
+$\begin{aligned} &\mathcal{Z}^{-1}(t_{0})F_{p,0}\mathcal{Z}(t_{0})\\ =&\mathcal{Z}^{-1}(t_{0})\mathcal{Z}(t_{p})\left[\mathcal{Z}^{-1}(t_{p})L^{(p)}\mathcal{Z}(t_{p-1})\right]\cdots\left[\mathcal{Z}^{-1}(t_{1})L^{(1)}\mathcal{Z}(t_{0})\right]\\ =&\mathcal{Z}^{-1}(t_{0})\mathcal{Z}(t_{p})\left(I_{dn}+\Delta_{p}\right)\cdots\left(I_{dn}+\Delta_{1}\right)\\ =&\begin{bmatrix}e^{MT}&\\ &D^{p}\end{bmatrix}\left(I_{dn}+\Delta_{p}\right)\cdots\left(I_{dn}+\Delta_{1}\right)\end{aligned}.$
+We then define the global error matrix and apply a telescoping sum expansion:
+$\begin{bmatrix}E_{\mathcal{X}\mathcal{X}}&E_{\mathcal{X}\mathcal{Y}}\\ E_{\mathcal{Y}\mathcal{X}}&E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}\equiv$ $\left(I_{dn}+\Delta_{p}\right)\cdots\left(I_{dn}+\Delta_{1}\right)-I_{dn}$
+$=$ $\sum_{i=0}^{p-1}(I_{dn}+\Delta_{p})\cdots(I_{dn}+\Delta_{p-i+1})\Delta_{p-i}.$
+We now estimate the two block columns separately. For the first block column, the telescoping representation yields
+$\begin{bmatrix}E_{\mathcal{X}\mathcal{X}}\\ E_{\mathcal{Y}\mathcal{X}}\end{bmatrix}$= $\sum_{i=0}^{p-1}(I_{dn}+\Delta_{p})\cdots(I_{dn}+\Delta_{p-i})\begin{bmatrix}E_{p-i,\mathcal{X}\mathcal{X}}\\ E_{p-i,\mathcal{Y}\mathcal{X}}\end{bmatrix}$.
+By Lemma˜4.2, there exists a constant $C>0$ such that
+$\left\lVert\left(I_{dn}+\Delta_{p}\right)\cdots\left(I_{dn}+\Delta_{p-i}\right)\right\rVert_{2}\leq\prod_{k=p-i}^{p}\left\lVert I_{dn}+\Delta_{k}\right\rVert_{2}\leq\prod_{k=p-i}^{p}(1+Ch^{(k)})\leq e^{CT},$
+which is uniformly bounded for all $i=1,\ldots,p$.
+Therefore, we obtain the estimate
+$\left\lVert\begin{bmatrix}E_{\mathcal{X}\mathcal{X}}\\ E_{\mathcal{Y}\mathcal{X}}\end{bmatrix}\right\rVert_{2}\leq$ $\sum_{i=1}^{p}e^{CT}\left\lVert\begin{bmatrix}E_{p-i,\mathcal{X}\mathcal{X}}\\ E_{p-i,\mathcal{Y}\mathcal{X}}\end{bmatrix}\right\rVert_{2}\leq e^{CT}\sum_{i=0}^{p-1}(h^{(p-k)})^{s+1}\leq e^{CT}h^{s}$
+where we used $\sum_{i=1}^{p}h^{(i)}=T$ and $h=\max_{i}h^{(i)}$.
+An analogous argument applies to the second block column,
+$\left\lVert\begin{bmatrix}E_{\mathcal{X}\mathcal{Y}}\\ E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}\right\rVert_{2}\leq$ $\sum_{i=1}^{p}e^{CT}\left\lVert\begin{bmatrix}E_{p-i,\mathcal{X}\mathcal{Y}}\\ E_{p-i,\mathcal{Y}\mathcal{Y}}\end{bmatrix}\right\rVert_{2}$
+[…]
+[…]
+By Assumption (A4), the cumulative deviation of the step ratios is uniformly bounded, and therefore the right-hand side is bounded by a constant. This completes the proof.
+The global difference estimates established above quantify the deviation between the one-period discrete propagator $F_{p,0}$ and its reference counterpart constructed $\Psi(t_{p},t_{0})$.
+To link it to estimates of eigenpairs difference, we recall classical results from invariant subspace perturbation theorem [22, 3].
+Theorem 4.6 (Invariant Subspace Perturbation Theorem [22]).
 Let $A,E\in\mathbb{C}^{n\times n}$, $P=[P_{1},P_{2}]$ is a unitary matrix.
@@
 $P^{\mathrm{H}}AP=\begin{bmatrix}A_{11}&A_{12}\\ &A_{22}\end{bmatrix},\quad P^{\mathrm{H}}EP=\begin{bmatrix}E_{11}&E_{12}\\ E_{21}&E_{22}\end{bmatrix}$
-If $\delta\equiv\operatorname{sep}(A_{11},A_{22})-\left(\left\lVert E_{11}\right\rVert+\left\lVert E_{22}\right\rVert\right)>0$, and $\left\lVert E_{21}\right\rVert\left(\left\lVert A_{12}\right\rVert+\left\lVert E_{12}\right\rVert\right)<\frac{1}{4}\delta^{2}$, then there exist a matrix $R\in\mathbb{C}^{(n-l)\times l}$ satisfying $\left\lVert R\right\rVert\leq 2\delta^{-1}\left\lVert E_{21}\right\rVert$, such that $\operatorname{span}\left\{P_{1}\begin{bmatrix}I\\ R\end{bmatrix}\right\}$ is an invariant subspace of $A+E$.
-3 Periodic-PEP and Discrete State Transition Matrix
-When a multistep method is applied to an LPTV system eq.˜1, it gives rise to a linear difference equation with periodic coefficient matrices, referred to as a periodic difference equation.
-In this section, we will show that this periodic difference equation corresponds to a periodic polynomial eigenvalue problem (periodic-PEP).
-Furthermore, the periodic-PEP can be transformed into a periodic eigenvalue problem whose coefficient matrix coincides with the discrete STM over a period, capturing the dynamics of the periodic difference equation.
-3.1 Derivation of the Periodic-PEP
-As in discussion of one-step method, we consider the period $[0,T]$ is partitioned at time points $0=t_{0}<t_{1}<\cdots<t_{p}=T$, In fact, this grid can be extended to an infinite sequence $\left\{t_{i}\right\}_{i\in\mathbb{Z}}$ satisfying $t_{i+p}=t_{i}+T$.
-This extension defines a periodic time grid.
-Throughout this paper, quantities that exhibit periodicity with period $p$ on this periodic discrete grid are indexed with a superscript $(i)$ to indicate their cyclic dependence.
-Hence, the coefficients $(\alpha_{j,i},\beta_{j,i})$ are denoted by $(\alpha_{j}^{(i)},\beta_{j}^{(i)})$, and the step size $\Delta t^{(i)}=t_{i}-t_{i-1}$, and the maximum step $\Delta t=\max_{i=1}^{p}\Delta t^{(i)}$.
-$\sum_{j=0}^{d}\left(\alpha_{j}^{(i)}-\Delta t^{(i)}\beta_{j}^{(i)}G^{(i-d+j)}\right)x_{i-d+j}=0,$ (12)
-Since $\left(I_{n}-\beta_{d}^{(i)}G^{(i)}\Delta t^{(i)}\right)$ is invertible for sufficiently small step size $\Delta t^{(i)}$, eq.˜12 can be expressed in the form of a $d$-step difference equation with $p$-periodic coefficients:
-$x_{i}-\sum_{j=0}^{d-1}A_{j}^{(i)}x_{i-d+j}=0,\quad i=1,\ldots,p,$ (13)
-where the coefficient matrices $A_{j}^{(i)}$ for $j=0,\ldots,d-1$ are defined by
-$A_{j}^{(i)}=-\left(I_{n}-\beta_{d}^{(i)}G^{(i)}\Delta t^{(i)}\right)^{-1}\left(\alpha_{j}^{(i)}I_{n}-\beta_{j}^{(i)}G^{(i-d+j)}\Delta t^{(i)}\right).$ (14)
-In this section, we only use the cyclicity of $A_{j}^{(i)}$ and do not require any of its specific properties from the multistep method.
-Next, we define the periodic-PEP corresponding to eq.˜13.
-Definition 4 (Periodic Polynomial Eigenvalue Problem).
-Let $\{A_{0}^{(i)},A_{1}^{(i)},\ldots,A_{d-1}^{(i)}\}_{i=1}^{p}$ be $p$ sets of coefficient matrices in $\mathbb{R}^{n\times n}$.
-The periodic polynomial eigenvalue problem (periodic-PEP) seeks scalars $\theta\in\mathbb{C}$ and nonzero vectors $u^{(i)}\in\mathbb{C}^{n}$, $i=1,\ldots,p$, satisfying the periodicity condition $u^{(i+p)}=u^{(i)}$ such that
-$u^{(i)}\theta^{d}-\sum_{j=0}^{d-1}A_{j}^{(i)}\,u^{(i-d+j)}\theta^{j}=0,\qquad i=1,\ldots,p.$ (15)
-If $(\theta,\{u^{(i)}\}_{i=1}^{p})$ is an eigenpair, then for each $k=0,\ldots,p-1$, $\left(\theta\zeta_{p}^{k},\,\{u^{(i)}\zeta_{p}^{ik-k}\}_{i=1}^{p}\right)$ is also an eigenpair, where $\zeta_{p}=e^{\jmath 2\pi/p}$ is a primitive $p$th root of unity.
-Among them, we define the periodic eigenvalue of the periodic-PEP as $\lambda=\theta^{p}$, and the periodic eigenvector as the sequence $\{u^{(i)}\}_{i=1}^{p}$ associated with the root $\theta$ whose argument $\arg(\theta)\in[0,2\pi/p)$ is minimal.
-For $d=1$, the periodic-PEP reduces to the standard periodic eigenvalue problem discussed in [11].
-For $d>1$, the problem takes a polynomial form, which we now illustrate.
-As an example, when $d=2$, and $p=3$, definition˜4 can be written as:
-$\left(\theta^{2}I_{3n}-\begin{bmatrix}&&A_{1}^{(1)}\\ A_{1}^{(2)}&&\\ &A_{1}^{(3)}&\\ \end{bmatrix}\theta-\begin{bmatrix}&A_{0}^{(1)}&\\ &&A_{0}^{(2)}\\ A_{0}^{(3)}&&\\ \end{bmatrix}\right)\begin{bmatrix}u^{(1)}\\ u^{(2)}\\ u^{(3)}\\ \end{bmatrix}=0.$ (16)
-This example shows that stacking the $p$ component equations in definition˜4 naturally yields a PEP whose coefficient matrices possess a block-cyclic structure.
-To express this structure compactly, we introduce the stacking notation
-$\operatorname{col}(x_{i})_{i=m+1}^{m+l}=\left[x_{m+1}^{\mathrm{H}},\,x_{m+2}^{\mathrm{H}},\,\ldots,\,x_{m+l}^{\mathrm{H}}\right]^{\mathrm{H}}\in\mathbb{R}^{ln},$ (17)
-which collects lm+1l-m+1 consecutive state vectors ending at step $i$.
-We then define the block-diagonal matrices $\mathcal{A}_{j}=\operatorname{diag}_{i}(A_{j}^{(i)}),\,j=0,\ldots,d-1,$ and the cyclic shift operator $\mathcal{C}=C_{p}\otimes I_{n}$, $C_{p}=[e_{2},\dots,e_{p},e_{1}],$ where $e_{i}$ denotes the $i$-th standard basis vector of $\mathbb{R}^{p}$.
-With these notations, the periodic-PEP in definition˜4 can be equivalently written in the compact matrix form.
-$\left(\theta^{d}I_{np}-\sum_{j=0}^{d-1}\mathcal{A}_{j}\mathcal{C}^{d-j}\theta^{j}\right)\operatorname{col}(u^{(i)})_{i=1}^{p}=0.$ (18)
-3.2 Equivalent Form of the Periodic-PEP using Discrete State Transition Matrix
-This subsection transforms the periodic-PEP in eq.˜18 into an equivalent form to reveal its connection with the general solution of the underlying difference equation.
-We show that through a suitable linearization and a structured reordering of the companion-form matrices, eq.˜18 becomes an $nd$-dimensional periodic eigenvalue problem.
-The coefficient matrices of this new problem are identical to the state transition matrices of the difference equation eq.˜13.
-This identity establishes a direct correspondence between the eigenpairs of the periodic-PEP and the fundamental solutions of the periodic difference system.
-The companion form of the PEP eq.˜18 is:
-$\begin{bmatrix}&I_{np}&&\\ &&\ddots&\\ &&&I_{np}\\ \mathcal{A}_{0}\mathcal{C}^{d}&\mathcal{A}_{1}\mathcal{C}^{d-1}&\cdots&\mathcal{A}_{d-1}\mathcal{C}\\ \end{bmatrix}\begin{bmatrix}\operatorname{col}(u^{(i)})_{i=1}^{p}\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta\\ \vdots\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta^{d-1}\end{bmatrix}=\begin{bmatrix}\operatorname{col}(u^{(i)})_{i=1}^{p}\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta\\ \vdots\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta^{d-1}\end{bmatrix}\theta,$ (19)
-The eigenvector in eq.˜19 exhibits a two-level structure: the inner structure encodes the periodicity of the coefficients, while the outer structure arises from the linearization of eq.˜18.
-To make this structure explicit, we introduce a permutation matrix $\mathcal{P}$ that reorders the stacked vector:
-$\mathcal{P}\begin{bmatrix}\operatorname{col}(u^{(i)})_{i=1}^{p}\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta\\ \vdots\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta^{d-1}\end{bmatrix}=\begin{bmatrix}\operatorname{col}(u^{(1-j)}\theta^{d-j})_{j=1}^{d}\\ \operatorname{col}(u^{(2-j)}\theta^{d-j})_{j=1}^{d}\\ \vdots\\ \operatorname{col}(u^{(p-j)}\theta^{d-j})_{j=1}^{d}\\ \end{bmatrix},\,\operatorname{col}(u^{(i-j)}\theta^{d-j})_{j=1}^{d}=\begin{bmatrix}u^{(i-d+1)}\\ u^{(i-d+2)}\theta\\ \vdots\\ u^{(i)}\theta^{d-1}\\ \end{bmatrix}$.
-After this reordering, the eigenvalue problem takes the following block-cyclic form:
-$\begin{bmatrix}&&&L^{(1)}\\ L^{(2)}&&&\\ &\ddots&&\\ &&L^{(p)}&\\ \end{bmatrix}\begin{bmatrix}\operatorname{col}(u^{(1-j)}\theta^{d-j})_{j=1}^{d}\\ \operatorname{col}(u^{(2-j)}\theta^{d-j})_{j=1}^{d}\\ \vdots\\ \operatorname{col}(u^{(p-j)}\theta^{d-j})_{j=1}^{d}\\ \end{bmatrix}=\begin{bmatrix}\operatorname{col}(u^{(1-j)}\theta^{d-j})_{j=1}^{d}\\ \operatorname{col}(u^{(2-j)}\theta^{d-j})_{j=1}^{d}\\ \vdots\\ \operatorname{col}(u^{(p-j)}\theta^{d-j})_{j=1}^{d}\\ \end{bmatrix}\theta,$ (20)
-where each block $L^{(i)}$ is the companion matrix
-$L^{(i)}=\begin{bmatrix}&I_{n}&&\\ &&\ddots&\\ &&&I_{n}\\ A_{0}^{(i)}&A_{1}^{(i)}&\cdots&A_{d-1}^{(i)}\\ \end{bmatrix}$. (21)
-The block-cyclic structure implies that, for every $i=1,\ldots,p$, the eigenvalue problem reduces to the $nd$-dimensional periodic eigenvalue problem:
-$L^{(i)}L^{(i-1)}\cdots L^{(i-p+1)}\operatorname{col}(u^{(i-j)}\theta^{d-j})_{j=1}^{d}=\operatorname{col}(u^{(i-j)}\theta^{d-j})_{j=1}^{d}\theta^{p}.$ (22)
-This transformation maps each $p$-fold set of eigenvalues of definition˜4 to a single eigenvalue $\theta^{p}$, effectively removing the rotational symmetry and revealing the essential dynamics.
-Recall eq.˜12, we observe that each companion matrix $L^{(i)}$ describes one-step propagation of the stacked state vectors:
-$\operatorname{col}(x_{j})_{j=i-d+1}^{i}=L^{(i)}\operatorname{col}(x_{j})_{j=i-d}^{i-1}.$ (23)
-To describe the evolution, we define the discrete STM in a $dn$-dimensional state space [21]:
-$F_{i+j,i}=L^{(i+j)}L^{(i+j-1)}\cdots L^{(i+1)}.$ (24)
-Because the coefficients $L^{(i)}$ are $p$-periodic, we have $F_{i+kp,i+(k-1)p}=F_{i+p,i}$ for all integers $k$.
-Therefore, we partition the stacked sequence $\operatorname{col}(x_{j})_{j=i-d+1}^{i}$ for all $i\in\mathbb{Z}$ into $p$ groups according to the remainder of the index $i$ modulo $p$.
-Each group forms a phase slice $\operatorname{col}(x_{j})_{j=i+kp-d+1}^{i+kp}$, and within each slice the evolution satisfies a constant-coefficient difference equation:
-$\operatorname{col}(x_{j})_{j=i+kp-d+1}^{i+kp}=F_{i+p,i}\,\operatorname{col}(x_{j})_{j=i+(k-1)p-d+1}^{i+(k-1)p}.$ (25)
-Here, $F_{i+p,i}$ acts as the discrete state-transition matrix over one period for the corresponding slice, and the associated eigenvalue problem is precisely eq.˜22.
-Moreover, these phase slices are not independent, because the transition matrices for different starting points are all similar:
-$F_{i+p,i}=F_{i,0}F_{p,0}F_{i,0}^{-1},$ (26)
-and therefore share the same spectrum. As a result, the eigen-decomposition of $F_{p,0}$ and state transition $F_{i,0}$ fully determines the eigen-structure of all $F_{i+p,i}$.
-In this section, we have shown the discrete STM over a period $F_{i+p,i}$ reveals structural properties that closely mirror those of the continuous STM $\Phi(T,0)$ discussed in section˜2.1.
-The key distinction is that the discrete STM acts on an $nd$-dimensional stacked state corresponding to the difference equation, whereas the continuous STM evolves in the original $n$-dimensional state space of the LPTV system.
-When the coefficient matrix $A_{j}^{(i)}$ are from multistep discretization of the LPTV system, we expect $n$ of the eigenvalues of $F_{p,0}$ to approximate the $n$ Floquet multipliers of the continuous system.
-However, the remaining n(d1)n(d-1) eigenvalues of $F_{p,0}$ require further explanation, which will be addressed by introducing spurious systems in the following section.
-4 Spurious Systems from Multistep Methods
-The continuous LPTV system eq.˜1 evolves in an $n$-dimensional state space.
-Discretizing it with a $d$-step multistep method yields a difference method whose state space is $nd$-dimensional and one-period STM is $F_{p,0}$.
-In this section, based on the given LPTV system and the adopted $d$-step multistep method, we design d1d-1 additional $n$-dimensional spurious systems together with their corresponding multistep discretizations.
-The discretizations of the physical and spurious systems share identical coefficient matrices, hence yield the same difference equation in coefficient form, but differ in the way the state variables enter the recursion.
-Utilizing the Floquet multipliers and eigenvectors of these spurious systems, we construct an $nd$-dimensional auxiliary STM $\Psi(t_{p},t_{0})$ that serves as the limiting counterpart of $F_{p,0}$ as $\Delta t\to 0$.
-4.1 Spurious Systems from Discretizing Scalar System
-To illustrate how multistep methods introduce the spurious systems, we first analyze a simple scalar system, $\dot{y}=0$, $t\in[0,T]$.
-Applying the $d$-step discretization on a periodic time grid leads to a high-order difference equation:
-$y_{i}=-\sum_{j=0}^{d-1}\alpha_{j}^{(i)}y_{i-d+j},\quad i=1,\ldots,p.$ (27)
-It corresponds to the following periodic-PEP:
-$\left(\nu^{d}I_{p}-\sum_{j=0}^{d-1}\mathbf{A}_{j}C^{d-j}\nu^{j}\right)\eta=0,$ (28)
-where $\mathbf{A}_{j}=-\operatorname{diag}\left\{\alpha_{j}^{(i)}\right\}_{i=1}^{p}$, and $\eta=[\eta^{(1)},\ldots,\eta^{(p)}]^{\mathrm{H}}$.
-According to definition˜4, the $d$ periodic eigenvalues of eq.˜28 are denoted by $\nu_{1}^{p},\ldots,\nu_{d}^{p}$.
-According to definition˜3, the trivial eigenvalue $\nu_{1}=1$ remains unchanged, while all other eigenvalues satisfy $\lvert\nu_{\tau}\rvert<1$ for $\tau=2,\ldots,d$.
-The corresponding eigenvector of $\nu_{1}$ is still $1,1,\ldots,1]^{\mathrm{H}}$, whereas the other eigenvectors are no longer uniform when the stepsize is not constant.
-4.2 Spurious Systems from Discretizing LPTV Systems
-Returning now to the general system eq.˜13, each coefficient matrix in eq.˜14 is a small perturbation of $\mathbf{A}_{j}\otimes I_{n}$ when $\Delta t$ is sufficiently small.
-Motivated by the scalar case, we assume that the general solution of eq.˜13 inherits the same structure, taking the form $x_{\tau,i}\eta^{(i)}\nu^{i}$.
-Substituting this back to eq.˜13 yields:
-$x_{\tau,i}\eta_{\tau}^{(i)}\nu_{\tau}^{i}-\sum_{j=1}^{d}A_{j}^{(i)}x_{\tau,i-d+j}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{i-d+j}=0,\quad\tau=1,\ldots,d.$ (29)
-The mode $\nu_{1}=1$ with $\eta_{1}^{(i)}\equiv 1$ corresponds to the physical dynamics and recovers eq.˜13.
-The remaining modes $\nu_{\tau}$ $(\tau>1)$ also describe valid dynamical components through $x_{\tau,i}$; however, their associated decaying envelopes render them spurious in the context of the original system.
-To formalize how $x_{\tau,i}$ captures these spurious dynamics, we associate each mode $\tau$ with a corresponding spurious LPTV system and a spurious multistep method.
-In doing so, we show that eq.˜29 can be obtained from the discretization of these spurious systems.
-This motivates the construction of the corresponding spurious multistep methods, whose coefficients are then defined as
-$\alpha_{\tau,j}^{(i)}=\frac{\alpha_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}}{\sum_{j=0}^{d}\alpha_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}h_{j}^{(i)}},\quad\beta_{\tau,j}=\frac{\beta_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}}{\sum_{j=0}^{d}\beta_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}}.$ (30)
-where the envelope terms are incorporated into the numerators, and the normalization ensures the first-order consistency condition (definition˜2).
-The scaling factor $h_{j}^{(i)}$ in the denominator of $\alpha_{\tau,j}^{(i)}$ is defined by
-$h_{j}^{(i)}\equiv\frac{t_{i+j}-t_{i}}{t_{i+d}-t_{i+d-1}}=\frac{\Delta t^{(i+1)}}{\Delta t^{(i+d)}}\sum_{l=1}^{j}\frac{\Delta t^{(i+l)}}{\Delta t^{(i+1)}}=\left(\prod_{l=1}^{d}{\omega^{(i+l)}}^{-1}\right)\sum_{l=1}^{j}\omega^{(i+l)}\cdots\omega^{(i+1)},$ (31)
-where $\omega^{(i)}=\Delta t^{(i)}/\Delta t^{(i-1)}$ denotes the local stepsize ratio.
-As seen from eq.˜31, $h_{j}^{(i)}$ depends only on the stepsize ratios $\omega^{(i+l)}$ for $l=1,\ldots,d$, and reduces to $h_{j}^{(i)}=j$ under constant stepsizes.
-We now construct the corresponding spurious LPTV systems:
-$\frac{\mathrm{d}x_{\tau}}{\mathrm{d}t}$ $=\rho_{\tau}(t)G(t)x_{\tau}(t),$ (32)
-$\rho_{\tau}(t)$ $=\frac{\sum_{j=0}^{d}\beta_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}}{\sum_{j=0}^{d-1}\alpha_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}h_{j}^{(i)}},\quad t\in[t_{i-1},t_{i}],\,\tau=2,\ldots,d.$
-Although the underlying system is locally defined on each interval $t_{i-1},t_{i}]$ with potentially distinct dynamics, the $d$-step multistep discretization requires evaluating the equation over the combined interval $t_{i-d},t_{i}]$.
-To this end, we formally extend the equation form from the last interval $t_{i-1},t_{i}]$ to the preceding intervals $t_{i-d},t_{i-1}]$, effectively using the definition on $t_{i-1},t_{i}]$ to cover the earlier steps.
-Applying the spurious multistep coefficients from eq.˜30 to this extended system then yields
-$\sum_{j=0}^{d}\left(\alpha_{j}^{(i)}I_{n}-\Delta t^{(i)}\beta_{j}^{(i)}G^{(i-d+j)}\right)x_{\tau,i-d+j}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}=0,\quad\tau=1,\ldots,d.$ (33)
-Left-multiplying by $\left(I_{n}-\Delta t^{(i)}\beta_{d}^{(i)}G^{(i)}\right)^{-1}$ then recovers eq.˜29, and we can write eq.˜23 in an extended version
-$L^{(i)}\operatorname{col}\left(x_{\tau,j}\,\eta^{(j)}\,\nu^{\,j}\right)_{j=i-d}^{i-1}=\operatorname{col}\left(x_{\tau,j}\,\eta^{(j)}\,\nu^{\,j}\right)_{j=i-d+1}^{i}.$ (34)
-At last, we estimate the local truncation error of the spurious multistep method. Accroding to eq.˜10, we have the following lemma.
-Lemma 3.
-Consider the spurious $d$-step discretization eq.˜33 applied to the spurious system.
-When the stepsize is constant, the local truncation error of eq.˜33 is second-order, but its local truncation error drops to first order when stepsize is variable.
-4.3 Construction of Auxiliary State Transition Matrix
-Similar to the structure of $F_{i+1,i}=L^{(i)}$, we now consider a continuous representation.
-Let $X_{\tau}(t)$ denote a fundamental solution matrix of eq.˜32 for any fixed full rank initial matrix condition, and we define an $n$-dimensional subspace of an $nd$-dimensional space spanned by:
-$\mathcal{X}_{\tau}\equiv\operatorname{span}\left\{\operatorname{col}\left(X_{\tau}^{(i)}(t_{j})\eta_{\tau}^{(j)}\nu_{\tau}^{j}\right)_{j=i-d+1}^{i}\right\}.$
-Next, we show that the subspaces $\mathcal{X}_{1},\ldots,\mathcal{X}_{d}$ form a direct sum spanning the entire $nd$-dimensional space, that is $\mathcal{X}_{1}\oplus\mathcal{X}_{2}\oplus\cdots\oplus\mathcal{X}_{d}=\mathbb{R}^{(nd)\times(nd)}.$ To see this, we collect the bases of all subspaces into the block matrix
-$\begin{bmatrix}X_{1}(t_{i-d+1})&X_{2}(t_{i-d+1})\eta_{2}^{(i-d+1)}&\cdots&X_{d}(t_{i-d+1})\eta_{d}^{(i-d+1)}\\ \vdots&\vdots&\ddots&\vdots\\ X_{1}(t_{i-1})&X_{2}(t_{i-1})\eta_{2}^{(i-1)}\nu_{2}^{i-1}&\cdots&X_{d}(t_{i-1})\eta_{d}^{(i-1)}\nu_{d}^{i-1}\\ X_{1}(t_{i})&X_{2}(t_{i})\eta_{2}^{(i)}\nu_{2}^{i}&\cdots&X_{d}(t_{i})\eta_{d}^{(i)}\nu_{d}^{i}\\ \end{bmatrix}$.
-(35)
-Assume that $X_{\tau}(t_{i})$ is nonsingular for $\tau=1,\ldots,d$.
-For a fixed $\tau$ and sufficiently small $\Delta t$, the matrices $X_{\tau}(t_{j})$ with $j=i-d+1,\ldots,i$ differ from $X_{\tau}(t_{i})$ only by $\mathcal{O}(\Delta t)$, and the scaling factor $\eta^{(j)}$ vary smoothly as well.
-Consequently, the assembled block matrix can be regarded as a small perturbation of a block Vander monde matrix, which remains nonsingular provided that $\nu_{\tau}$ are pairwise distinct.
-We define an auxiliary STM $\Psi(t_{i},t_{i-1})$ that satisfies
-$\operatorname{col}\left(x_{\tau}^{\mathrm{H}}(t_{j})\,\eta_{\tau}^{(j)}\,\nu_{\tau}^{\,j}\right)_{j=i-d+1}^{i}=\Psi(t_{i},t_{i-1})\operatorname{col}\left(x_{\tau}^{\mathrm{H}}(t_{j})\,\eta_{\tau}^{(j)}\,\nu_{\tau}^{\,j}\right)_{j=i-d}^{i-1}.$ (36)
-The block companion form of $\Psi(t_{i},t_{i-1})$ on subspace $\mathcal{X}_{\tau}$ is given by
-$\left.\Psi(t_{i},t_{i-1})\right|_{\mathcal{X}_{\tau}}=\begin{bmatrix}0&I_{n}&&\\ &\ddots&\ddots&\\ &&0&I_{n}\\ 0&\cdots&0&\frac{\eta_{\tau}^{(i)}}{\eta_{\tau}^{(i-1)}}\nu_{\tau}\Phi_{\tau}(t_{i},t_{i-1})\\ \end{bmatrix}$.
-(37)
-Repeated transition over one full period yields
-$\left.\Psi(t_{i+p},t_{i})\right|_{\mathcal{X}_{\tau}}=$ $\begin{bmatrix}0&\cdots&0&\frac{\eta_{\tau}^{(i-d+1)}}{\eta_{\tau}^{(i)}}\Phi_{\tau}(t_{i+p-d+1},t_{i})\nu_{\tau}^{1-d}\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&0&\frac{\eta_{\tau}^{(i-1)}}{\eta_{\tau}^{(i)}}\Phi_{\tau}(t_{i+p-1},t_{i})\nu_{\tau}^{-1}\\ 0&\cdots&0&\Phi_{\tau}(t_{i+p},t_{i})\\ \end{bmatrix}$.
-(38)
-Let $U_{\tau}(t_{i})$ be the eigenvector matrix of the STM of the $\tau$-th system $\Phi_{\tau}(t_{i+p},t_{i})$ associated with the Floquet multipliers $e^{M_{\tau}T}$ satisfying $\Phi_{\tau}(t_{i+p},t_{i})U_{\tau}(t_{i})=U_{\tau}(t_{i})e^{M_{\tau}T}.$ Denote the stacked eigenvector matrix:
-$\mathcal{U}_{\tau}(t_{i})=\operatorname{col}\left(\Phi_{\tau}(t_{j},t_{i})U_{\tau}(t_{i})\eta_{\tau}^{(j)}\nu_{\tau}^{j}\right)_{j=i-d+1}^{i}.$ (39)
-Then, it follows that $\mathcal{U}_{\tau}(t_{i+p})=\mathcal{U_{\tau}}(t_{i})\nu_{\tau}^{p}e^{M_{\tau}T}$, and
-$\Psi(t_{i+p},t_{i})\mathcal{U_{\tau}}(t_{i})=\mathcal{U_{\tau}}(t_{i})\nu_{\tau}^{p}e^{M_{\tau}T}.$ (40)
-The case $\tau=1$ corresponds to $\nu_{\tau}=1$, which corresponds to the physical system, whose eigenvalues $e^{M_{1}T}$ are the Floquet multipliers of the original LPTV system eq.˜1.
-The remaining $\tau=2,\ldots,d$ correspond to the spurious systems, whose eigenvalues $\nu_{\tau}^{p}e^{M_{\tau}T}$ are scaled versions of the Floquet multipliers of the spurious LPTV systems eq.˜32, and thus decay geometrically with increasing $p$ since $\lvert\nu_{\tau}\rvert<1$.
-5 Spectral Convergence Analysis of Periodic-PEP from Multistep Discretization
-To establish the convergence behavior of the discrete STM $F_{p,0}$ towards $\Psi(t_{p},t_{0})$, we analyze the difference $\Psi(t_{p},t_{0})-F_{p,0}$ acting on $\mathcal{U}_{\tau}$ for physical system $\tau=1$ and spurious systems $\tau=2,\ldots,d$.
-Since all spurious subspaces behave similarly, it suffices to present the proof for the representative two-step case ($d=2$).
-The proof proceeds in three steps:
-  1. 1.
-    Estimate the upper bound of $\left\lVert F_{i+l,i}\right\rVert$.
-  2. 2.
-    Estimate $\left\lVert[\Psi(t_{p},t_{0})-F_{p,0}][\mathcal{U}_{1},\mathcal{U}_{2}]\right\rVert$ by decomposing it into a sum of local differences across steps and bound each term.
-  3. 3.
-    Express the difference $\left\lVert[\Psi(t_{p},t_{0})-F_{p,0}][\mathcal{U}_{1},\mathcal{U}_{2}]\right\rVert$ in an orthonormal basis aligned with $\Psi(t_{p},t_{0})$’s invariant subspaces and apply the subspace perturbation theorem to conclude convergence of the physical and spurious subspaces.
-Before we start the convergence analysis, we first introduce the following assumptions for the LPTV systems, periodic time grid and the multistep methods.
-  1. A1.
-    Consistency and stability of the multistep method: The $d$-step multistep method is of consistency order $s$ and zero-stable.
-Its coefficients $\alpha_{j}^{(i)}$ and $\beta_{j}^{(i)}$ are sufficiently smooth functions of the step-size ratios $\omega^{(i)}$.
-  2. A2.
-    Bounded variable step sizes: The stepsize ratios $\omega^{(i)}$ of the periodic time grid satisfy uniform bounds, i.e., $\omega<\omega^{(i)}<\Omega$, for some constants $0<\omega<1<\Omega$.
-Consequently, the coefficients $\alpha_{j}^{(i)}$ and $\beta_{j}^{(i)}$ admit uniform upper and lower bounds
-    $\alpha_{\min}\leq\alpha_{j}^{(i)}\leq\alpha_{\max},\qquad\beta_{\min}\leq\beta_{j}^{(i)}\leq\beta_{\max},$
-    for all $j=0,1,\ldots,d$ and $i=1,\ldots,p$.
-  3. A3.
-    Smoothness and boundedness of the LPTV system: The periodic matrix $G(t)$ in eq.˜1 is $s$-times continuously differentiable on $[0,T]$.
-In particular, smoothness implies that both $G(t)$ and the state-transition matrix $\Phi(t,s)$ are uniformly bounded:
-    $\|G(t)\|_{\mathrm{F}}\leq C_{G},\qquad\|\Phi(t,s)\|\leq C_{\Phi},\quad\forall\,t-s\in[0,T].$
-We begin by establishing the boundedness of the discrete STM $F_{i+p,i}$.
-Lemma 4 (Stability and Boundedness of Discrete STM).
-Under Assumption A1-A3, let $F_{i+l,i}$ denote the discrete STM over $l$ steps for the multistep discretization of eq.˜1. Then there exists a constant $C_{F}>0$, independent of $l$, such that
-$\|F_{i+l,i}\|_{\mathrm{F}}\leq C_{F},\qquad l=1,\ldots,p.$
-Proof.
-Let $\mathbf{F}_{i+l,i}$ denote the discrete STM associated with the zero scalar system $\dot{y}=0$.
-By definition˜3, it satisfies $\|\mathbf{F}_{i+l,i}\|_{\mathrm{F}}\leq C_{0}$ for all $l\in\mathbb{N}$.
-Strictly speaking, the discrete STM acting on the $nd$-dimensional stacked state is $\mathbf{F}_{i+l,i}\otimes I_{n}$; however, for notational simplicity, we continue to denote it by $\mathbf{F}_{i+l,i}$.
-We now regard $F_{i+l,i}$ as a perturbation of $\mathbf{F}_{i+l,i}$. For each step,
-$\left\lVert F_{i+1,i}-\mathbf{F}_{i+1,i}\right\rVert_{\mathrm{F}}$ $\leq\sum_{j=0}^{d-1}\left\lVert\Delta A_{j}^{(i)}\right\rVert_{\mathrm{F}},$
-where $\Delta A_{j}^{(i)}$ for $j=0,\ldots,d-1$ are defined as
-$\Delta A_{j}^{(i)}$ $\equiv\left(\alpha_{d}^{(i)}I_{n}-\Delta t^{(i)}\beta_{d}^{(i)}G^{(i)}\right)^{-1}\left(\alpha_{j}^{(i)}I_{n}-\Delta t^{(i)}\beta_{j}^{(i)}G^{(i-d+j)}\right)-\alpha_{d}^{-1}\alpha_{j}I_{n}$
-$=\Delta t^{(i)}\left(\alpha_{d}^{(i)}I_{n}-\Delta t^{(i)}\beta_{d}^{(i)}G^{(i)}\right)^{-1}{\alpha_{d}^{(i)}}^{-1}\left(\beta_{d}^{(i)}\alpha_{j}^{(i)}G^{(i)}-\alpha_{d}^{(i)}\beta_{j}^{(i)}G^{(i+j)}\right).$
-For stepsize $\Delta t^{(i)}<\frac{\left\lvert\alpha_{d}^{(i)}\right\rvert}{2\left\lvert\beta_{d}^{(i)}\right\rvert C_{G}}\leq\frac{\alpha_{\max}}{2\beta_{\min}C_{G}}$, we have
-$\left\lVert\Delta A_{j}^{(i)}\right\rVert_{\mathrm{F}}$ $<\left\lvert\alpha_{d}^{(i)}\right\rvert^{-1}\frac{2\alpha^{(i)}\beta^{(i)}C_{G}\Delta t^{(i)}}{\left\lvert\alpha_{d}^{(i)}\right\rvert-\Delta t\left\lvert\beta_{d}^{(i)}\right\rvert C_{G}}\leq 4\alpha_{\min}^{-2}\alpha_{\max}\beta_{\max}C_{G}\Delta t^{(i)},$
-Hence, the bound on $\|\Delta A_{j}^{(i)}\|_{\mathrm{F}}$ holds uniformly for all $i$, leading to $\|F_{i+1,i}-\mathbf{F}_{i+1,i}\|_{\mathrm{F}}\leq C_{1}\,\Delta t^{(i)},$ for some constant $C_{1}$ independent of $i$.
-The cumulative deviation over $l$ steps can be written as
-$F_{i+l,i}-\mathbf{F}_{i+l,i}=\sum_{m=1}^{l}\mathbf{F}_{i+l,i+m}\,\big(F_{i+m,i+m-1}-\mathbf{F}_{i+m,i+m-1}\big)\,F_{i+m-1,i}.$
-Taking norms and applying the triangle inequality yields
-$\|F_{i+l,i}\|_{\mathrm{F}}$ $\leq\|\mathbf{F}_{i+l,i}\|_{\mathrm{F}}+\|\mathbf{F}_{i+l,i+m}\|_{\mathrm{F}}C_{1}\sum_{m=1}^{l}\Delta t^{(i+m-1)}\|F_{i+m-1,i}\|_{\mathrm{F}}$
-$\leq C_{0}+C_{0}C_{1}\sum_{m=1}^{l}\Delta t^{(i+m-1)}\|F_{i+m-1,i}\|_{\mathrm{F}}.$
-Finally, using the discrete Gronwall–Bellman inequality [22], we obtain
-$\|F_{i+l,i}\|_{\mathrm{F}}\leq C_{0}e^{C_{0}C_{1}T}\equiv C_{F}.$
-Thus the discrete STM remains uniformly bounded within each period. ∎
-Lemma 5.
-Under assumptions A1-A3, let $\Psi(t_{p},t_{0})$ and $F_{p,0}$ denote the auxiliary and discrete STM over $t_{0},t_{p}]$.
-Consider the difference $\Psi(t_{p},t_{0})-F_{p,0}$ acting on the normalized eigenvector matrices $\mathcal{U}_{\cdot}(t_{0})$ of $\Psi(t_{p},t_{0})$, with $\lVert\mathcal{U}_{\cdot}(t_{0})\rVert_{\mathrm{F}}=1$.
-Then the following bounds hold:
-  1. 1.
-    $\left\lVert\left[\Psi(t_{p},t_{0})-F_{p,0}\right]\mathcal{U}_{1}(t_{0})\right\rVert_{\mathrm{F}}\leq C_{1}\Delta t^{s}$,
-  2. 2.
-    $\left\lVert\left[\Psi(t_{p},t_{0})-F_{p,0}\right]\mathcal{U}_{2}(t_{0})\right\rVert_{\mathrm{F}}\leq C_{2}\Delta t$.
-with constants $C_{1}$, $C_{2}$ independent of $\Delta t$.
-Proof.
-We decompose the global difference into sums of local difference:
-$\left[\Psi(t_{p},t_{0})-F_{p,0}\right]\mathcal{U}_{\tau}(t_{0})$ (41)
-$=$ $\sum_{i=1}^{p}F_{p,i}\cdot\left[\Psi(t_{i},t_{i-1})-F_{i,i-1}\right]\cdot\Psi_{i-1,0}\cdot\mathcal{U}_{\tau}(t_{0})$
-$=$ $\sum_{i=1}^{p}F_{p,i}\cdot\left[\Psi(t_{i},t_{i-1})-F_{i,i-1}\right]\cdot\operatorname{col}\left(\Phi_{\tau}(t_{j},t_{0})U_{\tau}(t_{0})\eta_{\tau}^{(j)}\nu_{\tau}^{j}\right)_{j=i-d+1}^{i}.$
-From eq.˜33 and lemma˜3, each local difference acts only on the last block of the column, giving
-$\left[\Psi(t_{i},t_{i-1})-F_{i,i-1}\right]\operatorname{col}\left(\Phi_{\tau}(t_{j},t_{i})U_{\tau}(t_{i})\eta_{\tau}^{(j)}\nu_{\tau}^{j}\right)_{j=i-d+1}^{i}=\left[0,E_{i}\right],$
-$\left\lVert E_{1,i}\right\rVert=\left\lVert\Phi_{1}(t_{i},t_{0})U_{1}(t_{0})\right\rVert\mathcal{O}(\Delta t^{s+1}),$
-$\left\lVert E_{\tau,i}\right\rVert=\nu_{\tau}^{i}\left\lVert\Phi_{\tau}(t_{i},t_{0})U_{\tau}(t_{0})\right\rVert\mathcal{O}(\Delta t).$
-For the physical mode ($\tau=1$), applying the triangle inequality and the uniform boundedness of $F_{p,i}$ and $\Phi_{1}(t_{i},t_{0})U_{1}(t_{0})$, we obtain
-$\big\|\big[\Psi(t_{p},t_{0})-F_{p,0}\big]\mathcal{U}_{1}(t_{0})\big\|_{\mathrm{F}}$ $\leq\sum_{i=1}^{p}\|F_{p,i}\|_{\mathrm{F}}\,\|E_{1,i}\|\leq\sum_{i=1}^{p}C^{(i)}\,\Delta t^{s+1}\leq C_{1}\,\Delta t^{s},$
-where $C_{1}$ is independent of $\Delta t$. This shows that the global error for the physical mode is $\mathcal{O}(\Delta t^{s})$.
-For the spurious mode ($\tau=2$), we similarly have
-$\big\|\big[\Psi(t_{p},t_{0})-F_{p,0}\big]\mathcal{U}_{2}(t_{0})\big\|_{\mathrm{F}}$ $\leq\sum_{i=1}^{p}\|F_{p,i}\|_{\mathrm{F}}\,\|E_{2,i}\|\leq\sum_{i=1}^{p}C_{F}\,C_{\mathcal{U}}\,\nu_{2}^{i}\,\Delta t\leq C_{2}\,\Delta t,$
-where we have used that $\Phi_{2}(t_{i},t_{0})$ and the initial vector $U_{2}(t_{0})$ are uniformly bounded, and that the geometric series $\sum_{i=1}^{p}\nu_{2}^{i}$ converges to a constant independent of $\Delta t$.
-∎
-To apply the classical subspace perturbation theorem [19, Theorem 3.5], we express $\Psi(t_{p},t_{0})-F_{p,0}$ in an orthonormal basis of $\Psi(t_{p},t_{0})$’s invariant subspaces.
-Lemma 6.
-Under assumptions A1-A3, there exists an orthogonal matrix $P^{(0)}=[P_{1}^{(0)},P_{2}^{(0)}]$ of $\Psi(t_{p},t_{0})$, such that
-${P^{(0)}}^{\mathrm{H}}\left[\Psi(t_{p},t_{0})-F_{p,0}\right]P^{(0)}=\begin{bmatrix}E_{11}&E_{12}\\ E_{21}&E_{22}\end{bmatrix}$, (42)
-with blockwise norms: $\left\lVert E_{11}\right\rVert_{\mathrm{F}},\left\lVert E_{21}\right\rVert_{\mathrm{F}}=\mathcal{O}(\Delta t^{s}),\quad\left\lVert E_{12}\right\rVert_{\mathrm{F}},\left\lVert E_{22}\right\rVert_{\mathrm{F}}=\mathcal{O}(\Delta t).$
-Proof.
-Define $\operatorname{span}[V(t_{0})]:=\operatorname{span}[\mathcal{U}_{1}(t_{0}),\mathcal{U}_{2}(t_{0})]$ from the eigenvector matrix of $\Psi(t_{p},t_{0})$.
-Construct $P^{(0)}$ as its orthonormal basis of $\operatorname{span}[V(t_{0})]$ via QR decomposition, then $P^{(0)}R(t_{0})^{-1}=V(t_{0}),$ where $R(t_{0})^{-1}$ is an block upper-triangular matrix.
-Note that $R(t_{0})$ is uniformly bounded since $V(t_{0})$ has full column rank and $\|V(t_{0})\|_{\mathrm{F}}$ is uniformly bounded.
-From the definition of $P^{(0)}$, we have:
-${P^{(0)}}^{\mathrm{H}}\left(\Psi(t_{p},t_{0})-F_{p,0}\right)P^{(0)}$
-$=$ ${P^{(0)}}^{\mathrm{H}}\left(\Psi(t_{p},t_{0})-F_{p,0}\right)\left[\mathcal{U}_{1},\mathcal{U}_{2}\right](t_{0})\begin{bmatrix}R_{11}(t_{0})&R_{12}(t_{0})\\ &R_{22}(t_{0})\\ \end{bmatrix}$
-$\equiv$ $\begin{bmatrix}E_{11}&E_{12}\\ E_{21}&E_{22}\end{bmatrix}(t_{0}).$
-According to lemma˜5, the first block column can be bounded as $\mathcal{O}(\Delta t^{s})$ and the second block column as $\mathcal{O}(\Delta t)$. ∎
-Since $\left\lVert E_{21}\right\rVert_{\mathrm{F}}=\mathcal{O}(\Delta t^{s})$, $P_{1}^{(0)}$ is an approximate invariant subspace of $F_{p,0}$.
-Our next work is to show that the difference between the invariant subspace of $F_{p,0}$ and $P_{1}^{(0)}$.
-This needs the help of the subspace perturbation theory.
-From theorem˜1, we already have a quantitative lower bound for the subspace separation $\operatorname{sep}(\Psi_{11},\Psi_{22})$ based on the spectral information of $\Psi(t_{p},t_{0})$.
-This bound, together with the subspace perturbation theorem theorem˜2, yields the convergence of both the eigenvalues and corresponding subspace of $F_{p,0}$ towards those of $\Psi(t_{p},t_{0})$ as $\Delta t\to 0$.
-This result is summarized in theorem˜7.
-Theorem 7 (Convergence of Eigenvalues and Invariant Subspaces).
-Under assumptions A1-A3, the eigenpairs of $F_{p,0}$ admits the following convergence properties as $\Delta t\to 0$:
-  1. 1.
-    The $n$ largest eigenvalues of $F_{p,0}$ converge to the $n$ Floquet multipliers of $\Psi(t_{p},t_{0})$ with order $\mathcal{O}(\Delta t^{s})$.
-  2. 2.
-    The corresponding $n$-dimensional invariant subspace of $F_{p,0}$ converges to that of $\Psi(t_{p},t_{0})$ with order $\mathcal{O}(\Delta t^{s})$.
-  3. 3.
-    The remaining (d1)n(d-1)n eigenvalues of $F_{p,0}$ asymptotically converge to zero as $\Delta t\to 0$, with the convergence rate improving to $\mathcal{O}(\Delta t^{2})$ when the stepsize is constant.
-Proof.
-By theorem˜1, since $\Psi_{11}$ is independent of $\Delta t$ and the eigenvalues of $\Psi_{22}$ vanish as $\Delta t\to 0$, there exists a constant $C_{\text{sep}}$ such that $\operatorname{sep}(\Psi_{11},\Psi_{22})>2C_{\text{sep}}$ for sufficiently small $\Delta t$.
-From lemma˜6, we have $\left\lVert E_{11}\right\rVert+\left\lVert E_{22}\right\rVert=\mathcal{O}(\Delta t)$, hence $\left\lVert E_{11}\right\rVert+\left\lVert E_{22}\right\rVert\leq C_{\text{sep}},$ which ensures that the spectral gap condition $\delta\geq C_{\text{sep}}$ in theorem˜2 holds uniformly.
-Next, the perturbation requirement $\left\lVert E_{21}\right\rVert\big(\left\lVert E_{12}\right\rVert+\left\lVert A_{12}\right\rVert\big)<\tfrac{1}{4}\delta^{2}$ is also satisfied for sufficiently small $\Delta t$, since $\left\lVert E_{21}\right\rVert=\mathcal{O}(\Delta t^{s})\to 0$ while the remaining terms stay bounded.
-Therefore, the conditions of theorem˜2 are met, implying the existence of a matrix $R$ such that $\left\lVert R\right\rVert\leq\frac{2}{C_{\text{sep}}}C_{1}\Delta t^{s}\equiv C_{2}\Delta t^{s},$ and
-$F_{p,0}P_{1}^{(0)}\begin{bmatrix}I\\ R\end{bmatrix}=P_{1}^{(0)}\begin{bmatrix}I\\ R\end{bmatrix}(\Psi_{11}+\Delta\Psi_{11}),\qquad\Delta\Psi_{11}=E_{11}-(\Psi_{12}+E_{12})R.$ (43)
-Hence, the invariant subspace of $F_{p,0}$ associated with the $n$ dominant eigenvalues deviates from that of $\Psi(t_{p},t_{0})$ by $\mathcal{O}(\Delta t^{s})$.
-Finally, since $\Delta\Psi_{11}=\mathcal{O}(\Delta t^{s})$, the Bauer–Fike theorem [23] yields
-$\max_{\lambda_{i}\in\lambda(\Psi_{11})}\max_{\mu_{j}\in\lambda(\Psi_{11}+\Delta\Psi_{11})}\left\lvert\lambda_{i}-\mu_{j}\right\rvert$ $\leq\left\lVert(\Psi_{12}+E_{12})R-E_{11}\right\rVert_{\mathrm{F}}$ (44)
-$\leq\left[C_{R}\left\lVert\Psi_{12}+E_{12}\right\rVert_{\mathrm{F}}+C_{E}\right]\Delta t^{s}.$
-The same reasoning applies to the remaining eigenpairs, which similarly converge to zero as $\Delta t\to 0$. ∎
-The key mechanism underlying theorem˜7 is the use of the spectral separation between the physical and spurious subspaces of $\Psi(t_{p},t_{0})$: since all physical Floquet multipliers are strictly nonzero while the spurious ones vanish as $\Delta t\to 0$, a positive spectral gap guarantees the separation $\operatorname{sep}(\Psi_{11},\Psi_{22})>0$ and thus the convergence of the corresponding invariant subspaces.
-In many applications, only the invariant subspace of the top $k$ Floquet multipliers is of interest.
-Assuming this subspace of $\Psi(t_{p},t_{0})$ remains separated from its complement as $\Delta t\to 0$, the same perturbation argument yields:
-Corollary 8 (Convergence of Partial Floquet Multipliers and Subspaces).
-Under the assumptions of theorem˜7, if the subspace separation between the invariant subspaces associated with the $k$ largest eigenvalues of $\Psi(t_{p},t_{0})$ and the remaining ones remains nonzero, then the $k$ largest eigenvalues of $F_{p,0}$ and their invariant subspace converge to the corresponding Floquet multipliers and invariant subspace of $\Psi(t_{p},t_{0})$ with rate $\mathcal{O}(\Delta t^{s})$.
-In practice, the subspace separation $\operatorname{sep}$ is rarely computed explicitly.
-Instead, it is often estimated using the spectral gap, which, according to theorem˜1, serves as an upper bound for $\operatorname{sep}$.
-Hence, it is advisable to choose $k$ such that the $k$-th largest eigenvalue is well separated from the $(k+1)$-th one, with a spectral gap typically on the order of $\mathcal{O}(10^{-3})$ or larger.
-Empirically, when the spectral gap is sufficiently large, the actual separation tends to approach this upper bound rather than its lower limit, leading to markedly improved convergence behavior.
+If $\delta\equiv\operatorname{sep}(A_{11},A_{22})-\left(\left\lVert E_{11}\right\rVert+\left\lVert E_{22}\right\rVert\right)>0$, and $\left\lVert E_{21}\right\rVert\left(\left\lVert A_{12}\right\rVert+\left\lVert E_{12}\right\rVert\right)<\frac{1}{4}\delta^{2}$, then there exist a matrix $R\in\mathbb{C}^{(n-l)\times l}$ satisfying $\left\lVert R\right\rVert\leq 2\delta^{-1}\left\lVert E_{21}\right\rVert$, such that $\operatorname{span}\left\{P\begin{bmatrix}I\\ R\end{bmatrix}\right\}$ is an invariant subspace of $A+E$.
+Remark 4.7.
+The separation between two square matrices $A_{11}$ and $A_{22}$ is defined as
+$\operatorname{sep}_{2}(A_{11},A_{22}):=\inf_{\|X\|_{2}=1}\|A_{11}X-XA_{22}\|_{2}.$
+If $A_{11}$ and $A_{22}$ are both normal, then the separation coincides with the spectral gap between their spectra.
+$\operatorname{sep}_{2}(A_{11},A_{22})=\mathrm{d}(A_{11},A_{12})\equiv\min\{\left\lvert\lambda-\mu\right\rvert,\,\lambda\in\lambda(A_{11}),\mu\in\lambda(A_{22})\}.$
+With the global difference estimate and the invariant subspace perturbation theorem at hand, we are now in a position to complete the proof of Theorem˜4.1.
+Proof 4.8.
+The three statements are proved by the same perturbation argument based on the block representation of $\mathcal{Z}^{-1}(t_{0})F_{p,0}\mathcal{Z}(t_{0})$ and a spectral separation estimate.
+We therefore present the proof for part (3), which contains the main ideas.
+Parts (2) and (1) follow by the same reasoning with minor simplifications, as outlined at the end of the proof.
+Let $1\leq k\leq n$, we partition $\mathcal{X}(t_{0})$ as $\mathcal{X}(t_{0})=\begin{bmatrix}X_{k}(t_{0})&X_{n-k}(t_{0})\end{bmatrix}$ After the similarity transformation induced by the moving basis $\mathcal{Z}(t_{0})$, the one-period stacked propagator admits the block representation Eq.˜29:
+$\mathcal{Z}^{-1}(t_{0})F_{p,0}\mathcal{Z}(t_{0})$
+$=$ $\begin{bmatrix}e^{M_{k}T}&&\\ &e^{M_{n-k}T}\\ &&D^{p}\end{bmatrix}+\begin{bmatrix}e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{k}}&e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{n-k}}&e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{Y}}\\ e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{X}_{k}}&e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{X}_{n-k}}&e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{Y}}\\ D^{p}E_{\mathcal{Y}\mathcal{X}_{k}}&D^{p}E_{\mathcal{Y}\mathcal{X}_{n-k}}&D^{p}E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}$.
+For brevity, we set
+$E_{11}=e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{k}},\quad E_{12}=\begin{bmatrix}e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{n-k}}&e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{Y}}\end{bmatrix}$,
+$E_{21}=\begin{bmatrix}e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{X}_{k}}\\ D^{p}E_{\mathcal{Y}\mathcal{X}_{k}}\end{bmatrix},\quad E_{22}=\begin{bmatrix}e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{X}_{n-k}}&e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{Y}}\\ D^{p}E_{\mathcal{Y}\mathcal{X}_{n-k}}&D^{p}E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}$.
+Since all Floquet multipliers $e^{\mu_{k}T}$ are assumed semisimple and sorted by non-increasing order, thus $e^{\mu_{n}T}$ is the smallest magnitude Floquet multiplier.
+Since $\|D^{p}\|=\nu_{\max}^{p}\to 0$ as $p\to\infty$, there exists $p_{0}\in\mathbb{N}$ such that $\|D^{p}\|\leq e^{\mu_{n}T}/2$ for all $p\geq p_{0}$.
+Because both $e^{MT}$ and $D^{p}$ are diagonal (hence normal), their spectra are separated, and
+$\operatorname{sep}(e^{MT},D^{p})=\operatorname{dist}\!\bigl(\lambda(e^{MT}),\lambda(D^{p})\bigr)\geq e^{\mu_{n}T}/2,\qquad p\geq p_{0}.$
+For sufficient large $p$, we have
+$\begin{aligned} \delta=&\operatorname{sep}(e^{M_{k}T},\begin{bmatrix}e^{M_{n-k}}&\\ &D^{p}\end{bmatrix})-\left(\left\lVert e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{k}}\right\rVert+\left\lVert\begin{bmatrix}E_{11}&\\ &E_{22}\end{bmatrix}\right\rVert\right)\\ >&1/2\left\lvert e^{\mu_{m}T}-e^{\mu_{m+1}T}\right\rvert\end{aligned}$,
+and
+$\left\lVert E_{21}\right\rVert\left(\left\lVert 0\right\rVert+\left\lVert E_{21}+E_{12}\right\rVert\right)<\frac{1}{4}\delta^{2}.$
+Thus, the conditions of Theorem˜4.6 are satisfied, and there exists a matrix $R\in\mathbb{C}^{((d-1)n+k)\times k}$ with
+$\|R\|\leq C\,\nu_{\max}^{p},$
+such that
+$\mathcal{Z}^{-1}(t_{0})F_{p,0}\mathcal{Z}(t_{0})\begin{bmatrix}I_{k}\\ R\end{bmatrix}=\begin{bmatrix}I_{k}\\ R\end{bmatrix}B,$
+where the reduced $n\times n$ matrix $B$ is given by
+$B:=e^{M_{k}T}+e^{M_{k}T}E_{11}-e^{M_{k}T}E_{12}R.$
+Thus,
+$F_{p,0}\left(\mathcal{X}_{k}(t_{0})+\begin{bmatrix}\mathcal{X}_{n-k}(t_{0})&\mathcal{Y}(t_{0})\end{bmatrix}R\right)=\left(\mathcal{X}_{k}(t_{0})+\begin{bmatrix}\mathcal{X}_{n-k}(t_{0})&\mathcal{Y}(t_{0})\end{bmatrix}R\right)B,$
+that is $\mathcal{X}_{k,0}=\mathcal{X}_{k}(t_{0})+\begin{bmatrix}\mathcal{X}_{n-k}(t_{0})&\mathcal{Y}(t_{0})\end{bmatrix}R$, thus we proved (3).
+Using $\|E_{11}\|=O(h^{s})$, $\|E_{12}\|=O(h^{s})$ and $\|R\|=O(\nu_{\max}^{p})$, we obtain
+$\|B-e^{M_{k}T}\|\leq Ch^{s}+C\nu_{\max}^{p}.$
+Since $e^{M_{k}T}$ is diagonal (hence normal), its eigenvalues are stable under perturbations:
+$\operatorname{dist}\!\bigl(\lambda(B),\lambda(e^{M_{k}T})\bigr)\leq\|B-e^{M_{k}T}\|.$
+Therefore, after suitable ordering, the $n$ eigenvalues of $F_{p,0}$ associated with the dominant spectral cluster satisfy
+$\lambda_{k}(h)=e^{\mu_{k}T}+O(h^{s})+O(\nu_{\max}^{p}),\qquad k=1,\ldots,n.$
+Because $\nu_{\max}^{p}$ decays geometrically as $p\to\infty$, this yields the claimed $O(h^{s})$ convergence of the principal eigenvalues. Let $k=n$, we proved $(1)$.
+The remaining (d1)n(d-1)n eigenvalues correspond to the complementary block and arise from a perturbation of $D^{p}$ of size $O(\nu_{\max}^{p})$. Hence each parasitic eigenvalue satisfies
+$|\lambda_{k}|\leq C\|D^{p}\|\leq C\nu_{\max}^{p},\quad k=n+1,\ldots,dn,$
+which proves the geometric decay of the parasitic spectrum.
 5 Algorithm and Complexity Analysis
-We present pTOAR, an iterative method for computing the dominant $k$ eigenpairs of the periodic-PEP eq.˜18.
-It combines the periodic Arnoldi process [12] with TOAR-style basis compression [24] to reduce memory and computational cost, making it suitable for large-scale problems.
+We present pTOAR, an iterative method for computing the dominant $k$ eigenpairs of the pPEP Eq.˜14.
+It combines the periodic Arnoldi process [18] with TOAR style basis compression [19] to reduce memory and computational cost, making it suitable for large-scale problems.
 The remainder of this section details the algorithmic steps and its underlying principles.
-The periodic Krylov-Schur (pKS) algorithm [12] is well-suited for the linearized periodic-PEP eq.˜22.
+The periodic Krylov-Schur (pKS) algorithm [18] is well-suited for the linearized pPEP Eq.˜19.
 It avoids explicitly forming the full product $F_{i+p,i}=L^{(p)}\cdots\allowbreak L^{(1)}$, while maintaining accuracy even for eigenvalues spanning several orders of magnitude.
 The pKS algorithm constructs a periodic Arnoldi process:
-$L^{(i)}V^{(i-1)}=V^{(i)}T^{(i)},$ (45)
+(30) $L^{(i)}V^{(i-1)}=V^{(i)}T^{(i)},$
 where $V^{(i)}$ for $i=1,\ldots,p-1$ are $(nd)\times k$ orthonormal matrices, and $V^{(p)}\in\mathbb{C}^{nd\times(k+1)}$, which includes the additional Arnoldi residual vector.
@@
 We now show the details of the compression mechanism.
-We explicitly substitute the block form of $L^{(i)}$ into eq.˜45, and partition the $V^{(i)}$ following the block structure of $L^{(i)}$, as shown in eq.˜46:
-$\begin{bmatrix}&I_{n}&&\\ &&\ddots&\\ &&&I_{n}\\ A_{0}^{(i)}&A_{1}^{(i)}&\cdots&A_{d-1}^{(i)}\\ \end{bmatrix}\begin{bmatrix}V_{1}^{(i-1)}\\ \vdots\\ V_{d-1}^{(i-1)}\\ V_{d}^{(i-1)}\end{bmatrix}=\begin{bmatrix}V_{1}^{(i)}\\ \vdots\\ V_{d-1}^{(i)}\\ V_{d}^{(i)}\end{bmatrix}T^{(i)}.$ (46)
-The key observation of pTOAR from eq.˜46 is an information overlap relationship across adjacent basis matrices:
+We explicitly substitute the block form of $L^{(i)}$ into Eq.˜30, and partition the $V^{(i)}$ following the block structure of $L^{(i)}$, as shown in Eq.˜31:
+(31) $\begin{bmatrix}&I_{n}&&\\ &&\ddots&\\ &&&I_{n}\\ A_{0}^{(i)}&A_{1}^{(i)}&\cdots&A_{d-1}^{(i)}\\ \end{bmatrix}\begin{bmatrix}V_{1}^{(i-1)}\\ \vdots\\ V_{d-1}^{(i-1)}\\ V_{d}^{(i-1)}\end{bmatrix}=\begin{bmatrix}V_{1}^{(i)}\\ \vdots\\ V_{d-1}^{(i)}\\ V_{d}^{(i)}\end{bmatrix}T^{(i)}.$
+The key observation of pTOAR from Eq.˜31 is an information overlap relationship across adjacent basis matrices:
 $V_{j+1}^{(i)}T^{(i)}=V_{j}^{(i+1)},\,j=1,\ldots,d-1,\,i=1,\ldots,p.$
@@
 $L^{(i)}v^{(i-1)}=\left[{v_{2}^{(i-1)}}^{\mathrm{H}},\cdots,{v_{d}^{(i-1)}}^{\mathrm{H}},{v_{d}^{(i)}}^{\mathrm{H}}\right]^{\mathrm{H}},\,v_{d}^{(i)}={A_{d}^{(i)}}^{-1}\left(\sum_{j=0}^{d-1}A_{j}^{(i)}v_{j}^{(i-1)}\right),$
-where $v^{(i-1)}$ is the last column of $V^{(i-1)}$, and $v_{j}^{(i-1)}$ is its $j$-th block of $v^{(i-1)}$, also the last column of $V_{j}^{(i-1)}$ in eq.˜46.
+where $v^{(i-1)}$ is the last column of $V^{(i-1)}$, and $v_{j}^{(i-1)}$ is its $j$-th block of $v^{(i-1)}$, also the last column of $V_{j}^{(i-1)}$ in Eq.˜31.
 The second step in periodic Arnoldi basis construction is orthogonalization.
@@
 At last, we evaluate memory requirement for storing the input data.
-For each $L^{(i)}$, it suffices to store the sampled matrices $G^{(i)}$ instead of the coefficient matrix $A_{j}^{(i)}$ in eq.˜14 and the full $nd\times nd$ matrix $L^{(i)}$.
+For each $L^{(i)}$, it suffices to store the sampled matrices $G^{(i)}$ instead of the coefficient matrix $A_{j}^{(i)}$ in Eq.˜12 and the full $nd\times nd$ matrix $L^{(i)}$.
 This leads to a memory cost of $n^{\gamma}$ per slice, where $\gamma=1$ if the matrices $A_{j}^{(i)}$ are sparse, and $\gamma=2$ if dense.
@@
 Algorithm 1 Algorithm for Computing Dominant Floquet Eigenpairs with Multistep Discretization and Periodic Iterative Solvers
-1:
-2:Time samples $\{t_{i}\}_{i=0}^{p}$ with $0=t_{0}<t_{1}<\cdots<t_{p}=T$, where $T$ is the period;
-3:Discrete evaluations $\{G(t_{i})\}_{i=0}^{p}$ of the system matrix;
-4:Target number of eigenpairs $k$.
-5:
-6:Dominant $k$ Floquet multipliers and their corresponding eigenvectors.
-7:Discretization: At each time $t_{i}$, define $\left\{A_{j}^{(i)}\right\}_{j=0}^{d}$ in eq.˜14 implicitly via matrix-vector products routines, and invoke them to define matrix-vector products of $L^{(i)}$, as ineq.˜21.
-8:Eigenpair Computation: Apply the pTOAR to compute the dominant $k$ eigenpairs of the linearized periodic EVP eq.˜22.
+0:Time samples $\{t_{i}\}_{i=0}^{p}$ with $0=t_{0}<t_{1}<\cdots<t_{p}=T$, where $T$ is the period;Discrete evaluations $\{G(t_{i})\}_{i=0}^{p}$ of the system matrix; Target number of eigenpairs $k$.
+0:Dominant $k$ Floquet multipliers and their corresponding eigenvectors.
+1:Discretization: At each time $t_{i}$, define $\left\{A_{j}^{(i)}\right\}_{j=0}^{d}$ in Eq.˜12 implicitly via matrix-vector products routines, and invoke them to define matrix-vector products of $L^{(i)}$, as inEq.˜18.
+2:Eigenpair Computation: Apply the pTOAR to compute the dominant $k$ eigenpairs of the linearized periodic EVP Eq.˜19.
 6 Numerical Results
-This section presents three numerical experiments designed to validate our proposed method.
-The first is an artificial example constructed from known Floquet eigenpairs to verify convergence theory and the distribution of Spurious eigenvalues.
-The second is a benchmark problem drawn from [25, 6], representative of scenario 1.
-We compare our pTOAR method with existing packages MATCONT [10] and AUTO-07p [9].
-The third experiment targets an industrial circuit application, representative of scenario 2, where fixed sampling makes collocation infeasible and highlights the advantage of our multistep-based method.
+This section presents three numerical experiments.
+The first is an artificial example constructed from known Floquet eigenpairs to verify convergence theory and the distribution of parasitic eigenvalues.
+The second and third experiments address two common scenarios where Floquet multipliers are needed.
+In the first scenario, the matrix $G(t;\delta)$ is given explicitly, and tracking the evolution of multipliers with respect to $\delta$ helps detect bifurcations or instability onset [20].
+We test this scenario using a benchmark problem and compare our pTOAR method with existing packages MATCONT and AUTO-07p.
+In the second scenario, $G(t)$ is only available at discrete time samples, without an analytic form; here the focus is often on computing the Floquet eigenvector associated with the dominant multiplier (e.g., in circuit oscillator design) [9, 10].
+We illustrate this scenario with an industrial circuit application, where fixed sampling makes collocation infeasible and highlights the advantage of our multistep-based approach.
 All numerical experiments are conducted using MATLAB 2024b on a system equipped with an Intel Core i5-1135G7 CPU at 2.4 GHz and 16 GB of RAM.
 We assess the accuracy of the computed Floquet multipliers $\lambda$ and eigenvectors $x$ using relative errors:
-$e_{\text{val}}=\frac{\left|\lambda-e^{\mu T}\right|}{\left|e^{\mu T}\right|},\quad e_{\text{vec}}=\frac{\|x_{p}-x(T)\|_{2}}{\|x(T)\|_{2}},$ (47)
+(32) $e_{\text{val}}=\frac{\left|\lambda-e^{\mu T}\right|}{\left|e^{\mu T}\right|},\quad e_{\text{vec}}=\frac{\|x_{p}-x(T)\|_{2}}{\|x(T)\|_{2}},$
 where $e^{\mu T}$ and $x(T)$ are the exact Floquet multiplier and eigenvector.
 For the invariant subspace spanned by several eigenvectors, we define the subspace error between computed subspace $U$ and exact subspace $V$ by
-$e_{\text{subspace}}=\Theta(U,V),$ (48)
+(33) $e_{\text{subspace}}=\Theta(U,V),$
 where $\Theta(U,V)$ is the principal angle between the two subspaces.
 When exact solutions are not available, we use higher-order multistep methods to generate reference solutions, whose discrepancy from the exact operator is negligible for our evaluation purposes.
-7.1 Convergence Behavior of Floquet multipliers and Spurious Eiegnvalues
-Our first experiment is based on a two-dimensional LPTV system given by eq.˜49,
-$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}\alpha\cos^{2}t+(\beta-2)\sin^{2}t&1+(\beta-\alpha-2)\sin t\cos t\\ -1+(\beta-\alpha-2)\sin t\cos t&\alpha\sin^{2}t+(\beta-2)\cos^{2}t\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$.
-(49)
+6.1 Convergence Behavior of Floquet multipliers and Parasitic Eiegnvalues
+Our first experiment is based on a two-dimensional LTP system given by Eq.˜34,
+(34) $\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}\alpha\cos^{2}t+(\beta-2)\sin^{2}t&1+(\beta-\alpha-2)\sin t\cos t\\ -1+(\beta-\alpha-2)\sin t\cos t&\alpha\sin^{2}t+(\beta-2)\cos^{2}t\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$.
 This model is obtained by linearizing the Stuart-Landau oscillator along its limit cycle.
 Its Floquet multipliers are $e^{2\pi\alpha}$ and $e^{2\pi(\beta-2)}$, and their corresponding eigenvectors are columns of
-$U(t)=\begin{bmatrix}\cos t&\sin t\\ -\sin t&\cos t\end{bmatrix}$. (50)
-We discretize one period $0,2\pi]$ with variable stepsize, and apply Gear2, Gear3 methods to obtain the periodic-EVP eq.˜22.
-The Floquet multiplier, its associated eigenvector, and the spurious eigenvalues from multistep methods are then computed using the periodic QR algorithm [11] with full precision.
+(35) $U(t)=\begin{bmatrix}\cos t&\sin t\\ -\sin t&\cos t\end{bmatrix}$.
+We discretize one period $0,2\pi]$ with variable stepsize, and apply Gear2, Gear3 methods to obtain the periodic-EVP Eq.˜19.
+The Floquet multiplier, its associated eigenvector, and the parasitic eigenvalues from multistep methods are then computed using the periodic QR algorithm [5] with full precision.
 For reference, we also apply the backward Euler (BE) discretization, which does not produce spurio us eigenvalues, to serve as a baseline for convergence comparison.
@@
 The system is discretized over one period into $2^{k}$ points with stepsizes alternating in the ratio $1:2:1:2:$...
+We study the convergence of the computed Floquet eigenpairs with respect to the largest stepsize.
+Through this experiment, we set $\alpha=0.1$ and $\beta=0.1$.
+The system is discretized over one period into $2^{k}$ grid points with variable stepsizes.
+To ensure that the resulting time grid satisfies Assumption (A4), we evenly insert four steps where the stepsize is doubled and four steps where the stepsize is halved, so that the total deviation of the step ratios from unity remains bounded over one period.
 We study the convergence of the computed Floquet eigenpairs with respect to the largest stepsize.
@@
 Only the errors for the dominant Floquet multiplier and its eigenvector are shown here, as the results for the other mode are similar.
-(a) Spurious Eigenvalues of Gear2
-(b) Spurious Eigenvalues of Gear3
-Figure 2: Convergence Behavior of Spurious Eigenvalues
-Figures˜2(a) and 2(b) illustrate the convergence of the spurious eigenvalues associated with Gear2 (two spurious eigenvalues) and Gear3 (four spurious eigenvalues), respectively, as the discretization Stepsize decreases.
-Since these spurious eigenvalues are extremely small, their logarithms are plotted on the vertical axis, and different solid markers are used to distinguish individual eigenvalues.
-In addition, based on the multistep method and eq.˜28, we compute the theoretical reference $\nu^{p}$, which represents the expected magnitude of these spurious eigenvalues from the corresponding auxiliary STM, and plot its logarithm as red lines in Figures˜2(a) and 2(b).
-Strictly speaking, each spurious eigenvalue is equal to the product of the Floquet multiplier of the Spurious system and $\nu^{p}$, where the Floquet multiplier is of moderate size.
-Therefore, we only compare the order of magnitude, rather than the exact value.
-Consequently, the slope of the red line corresponds to $\log(\nu)$ and serves as a reference for expected convergence.
-The figures show that the computed spurious eigenvalues match the expected magnitude very well, far exceeding the prediction of theorem˜7, which guarantees only linear decay with $\Delta t$.
+(a) Parasitic Eigenvalues of Gear2
+(b) Parasitic Eigenvalues of Gear3
+Figure 2: Convergence Behavior of Parasitic Eigenvalues
+Figures˜2(a) and 2(b) illustrate the convergence of the parasitic eigenvalues of Gear2 (two eigenvalues) and Gear3 (four eigenvalues), respectively, as the stepsize decreases.
+Since these eigenvalues are very small, their logarithms are shown, with different markers indicating individual modes.
+The red lines represent $\log(\nu)$ and serve as a reference for the expected decay rate.
+As each parasitic eigenvalue equals a Floquet multiplier of moderate magnitude times $\nu^{p}$, we compare only orders of magnitude rather than exact values.
+The figures show that the computed parasitic eigenvalues match the expected magnitude very well, far exceeding the prediction of Theorem˜4.1, which guarantees only linear decay with $h$.
 6.2 Scenario 1: Coupled Stuart-Landau Oscillators
@@
 We fix $\beta=0.5$ and use $\delta$ as the continuation parameter, approaching $\delta=0.25$ from below.
-Our focus is on the Floquet multipliers of the variational system associated to a limit cycle in eq.˜51, where two oscillators are $\pi$ radians out of phase, i.e., $x_{1}=-x_{2},y_{1}=-y_{2}$.
+Our focus is on the Floquet multipliers of the variational system associated to a limit cycle in Eq.˜36, where two oscillators are $\pi$ radians out of phase, i.e., $x_{1}=-x_{2},y_{1}=-y_{2}$.
 As $\delta$ approaches $0.25$, this limit cycle approaches a heteroclinic loop, and the period grows raplidly, and the Floquet multipliers vary sharply.
 Our goal is to compute its Floquet multipliers with high accuracy to determine their position relative to the unit circle in the complex plane.
-For a given pair of $(\delta,\beta)$, the periodic solution is first computed using existing continuation software—MATCONT [10] in MATLAB and AUTO-07p [26, 9] in FORTRAN.
+For a given pair of $(\delta,\beta)$, the periodic solution is first computed using existing continuation software—MATCONT [11] in MATLAB and AUTO-07p [12] in FORTRAN.
 These tools provide not only the periodic orbit but also the Jacobian along the trajectory, enabling us to study the associated variational system.
-We test three discretization and eigensolver combinations on the coupled Stuart-Landau system eq.˜51, as listed in table˜2.
+We test three discretization and eigensolver combinations on the coupled Stuart-Landau system Eq.˜36, as listed in Table˜2.
 Table 2: Comparison of Software, Discretization Methods, and Algorithms
@@
 In comparison, our Gear4 method performs worse.
-The reason is as the period increases and the smallest multiplier approaches zero, the spectral gap between Floquet multipliers and spurious eigenvalues decreases.
-According to theorem˜1, this leads to the separation between Floquet subspace and spurious subspaces decreases, and thus reducing accuracy of each Floquet multipliers.
+The reason is as the period increases and the smallest multiplier approaches zero, the spectral gap between Floquet multipliers and parasitic eigenvalues decreases.
+This leads to the separation between Floquet subspace and parasitic subspaces decreases, and thus reducing accuracy of each Floquet multipliers.
 Although the accuracy of the oscillatory multiplier computed by our method is slightly inferior, we are still able to maintain its value close to 1.
@@
 In radio frequency (RF) circuits, accurately predicting the phase noise of oscillators is essential for evaluating the performance and stability of oscillatory systems [9].
-A critical step in this prediction involves computing the perturbation projection vector (PPV) [5], which corresponds to the eigenvector associated with the dominant Floquet multiplier of an LPTV system.
-The LPTV system takes the form:
-$C(t)^{\mathrm{T}}\frac{\mathrm{d}y}{\mathrm{d}t}+G(t)^{\mathrm{T}}y=0,$ (52)
+A critical step in this prediction involves computing the perturbation projection vector (PPV) [9], which corresponds to the eigenvector associated with the dominant Floquet multiplier of an LTP system.
+The LTP system takes the form:
+(37) $C(t)^{\mathrm{T}}\frac{\mathrm{d}y}{\mathrm{d}t}+G(t)^{\mathrm{T}}y=0,$
 where $C(t)$ and $G(t)$ are $T$-periodic matrices with the following structure:
@@
 The vector $y_{1}$ represents the part of the nodal voltages, and $y_{2}$ denotes the rest nodes.
-According to the DAE theory presented in [27], equation eq.˜52 is classified as a semi-explicit index-1 differential-algebraic equation (DAE).
+According to the DAE theory presented in [21], equation Eq.˜37 is classified as a semi-explicit index-1 differential-algebraic equation (DAE).
 It can be systematically decoupled into an ODE and an algebraic constraint:
-$\left\{\begin{aligned} &-\frac{\mathrm{d}}{\mathrm{d}t}y_{1}(t)+C_{11}^{-\mathrm{T}}(t)\left[G_{11}^{\mathrm{T}}(t)-G_{21}^{\mathrm{T}}(t)G_{22}^{-T}(t)G_{12}^{\mathrm{T}}(t)\right]y_{1}=0,\\ &y_{2}(t)=-G_{22}^{-T}(t)G_{12}^{\mathrm{T}}(t)y_{1}(t).\end{aligned}\right$.
-(53)
-Our task is to solve the Floquet multiplier closest to 1 and its eigenvectors of the ODE part of eq.˜53, and thus recover the PPV.
-The physical nature of RF circuits leads to the following three characteristics of eq.˜53:
+(38) $\left\{\begin{aligned} &-\frac{\mathrm{d}}{\mathrm{d}t}y_{1}(t)+C_{11}^{-\mathrm{T}}(t)\left[G_{11}^{\mathrm{T}}(t)-G_{21}^{\mathrm{T}}(t)G_{22}^{-T}(t)G_{12}^{\mathrm{T}}(t)\right]y_{1}=0,\\ &y_{2}(t)=-G_{22}^{-T}(t)G_{12}^{\mathrm{T}}(t)y_{1}(t).\end{aligned}\right$.
+Our task is to solve the Floquet multiplier closest to 1 and its eigenvectors of the ODE part of Eq.˜38, and thus recover the PPV.
+The physical nature of RF circuits leads to the following three characteristics of Eq.˜38:
   1. 1.
@@
 In Fig.˜5(a), the Floquet multipliers are plotted in blue relative to the unit circle in black. Most multipliers are clustered near the origin, except for a single one lying close to 1.
-According to Theorem theorem˜1, the subspace separation between dominant $k$-dimension subspace (the invariant subspace of the dominant $k$ eigenvalues) and the remainder is bounded above by the spectral gap between the $k$-th and $(k+1)$-th largest eigenvalues.
+The subspace separation between dominant $k$-dimension subspace (the invariant subspace of the dominant $k$ eigenvalues) and the remainder is bounded above by the spectral gap between the $k$-th and $(k+1)$-th largest eigenvalues.
 Consequently, the dominant one-dimensional subspace is expected to be well isolated, while this separation rapidly decay as more multipliers are included.
@@
 7 Conclusion
-A multistep discretization of an LPTV system gives rise to a periodic-PEP for computing Floquet multipliers and their invariant subspaces.
-This periodic-PEP exhibits clear spectral separation: as the stepsize tends to zero, the dominant eigenvalues converge to the true Floquet multipliers, while the remaining ones decay to zero.
+A multistep discretization of an LPTV system gives rise to a pPEP for computing Floquet multipliers and their invariant subspaces.
+This pPEP exhibits clear spectral separation: as the stepsize tends to zero, the dominant eigenvalues converge to the true Floquet multipliers, while the remaining ones decay to zero.
 By subspace perturbation theory, the computed multipliers and invariant subspaces converge to their exact counterparts at a rate matching the consistency order of the multistep method.
@@
 Numerical experiments demonstrate that the algorithm effectively addresses the two common scenarios, as outlined in the introduction.
-Declarations
-  • Funding: The research leading to these results received funding from National Key R&D Program of China under Grant Agreement No 2020YFA0711900, 2020YFA0711902.
-  • Conflict of interest/Competing interests (check journal-specific guidelines for which heading to use): The authors have no conflicts of interest to declare that are relevant to the content of this article.
-  • Ethics approval and consent to participate: Not applicable
-  • Consent for publication:Not applicable.
-  • Data availability: The datasets generated during and/or analyzed during the current study are available from the corresponding author on reasonable request.
-  • Materials availability: Not applicable.
-  • Code availability:The code used for the numerical experiments is available from the corresponding author upon reasonable request.
-  • Author contribution: Yehao Zhang led the research and contributed to all aspects of this work, including the theoretical development, algorithm design, and numerical experiments.
-Yuncheng Xu and Yichen Tan participated in the theoretical discussions and provided suggestions for the design of the numerical experiments.
-Yangfeng Su supervised the research and provided guidance on the overall structure and direction of the study.
-All authors reviewed and approved the final manuscript.
-If any of the sections are not relevant to your manuscript, please include the heading and write ‘Not applicable’ for that section.
+Acknowledgments
+We would like to acknowledge the assistance of volunteers in putting together this example manuscript and supplement.